### v1

# Parametric Mean-Field Empirical Bayes in High-dimensional Linear Regression

Abstract
In this paper, we consider the problem of parametric empirical Bayes estimation of an i.i.d. prior in high-dimensional Bayesian linear regression, with random design. We obtain the asymptotic distribution of the variational Empirical Bayes (vEB) estimator (motivated from [10] and [27]) which approximately maximizes a variational lower bound of the intractable marginal likelihood. We characterize a sharp phase transition behavior for the vEB estimator — namely that it is information theoretically optimal (in terms of limiting variance) up to $p=o(n^{2/3})$ while it suffers from a sub-optimal convergence rate in higher dimensions. In the first regime, i.e., when $p=o(n^{2/3})$, we show how the estimated prior can be calibrated to enable valid coordinate-wise and delocalized inference, both under the empirical Bayes posterior and the oracle posterior. In the second regime, we propose a debiasing technique as a way to improve the performance of the vEB estimator beyond $p=o(n^{2/3})$. Extensive numerical experiments corroborate our theoretical findings.
high-dimensional statistics,
keywords:
[class=MSC]
keywords:
1 Introduction
Consider a canonical linear regression model
$y\mid X,\beta\sim N(X\beta,\sigma^{2}I_{n}),$ (1)
where $y\in\operatorname{\mathbb{R}}^{n}$ denotes the responses, $X\in\operatorname{\mathbb{R}}^{n\times p}$ denotes the design matrix, $\beta\in\operatorname{\mathbb{R}}^{p}$ is the unknown regression coefficients, and $\sigma^{2}>0$ is the error variance (potentially unknown). We work under a high-dimensional asymptotic setting, where both the sample size $n$ and the number of parameters $p$ grow to infinity. Here, we additionally posit an i.i.d. prior on the coefficient vector $\beta=(\beta_{1},\ldots,\beta_{p})$, where the (random) coefficients ($\beta_{i}$) are generated i.i.d. from an unknown prior distribution $\mu_{\theta_{0}}$, and $\theta_{0}\in\Theta\subset\mathbb{R}^{k}$ is a (deterministic) finite-dimensional parameter. These models arise naturally in estimating effect size distributions and in applications involving genetic associations; see e.g. [63, 41, 65]. Our goal in this paper is two-fold:
  1. (a)
    To estimate the prior $\mu_{\theta}$ in a data-driven manner, and
  2. (b)
    To use the estimated prior for downstream inference under the oracle posterior on $\beta$.
While traditional Bayesian inference assumes that the prior $\mu_{\theta}$ is specified by the statistician, this can be problematic for high-dimensional noisy data. For example, assume the simple high-dimensional setting where $p=n$ and $X$ is the $n\times n$ identity matrix. This is the classical Gaussian sequence model
$y_{i}\overset{i.i.d.}{\sim}N(\beta_{i},\sigma^{2}).$
In this case, the posterior distribution of $\beta\mid y,X$, witnesses proportional contribution from the prior and the likelihood, which implies that there is no posterior contraction. In other words, all posterior based inference depends heavily on the prior, even asymptotically as $n\to\infty$. A misspecified prior from the statistician would therefore result in inconsistent results downstream. This motivates the development of procedures that can consistently estimate $\mu_{\theta_{0}}$. Empirical Bayes (EB) addresses this problem by using the observed data to estimate the prior. For simplicity, suppose the variance $\sigma^{2}$ is known. A natural strategy for estimating $\mu_{\theta_{0}}$ is to maximize the marginalized likelihood
$m_{\theta}(y):=$ $\Big(\frac{1}{2\pi\sigma^{2}}\Big)^{\frac{n}{2}}\int\exp\Big(-\frac{1}{2\sigma^{2}}\|{y}-{X}{\beta}\|^{2}\Big)\prod_{i=1}^{p}d\mu_{\theta}(\beta_{i})$ (2)
with respect to the ‘hyperparameter’ $\theta$. However, except in extreme cases (such as when $\mu_{\theta}$ is a normal distribution or $X^{\top}X$ is diagonal), the integral in (2) does not have an explicit form. The problem is particularly challenging under the high-dimensional setting with $p\to\infty$, as the marginal likelihood in (2) is extremely challenging to compute by standard numerical integration or sampling-based approaches.
Recent works (see [10, 27, 38, 20]) have addressed this issue by maximizing a more tractable variational lower bound, also commonly known as the evidence lower bound (ELBO), instead of (2), to obtain a variational empirical Bayes (vEB) estimator. Necessary conditions for the consistency of this estimator was recently established in [38]. However, to our knowledge, the explicit convergence rates, asymptotic distributions, and asymptotic efficiency of vEB estimators have not been studied in the literature. In this work, we take a first step towards this direction by analyzing an approximate vEB estimator in the parametric setting (see [10, 21, 22]). We provide a complete characterization of its asymptotic behavior when $p=o(n)$, under appropriate scaling of the design matrix. We then use the estimated prior for coordinate-wise and delocalized inference under the oracle posterior.
1.1 Organization and summary of main results
This work first analyzes frequentist inferential guarantees of the vEB estimator $\hat{\theta}^{\mathsf{vEB}}$ that maximizes a variational approximation of the marginal likelihood $m_{\theta}(y)$, assuming a random design matrix $X$ and regularity conditions for the prior parametric family. Interestingly, the optimality of $\hat{\theta}^{\mathsf{vEB}}$ exhibits a phase transition depending on the asymptotic scaling of the problem. When $p\ll n^{2/3}$, the vEB estimator is $\sqrt{p}$-consistent, asymptotically normal, and attains the information-theoretically optimal variance. In contrast, all of these properties fail when $p\gg n^{2/3}$. This is formally stated in the following theorem, which combines the separate upper and lower bound claims in Section 3. While the following result is stated under a correctly specified parametric family for the prior, we also derive analogous limit distributions under misspecified priors (see 3.4).
Theorem.
Suppose $y$ is generated from the Bayesian regression model (1) with a true prior parameter $\theta_{0}$. Assume a random design matrix with a low signal to noise ratio (SNR) scaling, alongside standard regularity conditions on the prior (see Assumptions 3.13.4). Let $\hat{\theta}^{\mathsf{vEB}}$ be the vEB estimator (see (11)). Under the asymptotic regime $p=o(n^{2/3})$, $\hat{\theta}^{\mathsf{vEB}}$ is $\sqrt{p}$-consistent with a limiting normal distribution, and attains the information-theoretic optimal limiting variance. However, when $p\gg n^{2/3}$, $\hat{\theta}^{\mathsf{vEB}}$ is no longer optimal and is only $n/p$-consistent with a non-normal limit.
Moving beyond frequentist recovery of the prior hyperparameter $\theta$, we illustrate in Section 4 that the vEB estimator can be used for downstream Bayesian inference on the regression coefficients $\beta$, under the asymptotic regime $p=o(n^{2/3})$. First, we bound the total variation distance between coordinate-wise distributions under the EB posterior (constructed using the estimated prior $\mu_{\hat{\theta}^{\mathsf{vEB}}}$) and the oracle posterior (constructed using the oracle prior $\mu_{\theta_{0}}$). This in particular implies that the mean/variance computed under the variational approximation of the EB posterior are consistent. Next, we address uncertainty quantification under the two posteriors. We mainly focus on the one-dimensional projection of the posterior, that is $q^{\top}\beta\mid y,X$, for some fixed vector $q\in\mathbb{R}^{p}$. Contrary to previous studies in high SNR settings which show that inference under the oracle posterior and the EB posterior are asymptotically equivalent (the so-called higher-order merging phenomena in [45]), we illustrate that this is not guaranteed in the low SNR setting. More specifically, the one-dimensional projection of the oracle posterior and the EB posterior merge if and only if the projection direction is an approximate contrast vector (i.e. $\frac{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}}{\|q\|}=o(\sqrt{p})$), illustrating the need for caution for utilizing the EB posterior. Nonetheless, we provide EB adjusted credible intervals for such one-dimensional projections which provide valid marginal coverage under the oracle posterior, using the estimated prior hyperparameter $\hat{\theta}^{\mathsf{vEB}}$.
In Section 5, we verify satisfactory finite-sample performance of our theoretical results, under various priors that include common choices such as spike-and-slab and Gaussian mixtures. In Section 5.4, we also illustrate that the slower convergence rate of $\hat{\theta}^{\mathsf{vEB}}$ in higher-dimensions can be improved by debiasing. To elaborate, we show that the debiased vEB estimator relaxes the requirement $p\ll n^{2/3}$ for the aforementioned normal limit to $p\ll n^{3/4}$. We conclude the paper in Section 6 by discussing avenues for future work.
1.2 Related works
Standard EB in denoising.
The standard EB setup for normal means [this corresponds to taking $X=I_{n}$ in (1), see e.g., 46, 24, 25, 51] lead to a product posterior, and hence results in independent observations under the posterior. This allows the marginal likelihood to factorize, and does not require a high-dimensional integral computation as in (2). Additionally, the exact factorization of the posterior results in tractable representations for individual coordinates of the posterior mean and variance, using Tweedie’s formula (see e.g. [46]). In contrast, our regression setup does not have any such representations of the posterior moments due to the intractable likelihood, creating additional theoretical challenges for downstream inference.
Statistical efficiency of MML for EB.
As stated earlier in (2), maximizing the marginal likelihood is a natural approach to empirical Bayes. Consequently, asymptotic properties of parametric MML was previously studied in the literature [42, 53, 47, 45]. In particular, [42] established consistency of the MML estimator to an “oracle set” as well as the consistency of the EB posterior. The recent work [45] additionally shows a higher-order merging of the EB posterior to the oracle posterior, which indicates that the EB posterior credible intervals enjoy coverage under the oracle posterior. However, this line of work considered the more classical finite-dimensional asymptotic scaling with high SNR. More importantly, they posit a frequentist true model that corresponds to fixing any true parameter $\beta^{\star}$ in the regression case, thus making the Bayesian model formulation to be misspecified. This is in sharp contrast to our high-dimensional low SNR scaling with correct model specification. Indeed, we will show that such higher-order merging may not happen under our setup. We also stress that going beyond consistency, no asymptotic normality or efficiency results have previously been established for the MML estimator, beyond simple settings.
Gradient flow based methods.
Another direction for empirical Bayes methods in regression involves using gradient flows to solve the complex optimization problem in (2) [31, 20, 21, 22]. In particular, the recent line of works [21, 22] consider a very similar problem setup as ours, but instead analyze an adaptive Langevin dynamics estimator under proportional asymptotics $p\asymp n$. We believe these results complement ours, as we analyze the vEB estimator, consider the scaling $p=o(n)$, and impose different assumptions for priors. In Section 5, we compare our vEB estimator with the adaptive Langevin algorithm from [21, 22]. We also note that in the parametric EB setup, no consistency guarantees have been established for the gradient flow estimator.
vEB for general latent variable models.
To our knowledge, vEB dates back to [9], which first proposed the notion of maximizing the variational approximation lower bound of the intractable likelihood in the context of topic modeling. This idea was also utilized for unsupervised learning in the context of variational autoencoders (VAEs, [28]) and item response theory [14] as well as various modern machine learning problems such as matrix factorization [60, 64], due to their excellent scalability. However, the theoretical understanding of such vEB procedures is limited, in particular for precise convergence rates and limiting distributions, which is precisely the gap we aim to fill in this paper.
1.3 Notation
[…]
$\mathrm{D_{KL}}(\mathbb{Q}\mid\operatorname{\mathbb{P}}):=\begin{cases}\operatorname{\mathbb{E}}_{\mathbb{Q}}\log{\frac{d\mathbb{Q}}{d\operatorname{\mathbb{P}}}}&\mbox{if }\mathbb{Q}\ll\operatorname{\mathbb{P}},\\ \infty&\mbox{otherwise},\end{cases}$
and the Total Variation (TV) distance as
$\mathrm{D_{TV}}(\operatorname{\mathbb{P}},\mathbb{Q}):=\sup_{E\subseteq\mathbb{R}}\left|\operatorname{\mathbb{P}}(E)-\mathbb{Q}(E)\right|.$
Let $I_{n}$ denote the $n\times n$ identity matrix. For two matrices $V_{1},V_{2}$, write $V_{1}\succ V_{2}$ when $V_{1}-V_{2}$ is positive definite. Let $\|\cdot\|$ denote the usual $L^{2}$ norm for matrices/vectors. We use the usual Bachmann-Landau notations. In particular, for deterministic sequences $\{a_{p}\}_{p\geq 1}$, $\{c_{p}\}_{p\geq 1}$ (with $c_{p}>0$) and a constant $r>0$, we write $a_{p}\lesssim_{r}c_{p}$ when there exists a constant $K(r)$ such that $a_{p}\leq K(r)c_{p}$. Write $a_{p}\ll c_{p}$ if $a_{p}=o(c_{p})$. Given a matrix/tensor $M_{p}$ we will use $\|M_{p}\|_{F}$ to denote its Frobenius norm. We also extend the $o/O$ notation to vectors/matrices/tensors: for a sequence of vectors/matrices/tensors $\{M_{p}\}_{p\geq 1}$ (with fixed dimensions), write $M_{p}=O(c_{p})$ when $\|M_{p}\|_{F}=O(c_{p})$, and write $M_{p}=o(c_{p})$ when $\|M_{p}\|_{F}=o(c_{p})$. We also use the notations $o_{P},O_{P}$, and in particular write $o_{P,X}(1)$ and $\mathnormal{O}_{\mathnormal{P},X}(1)$ to capture the randomness of a random variable $X$. The notations $\overset{d}{\longrightarrow},\overset{P}{\longrightarrow}$ denotes convergence in distribution and convergence in probability, respectively. For a vector $v\in\mathbb{R}^{k}$, write $v^{2}:=v\otimes v$ and $v^{3}:=v\otimes v\otimes v$, where $\otimes$ denotes the outer product. For a random variable $Z$, define its sub-Gaussian norm $\|Z\|_{\psi_{2}}$ as the smallest $K\geq 0$ that satisfies
$\operatorname{\mathbb{E}}e^{t(Z-\operatorname{\mathbb{E}}Z)}\leq e^{K^{2}t^{2}/2},\quad\forall t\in\mathbb{R}.$
Throughout the paper, we work on an asymptotic setting where both $n,p\to\infty$. Hence, we understand $n\equiv n(p)$ as a function of $p$ and write limits in terms of $p$. All probabilities and expectations are conditioned on $X,\beta^{\star}$ unless specified otherwise. When the expectation is respect to $X$ or $\beta^{\star}$, we write $\operatorname{\mathbb{E}}_{X}$ or $\operatorname{\mathbb{E}}_{\beta^{\star}}$, respectively. We also define the notion of linear-quadratic tilts for probability measures.
Definition 1.1 (Linear-quadratic tilt).
Given a base-measure $\mu_{\theta}$ and constants $d_{i}$ for each $i=0,1,\ldots,p$ (defined below in (3) and 3.1), let $\mu_{i,\theta}$ be a quadratic tilt of $\mu_{\theta}$ by setting the likelihood ratio as
$\frac{d\mu_{i,\theta}}{d\mu_{\theta}}(b):=e^{-\frac{d_{i}}{2}b^{2}-C_{\theta}(d_{i})},\quad\text{where}\quad C_{\theta}(d_{i}):=\log\int e^{-\frac{d_{i}}{2}b^{2}}d\mu_{\theta}(b).$
Next, for $t\in\mathbb{R}$, let $\mu_{i,t,\theta}$ be the linear tilt of $\mu_{i,\theta}$, i.e.,
$\frac{d\mu_{i,t,\theta}}{d\mu_{i,\theta}}(b):=e^{tb-\psi_{i,\theta}(t)},\quad\text{where}\quad\psi_{i,\theta}(t):=\log\int e^{tb}d\mu_{i,\theta}(b).$
By properties of exponential families, the mean and variance of $\mu_{i,t,\theta}$ are $\psi_{i,\theta}^{\prime}(t)$ and $\psi_{i,\theta}^{\prime\prime}(t)$, respectively.
2 Variational empirical Bayes
For our goal of estimating $\theta$, the natural strategy is to maximize the marginal likelihood $m_{\theta}(y)$ (see (2)) with respect to $\theta$. We first define some transformations of the data $(X,y)$, which will be useful in simplifying $m_{\theta}(y)$.
Definition 2.1 (Transformed variables).
Define $p$-dimensional vectors $w=(w_{1},\ldots,w_{p})$ and $d=(d_{1},\ldots,d_{p})$ as follows:
$w:=\frac{X^{\top}y}{\sigma^{2}},\quad d_{i}:=\frac{(X^{\top}X)(i,i)}{\sigma^{2}}\quad\text{for}\;\;i=1,\ldots,p.$ (3)
Also, let $A_{p}\in\operatorname{\mathbb{R}}^{p\times p}$ be the off-diagonal part of the matrix $\sigma^{-2}X^{\top}X$ (with diagonal entries set to zero), that is
$A_{p}(i,j)=-\frac{(X^{\top}X)(i,j)}{\sigma^{2}}\quad\text{if}\quad 1\leq i\neq j\leq p\quad\mbox{and}\quad A_{p}(i,i)=0\quad\text{if}\quad 1\leq i\leq p.$
As $d$ and $A_{p}$ respectively denote the diagonal and off-diagonal parts of $\sigma^{-2}X^{\top}X$, we can write
$\sigma^{-2}X^{\top}X=\text{{Diag}}(d)-A_{p}.$
Using these notation, we expand the quadratic term in the exponent of (2):
$\exp\Big(-\frac{1}{2\sigma^{2}}\|y-X\beta\|^{2}\Big)=\exp\Big(-\frac{\|y\|^{2}}{2\sigma^{2}}-\frac{1}{2}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}d_{i}\beta_{i}^{2}+\frac{\beta^{\top}A_{p}\beta}{2}+w^{\top}\beta\Big).$
Now, by absorbing each summand $d_{i}\beta_{i}^{2}$ to the base measure $\mu_{\theta}$ and using the notations $C_{\theta}(d_{i}),\mu_{i,\theta}$ in 1.1, we can re-write the scaled log-likelihood as
$\ell_{p}(\theta):=\frac{1}{p}\log m_{\theta}(y)=-\frac{n}{2p}\log(2\pi\sigma^{2})-\frac{1}{2p\sigma^{2}}\lVert y\rVert^{2}+\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}C_{\theta}(d_{i})+\frac{1}{p}\log Z_{p}(w,\theta),$ (4)
where
$Z_{p}(w,\theta):=\int\exp\Big(\frac{\beta^{\top}A_{p}\beta}{2}+w^{\top}\beta\Big)\prod_{i=1}^{p}d\mu_{i,\theta}(\beta_{i}).$ (5)
We motivate the vEB (variational empirical Bayes) estimator, initially assuming a known variance $\sigma^{2}>0$. While maximizing $\ell_{p}(\theta)$ as a function of $\theta$ is tempting, the normalizing constant $Z_{p}(w,\theta)$ in (5) involves an intractable high-dimensional integral which makes the maximization computationally challenging. To remedy this issue, we build our estimator of $\theta_{0}$ using ideas from Mean-Field variational inference. To wit, first observe that the Gibbs variational formula implies the following dual representation of $\log Z_{p}(w,\theta)$ (c.f. [59, 39]):
$\log Z_{p}(w,\theta)$ $=\sup_{\mathbb{Q}\in\mathcal{P}(\mathbb{R}^{p})}\left(\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\frac{1}{2}\beta^{\top}A_{p}\beta+w^{\top}\beta\right]-\mathrm{D_{KL}}\big(\mathbb{Q}|\prod_{i=1}^{p}\mu_{i,\theta}\big)\right),$
where $\mathcal{P}(\mathbb{R}^{p})$ denotes the set of all probability measures on $\mathbb{R}^{p}$. Now, by restricting $\mathcal{P}(\mathbb{R}^{p})$ to the set of product measures $\mathbb{Q}=\prod_{i=1}^{p}Q_{i}$, we get a lower bound for $\log Z_{p}(w,\theta)$:
$\log Z_{p}(w,\theta)$ $\geq\sup_{\mathbb{Q}=\prod_{i=1}^{p}Q_{i}}\left(\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\frac{1}{2}\beta^{\top}A_{p}\beta+w^{\top}\beta\right]-\mathrm{D_{KL}}\big(\mathbb{Q}|\prod_{i=1}^{p}\mu_{i,\theta}\big)\right)$ (6)
$=\sup_{u\in\mathbb{R}^{p}}\Bigg(\underbrace{\frac{u^{\top}A_{p}u}{2}+w^{\top}u-\sum_{\mathnormal{i}=1}^{\mathnormal{p}}I_{i,\theta}(u_{i})}_{=:M_{\theta}(u)}\Bigg),$ (7)
where $I_{i,\theta}(t):=\mathrm{D_{KL}}(\mu_{i,(\psi_{i,\theta}^{\prime})^{-1}(t),\theta}|\mu_{i,\theta})$ denotes the KL divergence between a linear tilt (with mean $t$) and the base measure of the $i$th coordinate. Here (7) follows from the fact that the supremum in (6) is attained by linear tilts $\mu_{i,\cdot,\theta}$, which allows reducing the optimization over measures to that over reals (see [39, 62]). Note that the vector $u$ effectively denotes the coordinate-wise mean of the product measure $\mathbb{Q}$.
Hence, by plugging-in (7) to (4), the re-scaled marginal log-likelihood $\ell_{p}(\theta)$ can be lower bounded by the “Mean-Field” log-likelihood $\ell_{p}^{\text{MF}}(\theta)$ (equivalently, a re-scaled ELBO):
$\ell_{p}(\theta)\geq\ell_{p}^{\text{MF}}(\theta):=-\frac{n}{2p}\log(2\pi\sigma^{2})-\frac{1}{2p\sigma^{2}}\lVert y\rVert^{2}+\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}C_{\theta}(d_{i})+\frac{1}{p}\sup_{u\in\mathbb{R}^{p}}M_{\theta}(u).$ (8)
The supremum in (8) is still an implicit function of $\theta$. We therefore approximate it further to reduce to a more tractable optimization problem in $\theta$.
Definition 2.2 (Mean-Field optimizers).
Under 3.1, by [34, Lemma 2.1], the optimizer attaining the supremum in (8) is unique. Let us denote it by $u_{\theta}=(u_{1,\theta},\ldots,u_{p,\theta})$. Then $u_{\theta}$ satisfies ${u}_{i,\theta}=\psi_{i,\theta}^{\prime}({s}_{i,\theta}+w_{i})$, where $s_{i,\theta}:=\sum_{j\neq i}A_{p}(i,j){u}_{i,\theta}$, for $i\leq p$.
For heuristics, assume that $\mu_{\theta}$ is compactly supported. Then under our random design condition (see 3.1 below), direct computations show that the empirical measure $p^{-1}\sum_{i=1}^{p}\delta_{s_{i,\theta}}$ converges weakly to 0. It is then natural to define the approximate optimizer $\tilde{u}_{\theta}=(\tilde{u}_{1,\theta},\ldots,\tilde{u}_{n,\theta})$ as $\tilde{u}_{i,\theta}:=\psi^{\prime}_{i,\theta}(w_{i}).$ As a result, by spelling out the KL divergences in $I_{i,\theta}$, we can write
$\frac{1}{p}\sup_{u\in\mathbb{R}^{p}}M_{\theta}(u)\approx\frac{\tilde{u}_{\theta}^{\top}A_{p}\tilde{u}_{\theta}}{2p}+w^{\top}\tilde{u}_{\theta}-\sum_{\mathnormal{i}=1}^{\mathnormal{p}}I_{i,\theta}(\tilde{u}_{i,\theta})\approx\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\psi_{i,\theta}(w_{i}),$
where $\psi_{i,\theta}$ is defined in 1.1. Now, define functions $F_{i}:\Theta\to\mathbb{R}$ for each $i\leq p$ as
$F_{i}(\theta):=\log\int e^{w_{i}\beta_{i}-\frac{d_{i}\beta_{i}^{2}}{2}}d\mu_{\theta}(\beta_{i}).$ (9)
Note that $F_{i}(\theta)=C_{\theta}(d_{i})+\psi_{i,\theta}(w_{i})$. By plugging in the above observations into the lower bound (8), we have
$\ell_{p}(\theta)\geq\ell_{p}^{\text{MF}}(\theta)\approx-\frac{n}{2p}\log(2\pi\sigma^{2})-\frac{1}{2p\sigma^{2}}\lVert y\rVert^{2}+\frac{1}{p}F_{i}(\theta).$ (10)
Therefore maximizing the above (approximate) lower bound, we get our vEB estimator
$\boxed{\hat{\theta}^{\mathsf{vEB}}=\operatorname*{arg\,max}_{\theta\in\Theta}\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}(\theta).}$ (11)
Remark 2.1 (Unknown variance).
So far, the vEB estimator has required the knowledge of the variance $\sigma^{2}>0$. Fortunately, even though consistent estimation for $\beta$ is impossible under the low SNR scaling with $\sigma^{-2}\|X^{\top}X\|=O(1)$, it is straightforward to consistently estimate $\sigma^{2}$ using the usual MSE estimator:
$\hat{\sigma}^{2}:=\frac{y^{\top}(I_{n}-X(X^{\top}X)^{-1}X^{\top})y}{n-p}.$ (12)
Thus, by plugging-in the above variance estimate, the vEB estimator can be further extended. Alternatively, one may choose to maximize the approximate likelihood in (10) with respect to both $(\theta,\sigma^{2})$, as suggested in [27].
3 Prior recovery via vEB and Fisher optimality
Now, we illustrate our main results regarding the statistical performance of $\hat{\theta}^{\mathsf{vEB}}$. We discuss its limiting distribution in Section 3.1, extension to the misspecified setting in Section 3.2, and its information-theoretic optimality in Section 3.3. To present our theoretical results, we require several assumptions. First, to understand dimension dependence, we work under the following random design matrix with independent entries.
Assumption 3.1 (Random design).
For some distribution $P$ with mean zero, variance one, and sub-Gaussian norm bounded by $K>0$, suppose that $X_{k,i}\overset{i.i.d.}{\sim}\frac{1}{\sqrt{n}}P$ for all $k\leq n,i\leq p$. We will work under the asymptotic scaling $p=o(n)$.
Here the $n^{-1/2}$ scaling for $X$ is chosen so that $\|X^{\top}X\|/\sigma^{2}=\Theta(1)$, placing the model in a low signal to noise ratio (SNR) regime where individual regression coefficients $\beta$ are not consistently estimable. In contrast, when the prior belongs to a low-dimensional parametric family, its hyperparameters remain identifiable through aggregation across coordinates. This makes our high-dimensional scaling particularly natural for empirical Bayes analysis. Such regimes are standard in theoretical studies of high-dimensional Bayesian regression [44, 11, 3, 33, 48], and in particular appealing for EB problems that analyze the consequences of estimating the prior [38, 20, 22]. For simplicity, we also assume that the regression variance $\sigma^{2}$ is known; we expect that all following results will go through when this is replaced by the plug-in estimator in (12) (since it is consistent at a much faster $\sqrt{n}$-rate; also see Remark 2.1).
Next, we impose the following regularity conditions on the prior, which are standard assumptions for efficient estimation in parametric models [e.g. see Theorems 2.6 or 5.1 in 35]. For example, exponential families satisfy these conditions.
Assumption 3.2 (Prior regularity).
Suppose that the prior $\mu_{\theta}$ is a parametric family with log-likelihood $\ell(\theta;\beta)$ with respect to a non-degenerate base measure $\nu$, such that satisfies the following regularity conditions:
  1. R0.
    $\mu_{\theta}$ is identifiable and has common support.
  2. R1.
    The parameter space $\Theta\subset\mathbb{R}^{k}$ is a compact set for some fixed integer $k$, and the true parameter $\theta_{0}$ lives in the interior of $\Theta$.
  3. R2.
    For all $\beta$ and $\theta\in\Theta$, the log-likelihood $\ell(\theta;\beta)$ is three-times differentiable.
  4. R3.
    The Bartlett’s identities for $\nabla\ell:=\frac{\partial\ell}{\partial\theta}$ and $\nabla^{2}\ell:=\frac{\partial^{2}\ell}{\partial\theta^{2}}$ hold (see conditions (g), (h) in Theorem 2.6, [35]), and the Fisher information $I(\theta):=\operatorname{\textnormal{Var}}_{B_{\theta}\sim\mu_{\theta}}[\nabla\ell(\theta;B_{\theta})]$ is positive definite for all $\theta\in U$.
In passing, we mention that R2 requires the derivatives to exist for all $\theta$, compared to the usual regularity condition that assumes differentiability only near $\theta_{0}$. This is mainly for technical convenience.
Finally, in this Section, we assume a well-specified model, in the sense that the prior parametric family $\mu_{\theta}$ and the regression variance $\sigma^{2}$ are both correct. We will discuss extensions to misspecified priors later in Section 3.2.
Assumption 3.3 (Well-specified model).
Assume that the data $y$ is generated from a correctly specified model with true prior parameter $\theta_{0}\in\textnormal{int}(\Theta)$. In other words, assume that $y$ is generated from the same model used to construct the posterior:
$\beta_{i}^{\star}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu_{\theta_{0}},\quad y\mid X,\beta^{\star}\sim N(X\beta^{\star},\sigma^{2}I_{n}).$
3.1 Upper bound
To analyze $\hat{\theta}^{\mathsf{vEB}}$, we first introduce additional notations that will be used to describe the limiting distribution.
Definition 3.1.
  1. (a)
    Recalling the regression variance $\sigma^{2}$, set $d_{0}:=\sigma^{-2}$.
  2. (b)
    For any $\theta\in\Theta$, define a pair of random variables $(B_{\theta},W_{\theta})$ by setting
    $B_{\theta}\sim\mu_{\theta},\quad W_{\theta}\mid B_{\theta}\sim N(d_{0}B_{\theta},~d_{0}).$
    Also define a deterministic mapping $V:\Theta\to\mathbb{R}^{k\times k}$ by setting
    $V(\theta):=I(\theta)-\operatorname{\mathbb{E}}\Big[\operatorname{\textnormal{Var}}\big(\nabla\ell(\theta;B_{\theta})\mid W_{\theta}\big)\Big]=\operatorname{\textnormal{Var}}\Big[\operatorname{\mathbb{E}}\big(\nabla\ell(\theta;B_{\theta})\mid W_{\theta}\big)\Big]\succ 0_{k\times k}.$ (13)
  3. (c)
    For $t\in\mathbb{R},~d>0$, let $B_{t,d,\theta}$ be a linear-quadratic tilted probability measure on $\mathbb{R}$ defined as $\frac{dB_{t,d,\theta}}{dB_{\theta}}(b)\propto e^{tb-\frac{db^{2}}{2}},$ so that $B_{\theta}\mid(W_{\theta}=t)\sim B_{t,d_{0},\theta}$. Finally, set $h_{\theta}:\mathbb{R}\to\mathbb{R}^{k}$ as $h_{\theta}(t):=\operatorname{\mathbb{E}}\nabla\ell(\theta;B_{t,d_{0},\theta})$ and define a mapping $\kappa:\Theta\to\mathbb{R}^{k}$ as
    $\kappa(\theta):=\frac{1}{2}\operatorname{\mathbb{E}}[B_{\theta}^{2}]\operatorname{\mathbb{E}}[h_{\theta}^{\prime\prime}(W_{\theta})].$
We motivate the notation $W_{\theta}$ and $V(\theta)$ in part (b) above, which will later play a crucial role as the Fisher information. First, recall from R3 in 3.2 that $I(\theta)=\operatorname{\textnormal{Var}}[\nabla\ell(\theta;B_{\theta})]$ is exactly the Fisher information for a single observed coefficient $B_{\theta}\sim\mu_{\theta}$. However, under the hierarchical model, we cannot directly observe the regression coefficient $\beta$ and we instead observe $y$. This is roughly equivalent to observing the transformed data $w$ in (3):
$w=\sigma^{-2}X^{\top}y\mid X,\beta^{\star}\stackrel{{\scriptstyle d}}{{\equiv}}N(d_{0}X^{\top}X\beta^{\star},d_{0}X^{\top}X).$ (14)
Approximating $X^{\top}X\approx I_{p}$, we can understand each coordinate $w_{i}$ as i.i.d. random variables with distribution $W_{\theta}$. Then, $V(\theta)$ can be understood as the Fisher information of the single sample $W_{\theta}$, which is formalized in the following lemma. Note that $V(\theta)\prec I(\theta)$ since $W_{\theta}$ has less information than $B_{\theta}$.
Lemma 3.1 (Fisher information).
Let $\tilde{\ell}(\theta;\cdot)$ be the log-likelihood of the random variable $W_{\theta}$. Then, the corresponding Fisher information satisfies:
$\operatorname{\mathbb{E}}[\nabla^{2}\tilde{\ell}(\theta;W_{\theta})]=V(\theta).$
The vector $\kappa(\theta)$ in part (c) is inevitably more technical. While we refrain from displaying the details, this quantity is crucial for understanding the mean of $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})$, which is the gradient of the objective function in (11). We will further motivate $\kappa(\theta)$ after presenting the main result (see the proof overview that precedes Remark 3.3).
To analyze the limiting distribution, we require additional technical assumptions for the prior in addition to the usual ones in 3.2.
Assumption 3.4 (Additional prior regularity).
Assume the following conditions on the prior likelihood $\ell(\theta;\beta)$ and tilted random variables $B_{t,d,\theta}$ (see part (iii) in 3.1).
  1. R4.
    The first three derivatives of the log-likelihood satisfy the following polynomial growth condition, where $r>0$ is some positive constant:
    $\|\nabla\ell(\theta;\beta)\|+\|\nabla^{2}\ell(\theta;\beta)\|_{F}+\|\nabla^{3}\ell(\theta;\beta)\|_{F}\lesssim 1+|\beta|^{r},\quad\forall\theta\in\Theta.$
  2. R5.
    All finite moments of $B_{\theta_{0}}$ exist.
  3. R6.
    The tilted random variables $B_{t,d,\theta}$ satisfy:
    $\sup_{d\in B_{\epsilon}(d_{0})}\sup_{\theta\in\Theta}\big|\operatorname{\mathbb{E}}[B_{0,d,\theta}]\big|\lesssim 1,\quad\sup_{d\in B_{\epsilon}(d_{0})}\sup_{\theta\in\Theta}\|B_{t,d,\theta}\|_{\psi_{2}}\lesssim 1+|t|,$
    where $B_{\epsilon}(d_{0}):=(d_{0}-\epsilon,d_{0}+\epsilon)$ for some $\epsilon<\frac{d_{0}}{2}$, and $\|\cdot\|_{\psi_{2}}$ denotes the sub-Gaussian norm.
Remark 3.1 (On Conditions R4 and R5).
R4 is a technical polynomial growth condition for the first three derivatives of the log-likelihood. Note that this can be easily satisfied since this is a requirement for functions which are already in the log-scale. This is required to control the remainder term in the Taylor expansion for analyzing the first-order condition of (11), and is reminiscent of the usual integrability assumption for $\nabla^{3}\ell(\theta;\beta)$. Here we need the additional regularity for $\nabla\ell(\theta;\beta)$ and $\nabla^{2}\ell(\theta;\beta)$ as the third derivative of $F_{i}(\theta)$ involves all three derivatives. Also, R5 is a standard moment assumption for the prior. Note that it is weaker than a sub-Gaussian assumption on the prior.
Remark 3.2 (On Condition R6).
R6 requires sub-Gaussian tails for all linear and quadratic tilts of $B_{\theta}$. This is weaker than assuming sub-Gaussianity directly on the prior $\mu_{\theta}$. Here, as $\Theta$ and $B_{\epsilon}(d_{0})$ are bounded sets, the uniform control over all $\theta,d$ is natural. On the other hand, the linear tilt parameter $t$ is unbounded, so instead of taking a supremum over $t$, we allow the sub-Gaussian norm to depend on $|t|$. This flexibility allows us to cover log-concave as well as non log-concave mixture priors (see Lemma 3.2 below). We note that our results go through with minor modifications if we replace the linear growth of the sub-Gaussian norm in R6 with a general polynomial order growth condition. The linear growth is primarily for expositional simplicity.
The following lemma illustrates concrete instances where conditions R5, R6 in 3.4 hold, which already generalizes existing assumptions in the literature. In particular, we can accommodate priors with unbounded support, mixture priors, and sparsity inducing priors. On the other hand, existing works on low SNR high-dimensional Bayesian regression require the prior to have (a) bounded support [38, 33, 20], (b) log-concave densities [32, 49], or (c) densities that satisfy the log-Sobolev inequality alongside a stronger version of our R4 [21]. These assumptions in the existing literature exclude common choices in practical EB methods, such as the adaptive shrinkage priors (this is essentially scale mixtures of Gaussians, see [52]).
Lemma 3.2.
Conditions R5, R6 in 3.4 hold when the prior $\mu_{\theta}$ satisfies any one of the following conditions.
  1. (a)
    $\mu_{\theta}$ has bounded support.
  2. (b)
    $\mu_{\theta}$ is a strongly log-concave density.
  3. (c)
    $\mu_{\theta}$ is a finite mixture of $\delta_{0}$ and Lebesgue densities which satisfy (R5), the first condition in (R6), and the stronger second condition
    $\sup_{t\in\mathbb{R}}\sup_{d\in B_{\epsilon}(d_{0})}\sup_{\theta\in\Theta}\|B_{t,d,\theta}\|_{\psi_{2}}\lesssim 1.$
    In particular, these conditions for the densities are satisfied by distributions from parts (a) and (b) above.
Now, we state our main result on limiting distributions of $\hat{\theta}^{\mathsf{vEB}}$.
Theorem 3.3.
Suppose Assumptions 3.1, 3.2, 3.3, 3.4 hold. Then, for general $p=o(n)$, we have $\hat{\theta}^{\mathsf{vEB}}-\theta_{0}=O_{P}\Big(\frac{1}{\sqrt{p}}+\frac{p}{n}\Big)$. In particular, the following limits hold.
  1. (a)
    When $p\ll n^{2/3}$, we have
    $\sqrt{p}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(0,V(\theta_{0})^{-1}\Big).$ (15)
  2. (b)
    When $\frac{p^{3/2}}{n}\to\delta\in(0,\infty)$, we have
    $\sqrt{p}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(\delta V(\theta_{0})^{-1}\kappa(\theta_{0}),V(\theta_{0})^{-1}\Big).$
  3. (c)
    In contrast, when $n^{2/3}\ll p\ll n$, we have
    $\frac{n}{p}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}V(\theta_{0})^{-1}\kappa(\theta_{0}).$
We elaborate on each part of 3.3. First consider the setting of part (a) which shows $\sqrt{p}$-consistency of the estimator, with a normal limit. This is the typical parametric rate expected in this problem, as justified in Remark 3.3 below. In our hierarchical model, as we only have indirect measurements of $\beta^{\star}=(\beta_{1}^{\star},\ldots,\beta_{p}^{\star})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu_{\theta_{0}}$, the limiting variance $V(\theta_{0})^{-1}$ is strictly larger than $I(\theta_{0})^{-1}$, the limiting variance for estimating $\theta_{0}$ directly from $\beta^{\star}$. We will further discuss the optimality of this limiting variance by deriving a matching lower bound in Section 3.3.
Part (b) considers higher dimensions with $p\propto n^{2/3}$. While the convergence rate and limiting variance is same as that in part (b), an additional bias appears in the limit. Note that this bias grows as $\delta=\lim_{p\to\infty}\frac{p^{3/2}}{n}$ increases. In part (c), we further increase the dimension so that $\delta\to\infty$, or equivalently $n^{2/3}\ll p\ll n$. The resulting limiting distribution exhibits a phase transition and results in a slower rate of convergence (since $\frac{n}{p}\ll\sqrt{p}$ under this regime). This is because now the bias term in part (b) dominates the variance.
To better understand this phase transition, we briefly discuss the proof strategy when $k=1$. Assuming that $\theta$ is an one-dimensional parameter, the first-order conditions for (11) gives
$\sqrt{p}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})\approx\frac{\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime}(\theta_{0})}{\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime\prime}(\theta_{0})}.$
We show that the denominator converges to $V(\theta_{0})$ as long as $p=o(n)$. To analyze the numerator, we additionally condition on the true coefficients $\beta^{\star}$ and show that
$\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime}(\theta_{0})\mid X,\beta^{\star}\approx N\big(\mu_{p}(X,\beta^{\star}),T_{2}\big),\quad\text{where}\quad\mu_{p}(X,\beta^{\star}):=\frac{1}{\sqrt{p}}\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime}(\theta_{0})\mid X,\beta^{\star}\Big]$ (16)
and $T_{2}>0$ is some constant variance (see (51)). The phase transition on $p$ arises from analyzing this mean term, where we use second-order Taylor expansions to write
$\mu_{p}(X,\beta^{\star})\approx\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi(d_{0}\beta_{i}^{\star})+\frac{p^{3/2}}{n}\kappa(\theta_{0})$ (17)
for some function $\Phi$ with $\operatorname{\mathbb{E}}\Phi(d_{0}\beta_{i}^{\star})=0$ (see (52), (53)). When $p\ll n^{2/3}$, the first summand in (17) dominates and gives the limit $\mu_{p}(X,\beta^{\star})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,T_{1})$. Combining this with the conditional normal limit (16) using the fact that $T_{1}+T_{2}=V(\theta_{0})$, we get $\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime}(\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,V(\theta_{0}))$. The conclusion in part (a) follows by Slutsky’s theorem.
In contrast, suppose $p\gg n^{2/3}$. Then, as the first term in the RHS of (17) is $O_{P}(1)$, the second term dominates. The resulting non-degenerate limit of $\mu_{p}$ requires a slower convergence rate:
$\frac{n}{p^{3/2}}\mu_{p}(X,\beta^{\star})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}\kappa(\theta_{0}).$
Consequently, we have $\frac{n}{p^{2}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime}(\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}\kappa(\theta_{0}),$ and the conclusion in part (c) follows again by Slutsky’s theorem.
Remark 3.3 ($\sqrt{p}$ is the optimal rate of convergence).
The convergence rate $\sqrt{p}$ is information-theoretically tight. To see this, consider the setting where one directly observes all regression coefficients $\beta^{\star}=(\beta_{1}^{\star},\ldots,\beta_{p}^{\star})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu_{\theta_{0}}$. The resulting convergence rate for estimating $\theta_{0}$ is $\sqrt{p}$, and hence a faster convergence rate is impossible when one is only given $y$.
We can also show that under our general asymptotic scaling $p=o(n)$, the method of moments estimator for $\theta$ is $\sqrt{p}$-consistent, which illustrates that the slower rate in part (c) is sub-optimal. For illustration, assume a simple setting with $k=1$ and suppose that the mean-mapping $M(\theta):=\operatorname{\mathbb{E}}B_{\theta}$ is bijective. Then, by setting $\hat{\beta}:=(X^{\top}X)^{-1}X^{\top}y$ as the usual least squares estimator, we can define the first moment-based estimator as:
$\hat{\theta}^{\mathsf{MoM}}_{p}:=M^{-1}\Big(\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\hat{\beta}_{i}\Big).$
By summing up the separate limiting distributions of $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\hat{\beta}_{i}-\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}{\beta}^{\star}_{i}\mid y,X$, $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}{\beta}^{\star}_{i}-M(\theta_{0})\mid X$ and applying the Delta method, we can show the following limit:
$\sqrt{p}(\hat{\theta}^{\mathsf{MoM}}_{p}-\theta_{0})\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(0,\frac{\sigma^{2}+\operatorname{\textnormal{Var}}(B_{\theta_{0}})}{M^{\prime}(\theta_{0})^{2}}\Big).$
While this limit generally holds under $p=o(n)$, the limiting variance is sub-optimal and strictly greater than $V(\theta_{0})^{-1}$.
Remark 3.4 (Faster rate of convergence at singularity).
In part (c) of 3.3, note that the $k$-dimensional vector $\kappa(\theta_{0})$ may be zero. While this singularity is typically a measure-zero event (see Figure 1 below for a visual illustration), the convergence rate at such points is faster than the $\frac{n}{p}$-rate in part (c). In particular, assuming $\kappa(\theta_{0})=0$, we can extend the proof of 3.3 to show that $\hat{\theta}^{\mathsf{vEB}}-\theta_{0}=O_{P}\Big(\frac{1}{\sqrt{p}}+\big(\frac{p}{n}\big)^{3/2}\Big)$. Consequently, (15) extends to the higher-dimensional asymptotic regime of $p=o(n^{3/4})$, and the phase transition no longer happens at $p\asymp n^{2/3}$.
Figure 1: Visualization of $\kappa(\theta)$ under various priors: (first) $\textsf{Ber}(\theta)$, (second) Normal with unknown variance: $N(0,1/\theta)$, (third/fourth) Gaussian mixtures $\frac{1}{2}N(\theta_{1},1)+\frac{1}{2}N(\theta_{2},0.25)$ with $\theta_{2}=-1$ fixed throughout the visualization. For the mixture prior, $\kappa(\theta)=(\kappa_{1}(\theta),\kappa_{2}(\theta))$ is a two-dimensional vector, and each component is plotted in the third/fourth figure. We see that the singularity $\kappa(\theta)=0$ is atypical. In particular, for the Gaussian mixture case, the singularity occurs when the two mixture components share a common mean $\theta_{1}=\theta_{2}=-1$.
To illustrate the limits in 3.3, we consider a toy example of the Gaussian prior with unknown mean. Here, the conjugacy allows directly computing the limit distribution of $\hat{\theta}^{\mathsf{vEB}}$.
Example 3.1 (Gaussian mean).
As a toy example, suppose that the prior is a Gaussian with unknown mean and known variance: $\mu_{\theta}\stackrel{{\scriptstyle d}}{{\equiv}}N(\theta,1)$. Then, the integral for defining the functions $F_{i}$ becomes tractable (see (11)), and we can write
$F_{i}(\theta)=\frac{1}{2}\log\Big(\frac{2\pi}{d_{i}+1}\Big)-\frac{\theta^{2}}{2}+\frac{(w_{i}+\theta)^{2}}{2(d_{i}+1)}.$
As this is a quadratic function in $\theta$, the vEB estimator in (11) can be explicitly written as
$\hat{\theta}^{\mathsf{vEB}}=\frac{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}w_{i}/(d_{i}+1)}{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}d_{i}/(d_{i}+1)}=\frac{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}w_{i}}{pd_{0}}+O_{P}\Big(\frac{1}{\sqrt{n}}\Big),$
where the second equality used $d_{i}=d_{0}+O_{P,X}\Big(\frac{1}{\sqrt{n}}\Big).$ Using the Gaussianity $w\mid X,\beta^{\star}\stackrel{{\scriptstyle d}}{{\equiv}}N(\Sigma\beta^{\star},\Sigma)$ where $\Sigma:=d_{0}X^{\top}X$ (where $d_{0}=\sigma^{-2}$ is as in 3.1), we can write
$\hat{\theta}^{\mathsf{vEB}}\mid X,\beta^{\star}\stackrel{{\scriptstyle d}}{{\equiv}}\frac{1}{pd_{0}}N\Big({1_{p}^{\top}\Sigma\beta^{\star}},1_{p}^{\top}\Sigma 1_{p}\Big).$
Then, using the Gaussianity of $\beta^{\star}\stackrel{{\scriptstyle d}}{{\equiv}}N(\theta_{0}1_{p},I_{p})$, we can spell out the limit of the mean as
$1_{p}^{\top}\Sigma\beta^{\star}\mid X\stackrel{{\scriptstyle d}}{{\equiv}}N\Big(\theta_{0}{1_{p}^{\top}\Sigma 1_{p}},{1_{p}^{\top}\Sigma^{2}1_{p}}\Big).$
By combining the two displays above and noting $\frac{1_{p}^{\top}\Sigma 1_{p}}{p}=d_{0}+O_{P,X}\Big(\frac{1}{\sqrt{n}}\Big)$ (standard mean and variance computations using the design matrix randomness), we have obtained the limit distribution of $\hat{\theta}^{\mathsf{vEB}}$:
$\hat{\theta}^{\mathsf{vEB}}\mid X\stackrel{{\scriptstyle d}}{{\equiv}}N\Big(\theta_{0},\frac{1+d_{0}}{pd_{0}}\Big)+O_{P}\Big(\frac{1}{\sqrt{n}}\Big).$
Under the asymptotic scaling $p=o(n)$, the error terms are negligible and we can write:
$\sqrt{p}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,1+\sigma^{2}).$ (18)
To see that this matches 3.3, note that the corresponding log-likelihood derivative can be written as $\ell^{\prime}(\theta;b)=b-\theta$ and $\operatorname{\textnormal{Var}}(B_{\theta}\mid W_{\theta})=\frac{1}{1+1/\sigma^{2}}$. We first compare with part (a) by computing the limiting variance under 3.3. Direct computations give
$I(\theta)=\operatorname{\textnormal{Var}}(B_{\theta})=1,\quad\operatorname{\mathbb{E}}\operatorname{\textnormal{Var}}(B_{\theta}\mid W_{\theta})=\frac{1}{1+1/\sigma^{2}},$
so we have $V(\theta)=\frac{1}{1+\sigma^{2}}$. This is exactly inverse of the limiting variance computed in (18).
To check consistency with parts (b) and (c) in 3.3, note that $h_{\theta}(t)=\operatorname{\mathbb{E}}\ell^{\prime}(\theta;B_{t,d_{0},\theta})=\operatorname{\mathbb{E}}[B_{t,d_{0},\theta}]-\theta=t$ is exactly linear. Thus, $h_{\theta}^{\prime\prime}\equiv 0$ for all $\theta$, and $\kappa\equiv 0$. In other words, every $\theta$ is a singular point in the sense of Remark 3.4. The limit in part (c) becomes degenerate, and does not reflect the tight order as the actual limit in (18).
3.2 Extension to misspecified priors
One natural question would be to go beyond the well-specified requirement in 3.3 by allowing $\beta^{\star}$ to be generated from a misspecified prior, say $\mu^{\star}\notin\{\mu_{\theta}\}_{\theta\in\Theta}$. Standard finite-dimensional Bayesian asymptotics [6, 29] consider a misspecified likelihood and showed that the resulting posterior contracts around a KL minimizer. Compared to such results, we consider a misspecified prior while still assuming a correctly specified likelihood (see (1)). Interestingly, as we shall illustrate, our results more resemble model misspecification in the frequentist literature [61].
To formalize the setup, instead of requiring a well-specified class of priors as in 3.3, assume that $\beta_{i}^{\star}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu^{\star}$. Here, $\mu^{\star}$ is a true “data-generating” measure, which may not be parametrized as $\mu_{\theta}$. We also modify the notations in parts (b) and (c) of 3.1 accordingly as follows.
Definition 3.2 (Extension of 3.1 to misspecified priors).
  1. (a)
    For any $\theta\in\Theta$, define random variables $(\beta^{\star},W^{\star},B_{\theta}^{\star})$ as
    $\beta^{\star}\sim\mu^{\star},\quad W^{\star}\mid\beta^{\star}\sim N(d_{0}\beta^{\star},d_{0}),\quad B_{\theta}^{\star}\mid W^{\star}\sim B_{W^{\star},d_{0},\theta}.$
    Here, recall the notation $B_{t,d,\theta}$ from part (c) in 3.1. Set $V(\theta)$ as in (13), and also define
    $V^{\star}(\theta):=V(\theta)-\operatorname{\mathbb{E}}\Big[\nabla^{2}\ell(\theta;B_{\theta}^{\star})+\big(\nabla\ell(\theta;B_{\theta}^{\star})\big)\big(\nabla\ell(\theta;B_{\theta}^{\star})\big)^{\top}\Big].$
  2. (ii)
    Define a mapping $\kappa^{\star}:\Theta\to\mathbb{R}^{k}$ as
    $\kappa^{\star}(\theta):=\frac{1}{2}\operatorname{\mathbb{E}}[(\beta^{\star})^{2}]\operatorname{\mathbb{E}}[h_{\theta}^{\prime\prime}(W^{\star})].$
Recalling the discussion in (14), $\beta^{\star}$ denotes a sample coefficient, and $W^{\star}$ denotes the approximation of each $w_{i}$. Also recalling from 3.1 that $\frac{dB_{W^{\star},d_{0},\theta}}{dB_{\theta}}(b)\propto e^{-\frac{(W^{\star}-d_{0}b)^{2}}{2d_{0}}}$, $B_{\theta}^{\star}$ denotes a draw from the posterior $\beta\mid W^{\star}$ (while using $\mu_{\theta}$ as the prior). Note that these modified notations reduce to that in 3.1 under correct specification, in the sense that we have $B_{\theta}^{\star}\stackrel{{\scriptstyle d}}{{\equiv}}B_{\theta}$ and $V^{\star}(\theta)=V(\theta)$ (by Bartlett’s second identity) when $\mu^{\star}=\mu_{\theta}$. However such reductions are not possible for a general $\mu^{\star}\notin\{\mu_{\theta}\}_{\theta\in\Theta}$.
Finally, define $\theta^{\star}$ as the parameter value that minimizes the KL divergence between $W^{\star}$ and $W_{\theta}$, i.e.
$\theta^{\star}\in\operatorname*{arg\,min}_{\theta\in\Theta}\mathrm{D_{KL}}(W^{\star}\mid W_{\theta}).$ (19)
Here, viewing $W^{\star}$ as an approximation of each $w_{i}$, $\theta^{\star}$ is the parameter value that best approximates this “data-generating” distribution $W^{\star}$. We formally state limit distributions of $\hat{\theta}^{\mathsf{vEB}}$ under misspecified priors below.
Corollary 3.4 (Generalization of 3.3 to misspecified priors).
Suppose Assumptions 3.1, 3.2 hold (where we replace all instances of $\theta_{0}$ to $\theta^{\star}$ in 3.2), and suppose that (19) has a unique minimizer. Assume R4,R6 for the class of priors $\mu_{\theta}$, and assume R5 for the random variable $B_{\theta^{\star}}^{\star}$. Then, for general $p=o(n)$, $\hat{\theta}^{\mathsf{vEB}}-\theta^{\star}=O_{P}\Big(\frac{1}{\sqrt{p}}+\frac{p}{n}\Big)$. In particular, the following limits hold.
  1. (a)
    When $p\ll n^{2/3}$, we have
    $\sqrt{p}(\hat{\theta}^{\mathsf{vEB}}-\theta^{\star})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(0,V^{\star}(\theta^{\star})^{-1}V(\theta^{\star})V^{\star}(\theta^{\star})^{-1}\Big).$
  2. (b)
    When $\frac{p^{3/2}}{n}\to\delta\in(0,\infty)$, we have
    $\sqrt{p}(\hat{\theta}^{\mathsf{vEB}}-\theta^{\star})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(\delta V^{\star}(\theta^{\star})^{-1}\kappa^{\star}(\theta^{\star}),V^{\star}(\theta^{\star})^{-1}V(\theta^{\star})V^{\star}(\theta^{\star})^{-1}\Big).$
  3. (c)
    When $n^{2/3}\ll p\ll n$, we have
    $\frac{n}{p}(\hat{\theta}^{\mathsf{vEB}}-\theta^{\star})\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}V^{\star}(\theta^{\star})^{-1}\kappa^{\star}(\theta^{\star}).$
As a sanity check, we recover the conclusion of 3.3 by taking $\mu=\mu^{\star}$ in 3.4. Here, note that the normal limit in part (a) is analogous to the limit theory for the frequentist MLE under likelihood misspecification (e.g. see Theorem 3.2 in [61]).
3.3 Lower bound
We show that the limiting variance in (15) is optimal by deriving a matching lower bound. For this goal, define the Fisher information for the observed responses $y$. While $y$ is $n$-dimensional, the Fisher information matrix $\mathcal{I}_{p}(\theta)$ is still indexed by $p$ for notational consistency.
Definition 3.3 (Fisher information).
Let $\mathcal{I}_{p}(\theta)$ denote the $k\times k$ Fisher information computed from the marginal likelihood $m_{\theta}$ (see (2)):
$\mathcal{I}_{p}(\theta):=-\operatorname{\mathbb{E}}\Big[\frac{\partial^{2}\log m_{\theta}(y)}{\partial\theta^{2}}\mid X\Big].$
Here, the information is conditioned on a realized design matrix (as the above expectation is taken with respect to $y\mid X$), and thus implicitly depends on $X$. Also, note the distinction from the notation $I(\theta)$ in 3.1, which denotes the Fisher information for a single sample from the prior $\mu_{\theta}$.
By the multivariate Cramér–Rao bound [e.g. see Theorem 2.6.1 in 35], for any unbiased estimator $\hat{\theta}_{\mathnormal{p}}$, its variance is bounded below by the inverse Fisher information:
$\operatorname{\mathbb{E}}\left[(\hat{\theta}_{p}-\theta_{0})(\hat{\theta}_{p}-\theta_{0})^{\top}\mid X\right]\succeq{\mathcal{I}_{p}(\theta_{0})}^{-1},\quad\forall\theta_{0}\in\Theta.$
Thus, understanding the fundamental limits of estimation boils down to evaluating the Fisher information $\mathcal{I}_{p}(\theta_{0}).$ Below we state our main lower bound result, under a technical bounded support assumption. [^1] This assumption follows as we use the moment bounds developed in [34]; we believe this can be relaxed to assuming $\operatorname{\textnormal{Var}}[B_{t,d,\theta}]\leq K$ for some uniform constant $K$, which also allows log-concave priors.
Assumption 3.5 (Bounded support).
Assume that the common support of the prior $\mu_{\theta}$ is bounded and that the log-likelihood $\ell(\theta;b)$ is continuous with respect to both $\theta,b$. Note that this assumption is one sufficient condition for 3.4.
Theorem 3.5.
Suppose Assumptions 3.1, 3.2, 3.3, 3.5 hold. Under the asymptotic scaling $p\ll n^{2/3}$, we have $\frac{\mathcal{I}_{p}(\theta_{0})}{p}\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}V(\theta_{0}).$
As the scaled limiting Fisher information matches the limiting variance in (15), we can conclude that the vEB estimator $\hat{\theta}^{\mathsf{vEB}}$ attains the optimal limiting variance when $p\ll n^{2/3}$. The proof of 3.5 builds upon the crucial observation that the Fisher information $\mathcal{I}_{p}(\theta_{0})$ can be expressed as moments of Random Field Ising Models (RFIMs, see A.1). Utilizing recent developments for RFIMs in [34], we establish moment bounds for the generalized summand $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(B_{i})$ (see Lemma A.8) for some smooth function $g(\cdot)$, which may be of independent interest.
4 Consequences of EB in downstream posterior inference
Moving on from the frequentist guarantees for recovering the prior $\mu_{\theta_{0}}$ discussed in Section 3, now we investigate consequences of EB (Empirical Bayes) in downstream posterior inference for the regression coefficients $\beta=(\beta_{1},\ldots,\beta_{p})$. To set notation, we formally define the oracle posterior and the EB posterior.
Definition 4.1 (Oracle and EB posterior).
Given a prior $\mu_{\theta}$, let $\operatorname{\mathbb{P}}_{\theta}(\beta)$ denote the posterior distribution under the likelihood in (1):
$\operatorname{\mathbb{P}}_{\theta}(\beta)=\operatorname{\mathbb{P}}_{\theta}(\beta\mid y,X)\propto\operatorname{\mathbb{P}}(y\mid X,\beta)\prod_{i=1}^{p}d\mu_{\theta}(\beta_{i}).$
In particular, we say that $\operatorname{\mathbb{P}}_{\theta_{0}}(\beta)$ is the oracle posterior. For any estimator $\hat{\theta}_{\mathnormal{p}}=\hat{\theta}_{\mathnormal{p}}(y,X)$ of $\theta_{0}$, we say that $\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}(\beta)$ is the EB posterior.
Here note that the EB posterior $\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}$ depends on the estimated prior $\hat{\theta}_{\mathnormal{p}}$. So, the data $y$ is double-dipped to construct $\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}$; once to write the likelihood $\operatorname{\mathbb{P}}(y\mid X,\beta)$ and second to estimate the prior $\hat{\theta}_{\mathnormal{p}}$. In this section, the notation $\hat{\theta}_{\mathnormal{p}}$ will denote any generic estimator, and $\hat{\theta}^{\mathsf{vEB}}$ will continue to denote the vEB estimator (11). Also, $\beta$ will denote a sample from the posterior (or its variational approximation), as opposed to the true regression coefficients $\beta^{\star}$ (see 3.3).
Our main goal is to understand whether the EB posterior is a good approximation of the oracle posterior. In this section, we answer this question in terms of posterior inference for (a) each coordinate, and (b) an one-dimensional projection. Due to technical reasons (mainly to use dependent concentration inequalities in [34]), this section will work under 3.5 and the asymptotic scaling $p=o(n^{2/3})$.
4.1 Estimating marginal distributions
Here, note that even understanding the oracle posterior itself is a nontrivial task due to its dependent and high-dimensional structure. In fact, it is extremely challenging to even compute the posterior mean and variance due to the intractable normalizing constant. Fortunately, the variational approximation turns out to be useful for understanding these intractable quantities. Below, we introduce the notion of variational posteriors, which are product measure approximations of the corresponding posterior.
Definition 4.2 (Variational posteriors).
Recall from 4.1 that $\operatorname{\mathbb{P}}_{\theta}$ denotes the usual posterior given a prior $\mu_{\theta}$. Define $\mathbb{Q}_{\theta}$ as the Mean-Field variational approximation of $\operatorname{\mathbb{P}}_{\theta}($:
$\mathbb{Q}_{\theta}(\beta):=\operatorname*{arg\,min}_{\mathbb{Q}=\prod_{i=1}^{p}Q_{i}}\mathrm{D_{KL}}(\mathbb{Q}\mid\operatorname{\mathbb{P}}_{\theta}),$
where each $Q_{i}$ has identical support as $\mu_{\theta}$. Note that $\mathbb{Q}_{\theta}$ is uniquely determined when the design matrix satisfies 3.1, as discussed in 2.2. In particular, we say that $\mathbb{Q}_{\theta_{0}}$ is the variational oracle posterior, and $\mathbb{Q}_{\hat{\theta}_{\mathnormal{p}}}$ is the variational EB posterior (given an estimator $\hat{\theta}_{\mathnormal{p}}$ [^2] Here, $\hat{\theta}_{\mathnormal{p}}$ is not necessarily the vEB estimator $\hat{\theta}^{\mathsf{vEB}}$. Note that $\mathbb{Q}_{\hat{\theta}_{\mathnormal{p}}}$ should be understood as the “variational approximation” of the EB posterior defined in 4.1, and does not have to use the vEB estimator.). Under the variational oracle posterior, define the coordinate-wise mean and variance as
$u_{i}:=\operatorname{\mathbb{E}}_{\beta\sim\mathbb{Q}_{\theta_{0}}}[\beta_{i}],\quad\tau_{i}^{2}:=\operatorname{\textnormal{Var}}_{\beta\sim\mathbb{Q}_{\theta_{0}}}[\beta_{i}].$
Similarly, define analogous quantities under the variational EB posterior as
$\hat{u}_{i}:=\operatorname{\mathbb{E}}_{\beta\sim\mathbb{Q}_{\hat{\theta}_{\mathnormal{p}}}}[\beta_{i}],\quad\hat{\tau}_{i}^{2}:=\operatorname{\textnormal{Var}}_{\beta\sim\mathbb{Q}_{\hat{\theta}_{\mathnormal{p}}}}[\beta_{i}].$
Here, note that $u_{i},\tau_{i}^{2}$ are oracle quantities that are functions of $\theta_{0}$, whereas $\hat{u}_{i},\hat{\tau}_{i}^{2}$ depend on the estimated hyperparameter $\hat{\theta}_{\mathnormal{p}}$. Also, note that these quantities are easily computable [see e.g. 34, 33], as opposed to the exact posterior mean and variance. In fact, these quantities connect to the Mean-Field optimizers from 2.2 as $u_{i}=u_{i,\theta_{0}},\hat{u}_{i}=u_{i,\hat{\theta}_{\mathnormal{p}}}$ (we will use the new notations for brevity).
The following proposition establishes high probability bounds between these quantities.
Theorem 4.1.
Suppose $p\ll n^{2/3}$ and assume Assumptions 3.1, 3.5. For each $i\leq p$, the following conclusions hold.
  1. (a)
    For both the oracle posterior $\operatorname{\mathbb{P}}_{\theta_{0}}$ and EB posterior $\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}$, their variational approximation is accurate for each coordinate:
    $\mathrm{D_{TV}}\big(\operatorname{\mathbb{P}}_{\theta_{0}}(\beta_{i}),\mathbb{Q}_{\theta_{0}}(\beta_{i})\big)=O_{P,X}\Big(\sqrt{\frac{p}{n}}\Big),\quad\mathrm{D_{TV}}\big(\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}(\beta_{i}),\mathbb{Q}_{\hat{\theta}_{\mathnormal{p}}}(\beta_{i})\big)=O_{P,X}\Big(\sqrt{\frac{p}{n}}\Big),\quad\forall i\leq p.$
    In particular, the coordinate-wise mean and variance are consistently estimated:
    $\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}_{\theta_{0}}}[\beta_{i}]$ $=u_{i}+O_{P,X}\Big(\sqrt{\frac{p}{n}}\Big),\quad\operatorname{\textnormal{Var}}_{\operatorname{\mathbb{P}}_{\theta_{0}}}[\beta_{i}]=\tau_{i}^{2}+O_{P,X}\Big(\sqrt{\frac{p}{n}}\Big),$ (20)
    $\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}}[\beta_{i}]$ $=\hat{u}_{i}+O_{P,X}\Big(\sqrt{\frac{p}{n}}\Big),\quad\operatorname{\textnormal{Var}}_{\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}}[\beta_{i}]=\hat{\tau}_{i}^{2}+O_{P,X}\Big(\sqrt{\frac{p}{n}}\Big).$ (21)
  2. (b)
    The vEB posterior $\mathbb{Q}_{\hat{\theta}_{\mathnormal{p}}}$ is close to the oracle posterior $\operatorname{\mathbb{P}}_{\theta_{0}}$ for each coordinate:
    $\mathrm{D_{TV}}\big(\operatorname{\mathbb{P}}_{\theta_{0}}(\beta_{i}),\mathbb{Q}_{\hat{\theta}_{\mathnormal{p}}}(\beta_{i})\big)=O_{P,X}\Big(\sqrt{\frac{p}{n}}+\big(\frac{p}{\sqrt{n}}+1\big)\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|\Big),\quad\forall i\leq p.$
    In particular, the coordinate-wise mean and variance are consistently estimated:
    $\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}_{\theta_{0}}}[\beta_{i}]$ $=\hat{u}_{i}+O_{P}\Big(\sqrt{\frac{p}{n}}+\big(\frac{p}{\sqrt{n}}+1\big)\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|\Big),$ (22)
    $\operatorname{\textnormal{Var}}_{\operatorname{\mathbb{P}}_{\theta_{0}}}[\beta_{i}]$ $=\hat{\tau}_{i}^{2}+O_{P}\Big(\sqrt{\frac{p}{n}}+\big(\frac{p}{\sqrt{n}}+1\big)\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|\Big).$
Here, part (a) illustrates that the variational approximations give meaningful estimates of the coordinate-wise mean and variance, for both the oracle posterior and EB posterior. Note that this claim justifies the variational approximation of the posterior, but does not discuss the effect of estimating the prior. Part (b) addresses this, by providing error bounds for moments of the vEB posterior and corresponding oracle quantities. For illustration, take $\hat{\theta}_{\mathnormal{p}}=\hat{\theta}^{\mathsf{vEB}}$. Then, still under $p\ll n^{2/3}$, part (a) of 3.3 gives $\|\hat{\theta}^{\mathsf{vEB}}-\theta_{0}\|=O_{P}\Big(\frac{1}{\sqrt{p}}\Big),$ and (22) simplifies to:
$\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}_{\theta_{0}}}[\beta_{i}]=\hat{u}_{i}+O_{P}\Big(\sqrt{\frac{p}{n}}+\frac{1}{\sqrt{p}}\Big),\quad\operatorname{\textnormal{Var}}_{\operatorname{\mathbb{P}}_{\theta_{0}}}[\beta_{i}]=\hat{\tau}_{i}^{2}+O_{P}\Big(\sqrt{\frac{p}{n}}+\frac{1}{\sqrt{p}}\Big).$
Thus, the vEB posterior constructed under $\hat{\theta}^{\mathsf{vEB}}$ consistently estimates the marginal mean and variance.
4.2 Uncertainty quantification for linear statistics
Now we move beyond pointwise estimation and discuss the consequences of vEB on uncertainty quantification. We first consider the oracle case. Due to our low SNR scaling where the posterior crucially depend on the choice of prior, one cannot establish a normal limit for each coordinate of the posterior [48]. Recently, jointly with Sumit Mukherjee, the authors established normal limiting distributions for one-dimensional projections of the posterior [33]. We restate part (a) of Theorem 3.1 and Proposition 3.2 in [33] below.
Proposition 4.2 (CLT under the oracle posterior).
Let $q\in\mathbb{R}^{p}$ be a deterministic vector normalized as $\|q\|=1$, satisfying $\|q\|_{\infty}=o(1)$. Assume that the prior $\mu_{\theta_{0}}$ is compactly supported. Set $\upsilon_{p}=\upsilon_{p}(y):=\operatorname{\textnormal{Var}}_{\mathbb{Q}_{\theta_{0}}}(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\beta_{i})=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}^{2}\tau_{i}^{2}$. Under the scaling $p\ll n^{2/3}$, we have
$\upsilon_{p}\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}\upsilon(\theta_{0}):=\operatorname{\mathbb{E}}\operatorname{\textnormal{Var}}(W_{\theta_{0}}\mid B_{\theta_{0}}),$ (23)
and the following limit holds under both the oracle and variational posteriors:
$\frac{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}(\beta_{i}-u_{i})}{\sqrt{\upsilon_{p}}}\mid y,X\xrightarrow[\operatorname{\mathbb{P}}_{\theta_{0}}]{d}N(0,1),\quad\frac{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}(\beta_{i}-u_{i})}{\sqrt{\upsilon_{p}}}\mid y,X\xrightarrow[\mathbb{Q}_{\theta_{0}}]{d}N(0,1).$
In other words, under the oracle posterior, a CLT with a normal limiting distribution holds for the statistic $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\beta_{i}$. Interestingly, the above result shows that one may replace the intractable mean and variance of the oracle posterior $\operatorname{\mathbb{P}}_{\theta_{0}}$ with the tractable counterparts that arise from the Mean-Field approximation $\mathbb{Q}_{\theta_{0}}$. This allows one to perform posterior inference, such as constructing credible intervals with explicit expressions. Note that establishing limit distributions of the oracle posterior requires the asymptotic requirement $p\ll n^{2/3}$.
Armed with this result, our goal is to perform inference for one-dimensional projections $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\beta_{i}$ when $\theta_{0}$ is unavailable. Given an EB estimator $\hat{\theta}_{\mathnormal{p}}$, two immediate questions are:
  • Does 4.2 generalize to the EB posterior?
  • Can one modify the centering and variance in 4.2 to data-driven quantities (for example the plug-in counterparts using $\hat{\theta}_{\mathnormal{p}}$)?
The following result answers both questions, where the first question is true (see part (b)) but the plug-in part of the second question is generally false (see part (c)). Recalling the notation $B_{t,d,\theta}$ and $W_{\theta_{0}}$ from 3.1, define the following Jacobian $J:\textnormal{int}(\Theta)\to\mathbb{R}^{k}$:
$J(\theta):=\operatorname{\mathbb{E}}\left[\frac{\partial\operatorname{\mathbb{E}}[B_{W_{\theta},d_{0},\theta}\mid W_{\theta}]}{\partial\theta}\right],\quad\forall\theta\in\textnormal{int}(\Theta).$ (24)
Theorem 4.3.
Let $q=(q_{1},\ldots,q_{p})\in\mathbb{R}^{p}$ denote a standardized deterministic vector with $\|q\|_{\infty}=o(1)$, and set $\hat{\upsilon}_{p}:=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}^{2}\hat{\tau}_{i}^{2}$. Suppose $p\ll n^{2/3}$ and assume $\hat{\theta}_{\mathnormal{p}}\operatorname*{\xrightarrow{\mathnormal{P}}}\theta_{0}$.
  1. (a)
    We have $\hat{\upsilon}_{p}\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}\upsilon(\theta_{0}),$ where $\upsilon(\theta_{0})$ is the value defined in (23).
  2. (b)
    Assuming $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}^{4}=o\Big(\frac{n}{p^{2}}\Big)$, the following holds under the EB posterior ${\operatorname{\mathbb{P}}}_{\hat{\theta}_{\mathnormal{p}}}$ (and also under the vEB posterior ${\mathbb{Q}}_{\hat{\theta}_{\mathnormal{p}}}$):
    $\frac{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}(\beta_{i}-\hat{u}_{i})}{\sqrt{\hat{\upsilon}_{p}}}\mid y,X\xrightarrow[\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}]{d}N(0,1).$ (25)
  3. (c)
    Additionally assume that $\sqrt{p}(\hat{\theta}_{\mathnormal{p}}-\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,S(\theta_{0}))$ for some continuous function $S:\Theta\to\mathbb{R}^{k\times k}$, and that $\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\to\gamma$ for some constant $1\leq\gamma\leq 1$. Then, when $\gamma=0$, (25) holds under the oracle posterior ${\operatorname{\mathbb{P}}}_{\theta_{0}}$. For general $\gamma$, the following holds:
    $\frac{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}(\beta_{i}-\hat{u}_{i})}{\sqrt{\hat{\upsilon}_{p}+\gamma^{2}J(\hat{\theta}_{\mathnormal{p}})^{\top}S(\hat{\theta}_{\mathnormal{p}})J(\hat{\theta}_{\mathnormal{p}})}}\mid X\xrightarrow[\operatorname{\mathbb{P}}_{\theta_{0}}]{d}N(0,1).$ (26)
    Here, note that the LHS is a sample-based quantity that only depends on $y,X$.
Remark 4.1.
Note that Cauchy-Schwartz gives $\frac{1}{\sqrt{p}}|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}|\leq\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}^{2}}=1$, which justifies the range $|\gamma|\leq 1$ in part (c). In particular, when $\gamma=0$, one may understand $q$ as an approximate contrast vector. We also note in passing that the centering $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\hat{u}_{i}$ in parts (b), (c) can be replaced by a more direct value of $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})$.
We elaborate on each conclusion. Part (a) extends the limit of the variational oracle posterior variance in (23) to the variance of the vEB posterior. Part (b) considers the EB posterior and shows that the oracle CLT in 4.2 can be extended by replacing $\theta_{0}$ with $\hat{\theta}_{\mathnormal{p}}$. Part (c) instead considers oracle posterior, and shows that the limit (25) extends when $\gamma=0$, or equivalently when $q$ is a contrast vector. However, for a general $\gamma\neq 0$, the limit no longer holds as the vEB posterior mean is biased. In other words, the estimation error of $\hat{\theta}_{\mathnormal{p}}$ results in a bias term with nontrivial fluctuations:
$\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}_{\theta_{0}}}[q^{\top}\beta]-\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}}[q^{\top}\beta]=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}u_{i}-\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\hat{u}_{i}+o_{P}(1)=\Theta_{P}(1).$
This is why we have to marginalize the data $y$ to get a standard normal limit in (26). The notation $J(\theta)$ in (24) also arises from this bias, as we use a first-order Taylor expansion to approximate
$\hat{u}_{i}-u_{i}=\operatorname{\mathbb{E}}_{\mathbb{Q}_{\hat{\theta}_{p}}}[\beta_{i}]-\operatorname{\mathbb{E}}_{\mathbb{Q}_{{\theta}_{0}}}[\beta_{i}]\approx(\hat{\theta}_{p}-\theta_{0})J(\hat{\theta}_{p}).$
The following corollary summarizes inferential consequences of 4.3 by using the vEB estimator $\hat{\theta}^{\mathsf{vEB}}$ to construct data-driven $100(1-\alpha)\%$ credible intervals with (conditional) coverage under the EB posterior, and (marginal) coverage under the oracle posterior. Here, for $\alpha\in(0,1)$, let $c_{\alpha/2}$ denote a constant such that $\operatorname{\mathbb{P}}(N(0,1)>c_{\alpha/2})=\alpha/2$.
Corollary 4.4.
  1. (a)
    The following interval
    $\mathcal{I}(y,X,\hat{\theta}^{\mathsf{vEB}})=\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\hat{u}_{i}\pm c_{\alpha/2}\sqrt{\hat{\upsilon}_{p}}\Big]$
    has asymptotic $1-\alpha$ coverage under the EB posterior $\beta\sim{\operatorname{\mathbb{P}}}_{\hat{\theta}^{\mathsf{vEB}}}$:
    $\operatorname{\mathbb{P}}_{\hat{\theta}^{\mathsf{vEB}}}\Big(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\beta_{i}\in{\mathcal{I}}(y,X,\hat{\theta}^{\mathsf{vEB}})\mid y,X\Big)\to 1-\alpha.$
    However, $\mathcal{I}(y,X,\hat{\theta}^{\mathsf{vEB}})$ has undercoverage under the oracle posterior $\operatorname{\mathbb{P}}_{\theta_{0}}$ if $\gamma\neq 0$.
  2. (b)
    The following interval with modified variance:
    $\tilde{\mathcal{I}}(y,X,\hat{\theta}^{\mathsf{vEB}})=\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\hat{u}_{i}\pm c_{\alpha/2}\sqrt{\hat{\upsilon}_{p}+\gamma^{2}J(\hat{\theta}^{\mathsf{vEB}})^{\top}V(\hat{\theta}^{\mathsf{vEB}})^{-1}J(\hat{\theta}^{\mathsf{vEB}})}\Big]$ (27)
    has asymptotic $1-\alpha$ average coverage under the oracle posterior $\beta\sim\operatorname{\mathbb{P}}_{\theta_{0}}$:
    $\operatorname{\mathbb{P}}_{\theta_{0}}\Big(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\beta_{i}\in\tilde{\mathcal{I}}(y,X,\hat{\theta}^{\mathsf{vEB}})\mid X\Big)\to 1-\alpha.$
Remark 4.2 (Comparison with high SNR posteriors).
As noted in the introduction, this is in sharp contrast with the “merging” properties of the EB posterior in high SNR settings where the EB posterior is a higher-order approximation of the oracle posterior [c.f. Prop 3.1, 45]. In other words, the cited result states that the EB credible region can serve as an oracle credible region, without the need for variance inflation as in part (b).
Remark 4.3 (Singularity).
Recall Remark 3.4, where we exhibit desirable properties under $\kappa(\theta_{0})=0$. A similar situation may happen when $J(\theta_{0})=0$, under which the CIs $\mathcal{I}$ and $\tilde{\mathcal{I}}$ in parts (a) and (b) asymptotically coincide. We will further discuss this in the simulations in Section 5.3.
5 Simulation studies
We conduct simulation studies to assess (i) the accuracy for recovering the prior hyperparameter $\theta_{0}$ and (ii) Bayesian uncertainty quantification for the regression coefficients $\beta$.
5.1 Setup
We vary the number of coefficients as $p=25,50,100,200$, and take $n=p^{2}$. All reported accuracy values that will follow are averaged over 400 replications, under a single realization of the random design $X$ with Gaussian entries (see 3.1). We consider the following priors, which allow both discrete, continuous, and mixed distributions:
  1. (a)
    $\textsf{Ber}(\pi)$ with $\pi_{0}=0.5$,
  2. (b)
    $\textsf{Spike-Slab}(\pi,\tau^{2})=\pi\delta_{0}+(1-\pi)N(0,\tau^{2})$ with $\pi_{0}=0.5,\tau_{0}=1$,
  3. (c)
    $\textsf{location-GMM}(\theta_{1},\theta_{2})=\frac{1}{2}N(\theta_{1},1)+\frac{1}{2}N(\theta_{2},0.25)$ with $\theta_{1,0}=-1,\theta_{2,0}=1$ [^3] Here, the MoM estimator suffers from a fundamental label switching issue, as the moment equations cannot identify $(\theta_{1},\theta_{2})=(-1,1)$ and (1,1)(1,-1), so we additionally imposed $\theta_{1}<\theta_{2}$. Interestingly, the same issue persists for the adaptive Langevin sampler, where (1,1)(1,-1) is a local minimizer of the free energy. Thus, we used a warm initialization for the Langevin sampler.,
The first is a discrete sparse prior, the second is a version of the spike and slab prior (where the prior parameters are treated as fixed), and the third is a non-sparse, bimodal Gaussian location mixture prior. Further details on implementation, computational time comparisons, and additional simulation results are provided in Appendix D.
5.2 Accuracy of prior recovery
We compare our vEB estimator $\hat{\theta}^{\mathsf{vEB}}$ with the following alternative EB methods:
  1. (a)
    Method of Moments (extension of Remark 3.3 by considering higher-order moments),
  2. (b)
    Adaptive Langevin Diffusion from [21, 22] for the continuous location-GMM prior. We did not implement this for other priors as the available implementation requires a Lebesgue density.
Prior Ber Spike-Slab location-GMM
$p$ \ Algo vEB MoM vEB MoM vEB MoM Langevin
25 0.042 0.048 0.173 0.391 0.339 0.318 0.713
50 0.023 0.023 0.125 0.373 0.168 0.237 0.400
100 0.011 0.012 0.086 0.335 0.065 0.203 0.182
200 0.006 0.006 0.059 0.299 0.028 0.187 0.089
Table 1: MSE $\|\hat{\theta}^{\mathsf{vEB}}-\theta_{0}\|^{2}$ for varying $p$ and $n=p^{2}$ (averaged over 400 replications). Note that the scaling varies across different parametric priors.
Table 1 reports the Mean Squared Error (MSE) for estimating the prior parameter $\theta_{0}$ under the aforementioned three prior families. The results validate that the vEB estimator is consistent in the sense that the MSE decreases as the dimension $p$ increases. In particular, for the Ber and location-GMM families, the MSE for the vEB estimator is roughly proportional to $1/p$, which aligns with the theoretical $\sqrt{p}$-convergence rate (here the Spike-Slab family is ruled out as the MSE aggregates two parameters which are of different scale). Compared to other estimators, the vEB estimator attains a smaller error, and empirically validates our asymptotic efficiency claim.
We also remark that the vEB estimator is computationally appealing. For estimating the location-GMM prior under $p=200$, the vEB estimator only took 0.05 seconds to compute, whereas the MoM and Langevin estimators took 1.6, 3.0 seconds, respectively [^4] Here, the vEB and MoM estimators are implemented in R whereas the Langevin estimator is implemented in Python. See Appendix D for details.. Complete results for computation time is reported in Table D.1 in the supplement, which illustrate that computing the vEB estimator took less than 0.1 seconds in all simulation settings and was always the fastest.
5.3 Accuracy of oracle posterior inference
Next, we consider posterior inference and evaluate the accuracy of the proposed vEB credible intervals (CIs) in part (b) of 4.4. We consider $90\%$ credible intervals for $q^{\top}\beta$, where $\beta$ is a draw from the oracle posterior $\operatorname{\mathbb{P}}_{\theta_{0}}$. We consider two choices for $q$:
  1. (a)
    $q^{\textsf{avg}}=\frac{1_{p}}{\sqrt{p}}$, which corresponds to the coordinate-wise average,
  2. (b)
    $q^{\textsf{con}}=\frac{1}{\sqrt{p}}(1_{p/2}^{\top},-1_{p/2}^{\top})^{\top}$, which is a contrast vector such that $\langle q^{\textsf{avg}},q^{\textsf{con}}\rangle=0$.
Note that each choice of $q$ yields the value $\gamma=\lim_{p\to\infty}\langle q,q^{\textsf{avg}}\rangle=1,0$, respectively. We compare the vEB credible interval $\tilde{\mathcal{I}}(y,X,\hat{\theta}^{\mathsf{vEB}})$ in (27) against its oracle analog that builds upon the oracle CLT in 4.2. While both theoretical coverage guarantees for the vEB and oracle CIs require bounded priors, we believe this is a technical requirement due to our proof technique and did not truncate the priors here. In addition to the coverage probability, we also report the average width of each credible interval.
Tables 2 and 3 respectively considers $q=q^{\textsf{avg}}$ and $q^{\textsf{con}}$. Looking at the oracle coverage guarantees, we see that the asymptotic CLT kicks in even for moderate sample sizes in the sense that it gives the correct coverage even when $p=25$. In terms of the vEB interval in Table 2, it is wider than the oracle interval for the Bernoulli and GMM prior. This is as expected by 4.4, since we take $\gamma=1$. Interestingly, the widths are comparable under the Spike-Slab prior, which is because $J(\hat{\theta}^{\mathsf{vEB}})\approx J(\theta_{0})=0$ and the width of the CI does not have to be increased for correct coverage (as mentioned in Remark 4.3). Similar to this, the widths of the EB-CI and oracle-CI in Table 3 are always comparable, as we have $\gamma=0$. In both tables, the vEB credible interval enjoys close to $90\%$ coverage and indicates accurate finite-sample performance. This suggests that the bounded prior requirement in our results in Section 4 may potentially be relaxed. Finally, we note that the vEB CIs’ coverage exhibits a larger standard error compared to the oracle CIs. This is more apparent in Table 2, as the vEB CIs only exhibit marginal coverage and no conditional coverage, theoretically.
Prior Ber Spike-Slab location-GMM
CI Oracle vEB Oracle vEB Oracle vEB
$p$ \ Object width cover width cover width cover width cover width cover width cover
25 1.47 0.902 (0.021) 3.12 0.860 (0.259) 1.87 0.899 (0.007) 1.72 0.783 (0.270) 2.50 0.897 (0.006) 3.32 0.899 (0.107)
50 1.47 0.901 (0.020) 3.20 0.873 (0.231) 1.87 0.900 (0.006) 1.77 0.819 (0.220) 2.52 0.898 (0.006) 3.29 0.907 (0.088)
100 1.47 0.900 (0.018) 3.22 0.890 (0.218) 1.87 0.900 (0.006) 1.82 0.862 (0.119) 2.52 0.900 (0.006) 3.26 0.894 (0.100)
200 1.47 0.901 (0.015) 3.24 0.907 (0.184) 1.87 0.900 (0.006) 1.87 0.885 (0.059) 2.52 0.900 (0.006) 3.25 0.901 (0.092)
Table 2: Average credible interval (CI) width and average coverage probability (standard error) under the oracle posterior, where $q^{\textsf{avg}}=\frac{1_{p}}{\sqrt{p}}$ is the coordinate-wise average.
Prior Ber Spike-Slab location-GMM
CI Oracle vEB Oracle vEB Oracle vEB
$p$ \ Object width cover width cover width cover width cover width cover width cover
25 1.47 0.899 (0.022) 1.27 0.822 (0.169) 1.87 0.892 (0.006) 1.72 0.774 (0.268) 2.50 0.890 (0.006) 2.47 0.874 (0.041)
50 1.47 0.900 (0.020) 1.39 0.876 (0.060) 1.87 0.899 (0.006) 1.77 0.815 (0.220) 2.52 0.899 (0.006) 2.46 0.884 (0.032)
100 1.47 0.900 (0.017) 1.44 0.891 (0.024) 1.87 0.899 (0.006) 1.82 0.860 (0.120) 2.52 0.900 (0.006) 2.49 0.893 (0.021)
200 1.47 0.900 (0.016) 1.45 0.898 (0.016) 1.87 0.899 (0.006) 1.87 0.885 (0.060) 2.52 0.900 (0.006) 2.50 0.893 (0.017)
Table 3: Average credible interval (CI) width and average coverage probability (standard error) under the oracle posterior, where $q^{\textsf{con}}=\frac{1}{\sqrt{p}}(1_{p/2}^{\top},-1_{p/2}^{\top})^{\top}$ is the contrast vector.
5.4 Improving the naive vEB estimator by debiasing
As the vEB estimator exhibits a sharp phase transition at $p\sim n^{2/3}$, a natural question arises: can one modify the vEB estimator to recover a $p^{-1/2}$ convergence rate beyond $p\sim n^{2/3}$? To this end, one natural strategy is debiasing, especially noting that the limiting distribution in part (c) of 3.3 is essentially a degenerate bias. The following proposition shows that debiasing the vEB estimator can allow the normal limit (15) to extend to higher dimensions.
Proposition 5.1.
Assume the usual assumptions 3.13.4, and consider the debiased estimator
$\tilde{\theta}^{\mathsf{d}}:=\hat{\theta}^{\mathsf{vEB}}-\frac{p}{n}V(\hat{\theta}^{\mathsf{vEB}})^{-1}\kappa(\hat{\theta}^{\mathsf{vEB}}).$
Then, the limit (15) holds for $\tilde{\theta}^{\mathsf{d}}$ as long as $p\ll n^{3/4}$.
The following Figure 2 considers the one-dimensional symmetric Gaussian mixture prior $\mu_{\theta}\equiv\frac{1}{2}N(\theta,1)+\frac{1}{2}N(-\theta,1)$, and illustrates that bias correction leads to better finite sample performance, which is more apparent for larger $p$. Our debiasing formula arises from correcting the second-order expansion of the likelihood. More specifically, noting that the slower rate of convergence of $\hat{\theta}^{\mathsf{vEB}}$ arises from the second term in the RHS of (17), we correct for this bias by replacing $\kappa(\theta_{0})$ with $\kappa(\hat{\theta}^{\mathsf{vEB}})$. This is in contrast to the more common first-order debiasing in the frequentist literature for estimating high-dimensional regression coefficients $\beta$ [55, 4, 5, 37]. We conjecture that it may be possible to extend the limit (15) to higher dimensions $p\gg n^{3/4}$ by additionally correcting for higher-order error terms in the likelihood expansion (17).
Figure 2: (Top row) Histogram of the estimators $\hat{\theta}^{\mathsf{vEB}}$ and $\tilde{\theta}^{\mathsf{d}}$, where the prior is the symmetric Gaussian mixture $\frac{1}{2}N(\theta,1)+\frac{1}{2}N(-\theta,1)$. We fix $n=1000$, and consider $p=\sqrt{n}$ and $n^{2/3}$ in each row. The solid red line shows the true value $\theta_{0}=1$ and the dotted blue line denotes the average across 400 replications (under a fixed realization of $X$). We see that $\hat{\theta}^{\mathsf{vEB}}$ has a positive bias and is right-skewed whereas the debiased estimator $\tilde{\theta}^{\mathsf{d}}$ has smaller bias. (Bottom row) Average bias and variance of $\hat{\theta}^{\mathsf{vEB}}$ and $\tilde{\theta}^{\mathsf{d}}$, where $p$ varies from $10$ to $250$ and $n=1000$ is fixed. As $p$ increases, the bias of $\hat{\theta}_{\mathnormal{p}}$ also increases.
6 Discussion
We discuss limitations of the current paper and avenues for future research.
  1. (a)
    Technical conditions.  It would be important to improve upon various technical conditions. One immediate direction involves our design matrix assumption (see 3.1), which requires entry-wise independence. While this is a natural starting point for theoretical understanding, it would be important to further relax this assumption to design matrices with less randomness, such as rotationally invariant matrices or fixed designs with correlated columns that were considered in the AMP literature [36, 18]. We believe that the definition of the vEB estimator (see (11)) as well as the limit distributions and debiasing formula should change appropriately. Additionally, some parts of our analysis require a bounded prior (see 3.5). As this restriction is inevitable due to limitations in existing literature for RFIMs [34], it would be important to further understand RFIMs with relaxed conditions on the base measure.
  2. (b)
    Singular or hierarchical priors.  One common choice for practical empirical Bayes procedures is to use finite (Gaussian) mixtures to model the prior $\mu_{\theta}$. While our results allow such mixtures, we require standard regularity assumptions (see 3.2), which rules out singular models such as overfitted mixtures. Indeed, it is well-known that overfitted mixtures have a slower $p^{1/4}$-rate of convergence for parameter estimation, even under the non-hierarchical setting with $p$ i.i.d. observations [23]. It would be interesting to explore whether our phase transition for the vEB estimator persists under such singular priors. On a related note, it would be interesting to view mixture priors as another layer of hierarchy and consider inference for each mixture component (for example, suppose $\beta_{i}\mid Z_{i}\sim P_{Z_{i}}$ and $Z_{i}\sim\textsf{Categorical}(\pi)$, where $P_{1},\ldots,P_{M}$ denotes each mixture distribution, $Z_{i}$ denotes the mixture component, and $\pi$ is a $M$-dimensional weight vector). Related hierarchical modeling approaches were recently studied under the classical normal means EB problem [26] as well as oracle structured variational inference in regression [50], and it would be interesting to explore additional hierarchical modeling under our setup.
  3. (c)
    Going beyond naive MF.  Finally, it would be interesting to utilize higher-order variational approximations (as opposed to the naive Mean-Field) such as the TAP free energy in statistical physics [54, 30] to estimate the prior. It is known that the TAP free energy can capture the leading-order asymptotics of the log-partition function under a proportional scaling [43], but its tightness compared to NMF has not been investigated when $p=o(n)$. It would be interesting whether using the TAP free energy can improve prior recovery, and whether it has connections with the debiasing strategy mentioned in 5.1.
Acknowledgments
The authors thank Sumit Mukherjee and Bodhisattva Sen for many helpful discussions throughout this work, and Zhou Fan for sharing the adaptive Langevin implementation in [21].
Appendix
The Appendix is organized as follows. We prove all main theorems in Appendix A. We prove main lemmas (including those in the main text) in Appendix B, and prove additional technical results in Appendix C. Appendix D discusses the simulation setups from the main paper in more detail and also presents additional numerical results.
Appendix A Proof of main theorems
A.1 Proof of 3.3 and 3.4
We begin with a modified version of Wald’s consistency result (see Theorem 5.14 in [56]), where we relax the independent sample requirement by directly assuming the pointwise LLNs that are used in the proof. This will enable us to establish the consistency of $\hat{\theta}^{\mathsf{vEB}}$.
Lemma A.1 (Wald’s consistency theorem).
Let $g:\mathbb{R}\times\Theta\to\mathbb{R}$ be a function such that the map $\theta\to g(v,\theta)$ is upper-semicontinuous for all $v$. Let $\{V_{i}\}_{i\geq 1}$ and $V_{\infty}$ be random variables such that satisfy the LLN
$\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(V_{i},\theta)\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}g(V_{\infty},\theta),\quad\forall\theta\,\in\Theta.$
Assume that the map $\theta\to\operatorname{\mathbb{E}}g(V_{\infty},\theta)$ is uniquely maximized at $\theta_{0}$, and that for every sufficiently small ball $U\in\Theta$ we have
$\operatorname{\mathbb{E}}\sup_{\theta\in U}g(V_{\infty},\theta)<\infty.$
Define $\mathcal{M}_{p}(\theta):=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(V_{i},\theta)$. Let $\hat{\theta}_{p}$ be any estimator such that $\mathcal{M}_{p}(\hat{\theta}_{\mathnormal{p}})\geq\mathcal{M}_{p}(\theta_{0})-o_{P}(1)$. Then, $\hat{\theta}_{\mathnormal{p}}\operatorname*{\xrightarrow{\mathnormal{P}}}\theta_{0}$.
To apply Lemma A.1, we require several additional lemmas whose proofs are deferred to Section B.2. Recall the notation $B_{\theta}$, $W_{\theta}$, and $B_{t,d,\theta}$ from 3.1. Also recall the definition of $w$ from (3). Our first Lemma will be used to prove that the limit of $p^{-1}\sum_{i=1}^{p}F_{i}(\theta)$ (see (11)) equals $F_{\mathrm{lim}}(\theta)$ where
$F_{\mathrm{lim}}(\theta):=\operatorname{\mathbb{E}}\Big[\log\Big(\int e^{W_{\theta_{0}}b-\frac{d_{0}b^{2}}{2}}d\mu_{\theta}(b)\Big)\Big].$ (28)
To wit, let $\operatorname{\mathsf{{Poly}}}(r)$ denote the set of all functions $h:\mathbb{R}\to\mathbb{R}$ with sub-polynomial growth $|h(z)|\lesssim 1+|z|^{r}$ for all $z\in\mathbb{R}$. Here $\lesssim$ hides constants free of $z$.
Lemma A.2 (LLN).
Suppose R5 from Assumption 3.4 holds.
  1. (a)
    Let $g:\mathbb{R}\to\mathbb{R}$ be a $\mathcal{C}^{1}$ function with $g^{\prime}\in\operatorname{\mathsf{{Poly}}}(r)$ for some $r\geq 0$ and $\operatorname{\mathbb{E}}g(W_{\theta_{0}})<\infty$. Then, the following LLN holds:
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(w_{i})\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}g(W_{\theta_{0}}).$
  2. (b)
    Additionally assuming R6, for any function $h\in\operatorname{\mathsf{{Poly}}}(r)$, we have
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}h(B_{w_{i},d_{i},\theta_{0}})\mid X$ $\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}h(B_{\theta_{0}}),\quad\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big[\operatorname{\mathbb{E}}h(B_{w_{i},d_{i},\theta_{0}})\big]^{2}\mid X$ $\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}\Big[\operatorname{\mathbb{E}}\big(h(B_{\theta_{0}})\mid W_{\theta_{0}}\big)\Big]^{2}.$
The next result shows how $F_{\mathrm{lim}}(\cdot)$ identifies $\theta_{0}$.
Lemma A.3.
Under the regularity condition R0 from 3.2, $F_{\mathrm{lim}}(\cdot)$ is uniquely maximized at $\theta=\theta_{0}$.
To go beyond consistency and establish limiting distributions, we use the following key lemma.
Lemma A.4 (CLT).
Suppose Assumptions 3.13.4 hold.
  1. (a)
    When $\frac{p^{3/2}}{n}\to\delta\in[0,\infty)$, we have
    $\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N\big(\delta\kappa(\theta_{0}),V(\theta_{0})\big),$
    where $\kappa(\theta_{0})$ is defined in 3.1. Here we note that if $p\ll n^{2/3}$, then $\delta=0$ and the mean of the limiting normal is 0.
  2. (b)
    When $p\gg n^{2/3}$, we have
    $\frac{n}{p^{2}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}\kappa(\theta_{0}).$
  3. (c)
    When $p\ll{n^{3/4}}$, we have
    $\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})-\frac{p^{3/2}}{n}\kappa(\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N\big(0,V(\theta_{0})\big).$
Finally, we state two technical moment-related lemmas using the sub-Gaussian assumption of the tilted prior (see R6 in 3.4). The following lemma provides convenient moment and cgf bounds under the sub-Gaussian assumption.
Lemma A.5.
Under R6 from 3.4, we have
$\sup_{\theta\in\Theta}\log\int e^{tb-\frac{db^{2}}{2}}d\mu_{\theta}(b)\lesssim 1+|t|^{4},\quad\operatorname{\mathbb{E}}|B_{t,d,\theta}|^{r}\lesssim_{r}1+|t|^{3r},\quad\forall r>0,~\theta\in\Theta,d>0.$
Additionally assuming R5 from 3.4, we have $\sup_{\theta\in\Theta}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}|B_{w_{i},d_{i},\theta}|^{r}=O_{P}(p)$.
The next lemma generalizes Lemma B.4 in [34], which required a bounded prior.
Lemma A.6.
Suppose R6 from 3.4 holds. For any $t\in\mathbb{R},\theta\in\Theta$ and $h\in\operatorname{\mathsf{{Poly}}}(r)$, we have
$\Big|\log\Big[\int e^{tb-\frac{db^{2}}{2}}d\mu_{\theta}(b)\Big]-\log\Big[\int e^{tb-\frac{d_{0}b^{2}}{2}}d\mu_{\theta}(b)\Big]\Big|$ $\lesssim|d-d_{0}|\big(1+|t|^{6}\big),$
$|\operatorname{\mathbb{E}}h(B_{t,d,\theta})-\operatorname{\mathbb{E}}h(B_{t,d_{0},\theta})|$ $\lesssim_{r}|d-d_{0}|\big(1+|t|^{3(r+2)}\big).$
Lemma A.7.
Suppose R5 from 3.4 holds. For $\mathcal{F}_{i}:=\sum_{j\neq i}A_{p}(i,j)\beta_{j}^{\star}$ and any fixed constant $r>0$, we have
$\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}e^{|\mathcal{F}_{i}|}\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}1,\quad\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}[{|w_{i}|^{r}}\mid X,\beta^{\star}]=O_{P}(1).$
Proof of 3.3.
Under the general asymptotic scaling $p=o(n)$, we establish common bounds before proving the individual conclusions in parts (a), (b), and (c). Throughout the proof, the design matrix $X$ will always be conditioned on unless specified otherwise.
  1. (i)
    Consistency: We begin by proving $\hat{\theta}^{\mathsf{vEB}}\operatorname*{\xrightarrow{\mathnormal{P}}}\theta_{0}$ by checking the conditions in Lemma A.1. To this extent, set $g(t,d,\theta):=\log\int e^{tb-\frac{db^{2}}{2}}d\mu_{\theta}(b)$ and recall that $\hat{\theta}^{\mathsf{vEB}}$ is a maximizer of $\tilde{M}_{p}(\theta)=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(w_{i},d_{i},\theta)$. Since
    $d_{i}=d_{0}+\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{1}{\sqrt{n}}\Big)$ (29)
    by [33, Equation (A.19)], define a simpler objective $M_{p}(\theta):=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(w_{i},d_{0},\theta)$ to which we apply Lemma A.1. First note that the derivative of the map $t\mapsto g(t,d_{0},\theta)$ belong to $\operatorname{\mathsf{{Poly}}}(r)$ for all $\theta\in\Theta$, by Lemma A.5. Therefore, the first two conditions in Lemma A.1, LLN and unique maximizer, directly follow from Lemmas A.2 and A.3 respectively. For the third condition, it suffices to show the following bound:
    $\operatorname{\mathbb{E}}\left[\sup_{\theta\in\Theta}\log\Big(\int e^{W_{\theta_{0}}b-\frac{d_{0}b^{2}}{2}}d\mu_{\theta}(b)\Big)\right]<\infty.$
    For this goal, applying the first bound in Lemma A.5 gives
    $\sup_{\theta\in\Theta}\log\int e^{wb-\frac{d_{0}b^{2}}{2}}d\mu_{\theta}(b)\lesssim 1+w^{4},$
    Hence, the expectation is upper bounded by $O(1+\operatorname{\mathbb{E}}W_{\theta_{0}}^{4})<\infty$. Finally, using Lemma A.6 coupled with (29), we can write $g(w_{i},d_{i},\theta)=g(w_{i},d_{0},\theta)+O_{P}\Big(\frac{1+|w_{i}|^{6}}{\sqrt{n}}\Big)$. By noting that $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|w_{i}|^{6}=O_{P}(p)$ (by Lemma A.7), we have
    $M_{p}(\hat{\theta}^{\mathsf{vEB}})=\tilde{M}_{p}(\hat{\theta}^{\mathsf{vEB}})+o_{P}(1)\geq\tilde{M}_{p}(\theta_{0})+o_{P}(1)=M_{p}(\theta_{0})+o_{P}(1).$
    Therefore consistency follows from Lemma A.1.
  2. (ii)
    Limiting distribution: For a $k\times k\times k$ tensor $T$ and a vector $v$, define $T\times_{a}v\in\mathbb{R}^{k\times k}$ as the mode-$a$ product of $T$ and $v$. By the first order conditions and Taylor approximation, we have
    $0_{k}$ $=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\hat{\theta}^{\mathsf{vEB}})$
    $=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})+\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}F_{i}(\theta_{0})(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})+\frac{1}{2p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{3}F_{i}(\tilde{\theta}_{n})\times_{2}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})\times_{3}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})$
    for some $\tilde{\theta}_{n}\in(\theta_{0},\hat{\theta}^{\mathsf{vEB}})$. Hence, we can write
    $(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})=-\Big(\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}F_{i}(\theta_{0})+\frac{1}{2p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{3}F_{i}(\tilde{\theta}_{n})\times_{3}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})\Big)^{-1}\Big(\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})\Big).$ (30)
    By direct computations, we can write the derivatives as:
    $\nabla F_{i}(\theta)$ $=\operatorname{\mathbb{E}}\nabla\ell(\theta;B_{w_{i},d_{i},\theta}),$
    $\nabla^{2}F_{i}(\theta)$ $=\operatorname{\mathbb{E}}[\nabla^{2}\ell(\theta;B_{w_{i},d_{i},\theta})]+\operatorname{\mathbb{E}}[\nabla\ell(\theta;B_{w_{i},d_{i},\theta})^{2}]-\big[\operatorname{\mathbb{E}}\nabla\ell(\theta;B_{w_{i},d_{i},\theta})\big]^{2},$
    $\nabla^{3}F_{i}(\theta)$ $=\operatorname{\mathbb{E}}[\nabla^{3}\ell(\theta;B_{w_{i},d_{i},\theta})+\nabla\ell(\theta;B_{w_{i},d_{i},\theta})^{3}]+2[\operatorname{\mathbb{E}}\nabla\ell(\theta;B_{w_{i},d_{i},\theta})]^{3}$
    $~~-3\operatorname{\mathbb{E}}[\nabla\ell(\theta;B_{w_{i},d_{i},\theta})]\otimes\operatorname{\mathbb{E}}[\nabla^{2}\ell(\theta;B_{w_{i},d_{i},\theta})+\nabla\ell(\theta;B_{w_{i},d_{i},\theta})^{2}].$
    In the above expression, for a vector $v\in\mathbb{R}^{k}$, we have abbreviated $v^{k}$ as the $k$-fold outer product, i.e. $v^{2}:=v\otimes v,v^{3}:=v\otimes v\otimes v$. Recall that $B_{w_{i},d_{i},\theta}$ is defined in 3.1, and note that all expectations are conditional on $w_{i}$s. We claim that the third derivative term satisfies
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{3}F_{i}(\tilde{\theta}_{n})\times_{3}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})=o_{P}(1),$
    and is negligible. To see this, use the growth-condition R4 from 3.4, to write:
    $\|\nabla^{3}F_{i}(\theta)\|_{F}\lesssim 1+\operatorname{\mathbb{E}}|B_{w_{i},d_{i},\theta}|^{r},$
    for some integer $r>0$, where the implied constant does not depend on $\theta$. Then, by the last bound in Lemma A.5, we have
    $\|\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{3}F_{i}(\tilde{\theta}_{n})\|_{F}\lesssim\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\sup_{\theta}\|\nabla^{3}F_{i}(\theta)\|_{F}\lesssim\sup_{\theta\in\Theta}\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(1+\operatorname{\mathbb{E}}|B_{w_{i},d_{i},\theta}|^{r})=O_{P}(1),$
    so the claim follows by Slutsky’s theorem.
    Regarding the second derivative term, we claim the limit
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}F_{i}(\theta_{0})\operatorname*{\xrightarrow{\mathnormal{P}}}-V(\theta_{0}).$
    Now, by applying the simplified LLNs in part (b) of Lemma A.2 with
    $h(b)=\nabla^{2}\ell(\theta_{0};b),[\nabla\ell(\theta_{0};b)]^{2},\nabla\ell(\theta_{0};b)$
    (corresponding to each line below), the following limits hold:
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}[\nabla^{2}\ell(\theta_{0};B_{w_{i},d_{i},\theta_{0}})]$ $\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}\nabla^{2}\ell(\theta_{0};B_{\theta_{0}}),$ (31)
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}[\nabla\ell(\theta_{0};B_{w_{i},d_{i},\theta_{0}})^{2}]$ $\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{\theta_{0}})^{2},$
    $-\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}[\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{w_{i},d_{i},\theta_{0}})]^{2}$ $\operatorname*{\xrightarrow{\mathnormal{P}}}-\operatorname{\mathbb{E}}\Big[\operatorname{\mathbb{E}}\big(\nabla\ell(\theta_{0};B_{\theta_{0}})\mid W_{\theta_{0}}\big)\Big]^{2}.$
    Therefore,
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}F_{i}(\theta_{0})$ $\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}\nabla^{2}\ell(\theta_{0};B_{\theta_{0}})+\sum_{\mathnormal{i}=1}^{\mathnormal{p}}[\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{w_{i},d_{i},\theta_{0}})]^{2}-\operatorname{\mathbb{E}}\Big[\operatorname{\mathbb{E}}\big(\nabla\ell(\theta_{0};B_{\theta_{0}})\mid W_{\theta_{0}}\big)\Big]^{2}$
    $=\operatorname{\mathbb{E}}[\nabla^{2}\tilde{\ell}(\theta_{0};W_{\theta_{0}}),$
    where $\tilde{\ell}(\theta_{0};\cdot)$ is the log-likelihood of the random variable $W_{\theta_{0}}$. The claim now follows via Lemma 3.1.
    By combining the above conclusions, we can simplify (30) as follows:
    $(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})=\big(V(\theta_{0})+o_{P}(1)\big)^{-1}\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0}).$ (32)
  3. (iii)
    Finally, the individual conclusions in the theorem follow by using the limiting distribution of $p^{-1}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})$ from Lemma A.4 alongside Slutsky’s theorem.
Next we prove 3.4 by making minor changes in the above argument.
Proof of 3.4.
We first modify Lemma A.3 to the following claim:
$F_{\mathrm{lim}}^{\star}(\theta):=\operatorname{\mathbb{E}}\Big[\log\Big(\int e^{W^{\star}b-\frac{d_{0}b^{2}}{2}}d\mu_{\theta}(b)\Big)\Big]$
is uniquely maximized at $\theta=\theta^{\star}$. To show this, note that
$\mathrm{D_{KL}}(W^{\star}\mid W_{\theta})=\operatorname{\mathbb{E}}\Big[\log\Big(\int e^{W^{\star}b-\frac{d_{0}b^{2}}{2}}d\mu^{\star}(b)\Big)\Big]-\operatorname{\mathbb{E}}\Big[\log\Big(\int e^{W^{\star}b-\frac{d_{0}b^{2}}{2}}d\mu_{\theta}(b)\Big)\Big].$
Thus, $\operatorname*{arg\,max}_{\theta}F_{\mathrm{lim}}^{\star}(\theta)=\operatorname*{arg\,min}_{\theta}\mathrm{D_{KL}}(W^{\star}\mid W_{\theta})=\theta^{\star}$, where we use (19) alongside the unique minimizer assumption. By re-writing the first order conditions for $\nabla F_{0}^{\star}(\theta^{\star})=0$, we get $\operatorname{\mathbb{E}}\nabla\ell(\theta^{\star};B_{\theta^{\star}}^{\star})=0$, which is the misspecified analog of Bartlett’s first identity.
Now, we can apply Wald’s consistency theorem similar to that in part (a) of proof of 3.3, and get $\hat{\theta}^{\mathsf{vEB}}\operatorname*{\xrightarrow{\mathnormal{P}}}\theta^{\star}$. To establish the limit of the numerator, we again use (30) where all instances of $\theta_{0}$ are replaced with $\theta^{\star}$:
$(\hat{\theta}^{\mathsf{vEB}}-\theta^{\star})=-\Big(\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}F_{i}(\theta^{\star})+o_{P}(1)\Big)^{-1}\Big(\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta^{\star})\Big).$ (33)
Noting that the proof of Lemma A.4 is still valid under the notation change from $(B_{\theta_{0}},W_{\theta_{0}})$ to $(B_{\theta^{\star}}^{\star},W^{\star})$, parts (a) and (b) can be modified to:
  • When $\frac{p^{3/2}}{n}\to\delta\in[0,\infty)$, we have
    $\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta^{\star})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N\big(\delta\kappa^{\star}(\theta^{\star}),V(\theta^{\star})\big).$
  • When $p\gg n^{2/3}$, we have
    $\frac{n}{p^{2}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta^{\star})\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}\kappa^{\star}(\theta^{\star}).$
Also, the limit of the denominator modifies to (see (31) and Lemma 3.1):
$\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}F_{i}(\theta^{\star})\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}\Big[\nabla^{2}\ell(\theta^{\star};B_{\theta}^{\star})+\big(\nabla\ell(\theta^{\star};B_{\theta}^{\star})\big)\big(\nabla\ell(\theta^{\star};B_{\theta}^{\star})\big)^{\top}\Big]-V(\theta^{\star})=-V^{\star}(\theta^{\star}).$
The only change here is that we cannot use Bartlett’s second identity to simplify the final equality in (42). Now, the individual conclusions in each part of 3.4 follow by combining the separate limits for the numerator and denominator of (33) via Slutsky’s theorem. ∎
A.2 Proof of 3.5
Without loss of generality, assume that the bounded support in 3.5 is the interval [1,1][-1,1]. 3.5 is based on two technical moment bounds for RFIMs. Before stating the lemmas, we introduce some additional notation.
Definition A.1 (Random field Ising model (see [34, 40, 1, 8])).
Given $\theta\in\Theta$ and any $c=(c_{1},\ldots,c_{p})\in\mathbb{R}^{p}$, let $B=(B_{1},\ldots,B_{p})\stackrel{{\scriptstyle d}}{{\equiv}}\tilde{\operatorname{\mathbb{P}}}=\tilde{\operatorname{\mathbb{P}}}^{A_{p},c}_{\theta}$ be a sample from the following RFIM:
$\frac{d\tilde{\operatorname{\mathbb{P}}}^{A_{p},c}_{\theta}}{d\prod_{i=1}^{p}\mu_{i,\theta}}(b):=\frac{\exp\Big(\frac{b^{\top}A_{p}b}{2}+c^{\top}b\Big)}{Z_{p}(c,\theta)},\quad\forall\,b\in[-1,1]^{p},$
where $Z_{p}(c,\theta)$ is the normalizing constant defined in (5) and $\mu_{i,\theta}$ are quadratic tilts in 1.1.
Recall from 1.1 that $\mu_{i,c_{i},\theta}$ denotes the exponential tilt of $\mu_{i,\theta}$ with $\frac{d\mu_{i,c_{i},\theta}}{d\mu_{i,\theta}}(B_{i})\propto\exp(c_{i}B_{i})$. Denote the $i$th marginal of the Mean-Field approximation (that attains the supremum in (6)) as
${Q}_{i,\theta,c}:=\mu_{i,s_{i,\theta}+c_{i},\theta},\quad\textnormal{where}\quad{s}_{i,\theta}:=\sum_{\mathnormal{j}=1}^{\mathnormal{p}}A_{p}(i,j){u}_{j,\theta},$
where $u_{j,\theta}$s are as in 2.2. For any function $f:[-1,1]\to\mathbb{R}^{k}$ and measure $Q_{i}$ on [1,1][-1,1], write the expectation of $f$ under $Q_{i}$ (assuming it is well-defined) as:
$\langle f\rangle_{Q_{i}}:=\operatorname{\mathbb{E}}^{B_{i}\sim Q_{i}}f(B_{i}).$
The following lemma provides finite-sample error bounds for the first and second moment of general RFIMs (with arbitrary random fields $c$).
Lemma A.8.
Let $A_{p}$ be the off-diagonal parts of the random design in 3.1, $c\in\mathbb{R}^{p}$ be an arbitrary vector, and consider the asymptotic regime $p=o(n^{2/3})$. For $B\sim\tilde{\operatorname{\mathbb{P}}}^{A_{p},c}_{\theta}$ and any vector-valued $\mathcal{C}^{2}$ function $g:[-1,1]\to[-1,1]^{k}$, the following limits hold:
  1. (a)
    $\frac{1}{p}\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},c}_{\theta}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big[g(B_{i})-\langle g\rangle_{{Q}_{i,\theta,c}}\big]\Big]$ $=\mathnormal{O}_{\mathnormal{P},X}\left({\frac{1}{\sqrt{n}}}+\frac{p}{n}\right),$
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[\langle g\rangle_{{Q}_{i,\theta,c}}-\langle g\rangle_{\mu_{i,c_{i},\theta}}\Big]$ $=O_{P,X}\Big(\sqrt{\frac{p}{n}}\Big).$
  2. (b)
    $\frac{1}{p}\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},c}_{\theta}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big[g(B_{i})-\langle g\rangle_{{Q}_{i,\theta,c}}\big]\Big]^{2}=$ $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[\langle g^{2}\rangle_{\mu_{i,c_{i},\theta}}-\big(\langle g\rangle_{\mu_{i,c_{i},\theta}}\big)^{2}\Big]+\mathnormal{O}_{\mathnormal{P},X}\left(\frac{1}{\sqrt{p}}+\sqrt{\frac{p^{3}}{n^{2}}}\right).$
Here, the hidden constants are absolute constants, and part (b) abbreviates $v^{2}:=v\otimes v$ for vectors $v\in\mathbb{R}^{k}$.
Proof of 3.5.
Recall that the marginal likelihood in (2). By direct computations noting that
$\tilde{\operatorname{\mathbb{P}}}_{\theta}^{A_{p},w}(B)\propto\exp\Big(\frac{B^{\top}A_{p}B}{2}+w^{\top}B\Big)\prod_{i=1}^{p}\mu_{i,\theta}(B_{i})\propto\exp\Big(-\frac{\|y-XB\|^{2}}{2\sigma^{2}}\Big)\prod_{i=1}^{p}\mu_{\theta}(B_{i})$
and $\mu_{\theta}(b)=e^{\ell(\theta;b)}$, we have
$\frac{\partial\log m_{\theta}}{\partial\theta}$ $=\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla\ell(\theta;B_{i})\Big],$
$\frac{\partial^{2}\log m_{\theta}}{\partial\theta^{2}}$ $=\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}\ell(\theta;B_{i})\Big]+\operatorname{\textnormal{Var}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla\ell(\theta;B_{i})\Big].$
Define $g^{(1)}(b):=\nabla\ell(\theta_{0};b)$, $g^{(2)}(b):=\nabla^{2}\ell(\theta_{0};b)$. Under 3.5, $\ell(\theta;b)$ is a continuous function defined on a compact set, and hence uniformly bounded. For $B\sim\tilde{\operatorname{\mathbb{P}}}_{\theta_{0}}^{A_{p},w},$ abbreviate the average of $g^{(a)}(B_{i})$ as
$\overline{g^{(a)}(B)}:=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g^{(a)}(B_{i}),\quad a=1,2.$
Then, we can write the Fisher information $\mathcal{I}_{p}(\theta_{0})$ as:
$\mathcal{I}_{p}(\theta_{0})=-\operatorname{\mathbb{E}}\left[\frac{\partial^{2}\log m_{\theta}}{\partial\theta^{2}}\mid_{\theta=\theta_{0}}\right]=-\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta_{0}}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}\ell(\theta_{0};B_{i})\Big]\right]-\operatorname{\mathbb{E}}\left[\operatorname{\textnormal{Var}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta_{0}}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla\ell(\theta_{0};B_{i})\Big]\right].$
We separately compute the limit of each term. To control the first term, we use part (a) of Lemma A.8 to write
$\frac{1}{p}\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta_{0}}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}\ell(\theta_{0};B_{i})\Big]$ $=\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta_{0}}}[\overline{g^{(2)}(B)}]$
$=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\langle g^{(2)}\rangle_{\mu_{i,w_{i},\theta_{0}}}+o_{P,X}(1)$
$\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}\Big[\operatorname{\mathbb{E}}\big[\nabla^{2}\ell(\theta_{0};B_{\theta_{0}})\mid W_{\theta_{0}}\big]\Big]=-I(\theta_{0}).$
The convergence in probability is conditional on the covariate matrix $X$ and it follows from limits in (31). By taking an outer expectation (over $y$), we get the final limit:
$\operatorname{\mathbb{E}}\left[\frac{1}{p}\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta_{0}}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}\ell(\theta_{0};B_{i})\Big]\right]\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}-I(\theta_{0}).$
Here we use that fact the inner term is bounded (by a uniform constant).
Next, we compute the limit of the second term of $\mathcal{I}_{p}(\theta_{0})$. Using both parts of Lemma A.8, we can write
$p\operatorname{\textnormal{Var}}^{\tilde{\operatorname{\mathbb{P}}}_{\theta_{0}}^{A_{p},w}}[\overline{g^{(1)}(B)}]$
$=$ $p\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}_{\theta_{0}}^{A_{p},w}}\Big[\overline{g^{(1)}(B)}-\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\langle g^{(1)}\rangle_{{Q}_{i,\theta_{0},w_{i}}}\Big]^{2}-\frac{1}{p}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\langle g^{(1)}\rangle_{{Q}_{i,\theta_{0},w}}-\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}_{\theta_{0}}^{A_{p},w}}\overline{g^{(1)}(B)}\Big]^{2}$
$=$ $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[\langle(g^{(1)})^{2}\rangle_{{\mu}_{i,w_{i},\theta_{0}}}-\big(\langle g^{(1)}\rangle_{{\mu}_{i,w_{i},\theta_{0}}}\big)^{2}\Big]+o_{P,X}(1)$
$\operatorname*{\xrightarrow{\mathnormal{P}}}$ $I(\theta)-V(\theta_{0}),$ (34)
where the convergence in probability in the last line is conditional on $X$. Here, the second line uses the basic identity $\operatorname{\textnormal{Var}}(V)=\operatorname{\mathbb{E}}[V-v]^{2}-[v-\operatorname{\mathbb{E}}V]^{2}$, where $V$ is a random vector and $v$ is a deterministic vector. The third line follows from Lemma A.8. The last limit follows from (31). Finally, taking an outer expectation (over $y$) noting that the LHS of (34) is bounded, we get
$\operatorname{\mathbb{E}}\left[\frac{1}{p}\operatorname{\textnormal{Var}}_{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta_{0}}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla\ell(\theta_{0};B_{i})\Big]\right]\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}I(\theta)-V(\theta_{0}).$
The proof is complete by combining the two limits.
A.3 Proofs from Section 4
We first state a stability result with respect to the prior parameter $\theta$.
Lemma A.9.
Suppose 3.5 holds. For any $t\in\mathbb{R},i\leq p$, $\theta,\theta^{\prime}\in\Theta$, and a bounded function $h:[-1,1]\to[-1,1]$, we have
$|\langle h\rangle_{\mu_{i,t,\theta}}-\langle h\rangle_{\mu_{i,t,\theta^{\prime}}}|\lesssim\|\theta-\theta^{\prime}\|,$
where the hidden constant only depends on the prior likelihood $\ell$. In particular, we have
$|\psi_{i,\theta}^{\prime}(t)-\psi_{i,\theta^{\prime}}^{\prime}(t)|,~|\psi_{i,\theta}^{\prime\prime}(t)-\psi_{i,\theta^{\prime}}^{\prime\prime}(t)|=O\big(\|\theta-\theta^{\prime}\|\big)$
The next technical result provides a way to quantify the distance between the true posterior and the EB posterior. As it involves the use of several technical lemmas from [34] to compute moment bounds for RFIMs, we state them in the following lemma. Each part follows from Theorem 2.3, Lemma 3.2(b), and Lemma 3.1 in [34], respectively. We also refer the reader to [7, Lemma 2,10] and [17, Lemma A.1] for related results.
Lemma A.10 (Facts about RFIMs).
For $B\sim\tilde{\operatorname{\mathbb{P}}}_{\theta}^{A_{p},c}$, let $m_{i}:=m_{i}(B)=\sum_{j\neq i}A_{p}(i,j)B_{j}$ denote the “local field” for each coordinate $i$. Then, under 3.1 and $p\ll n^{2/3}$, the following conclusions hold.
  1. (a)
    For a constant vector $\gamma\in\mathbb{R}^{p}$, we have $\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\gamma_{i}(B_{i}-u_{i,\theta})\Big]^{2}=\mathnormal{O}_{\mathnormal{P},X}\big(\|\gamma\|^{2}\big).$
  2. (b)
    $\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(m_{i}-s_{i,\theta})^{2}\Big]^{2}=\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{p^{3}}{n^{2}}\Big).$
  3. (c)
    For any measurable function $h:[-1,1]\to[-1,1]$, we have
$\operatorname{\mathbb{E}}\Big[\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big(h(B_{i})-\langle h\rangle_{\mu_{i,m_{i}+c_{i},\theta}}\big)\Big]$ $=O\Big(\frac{1}{\sqrt{p}}\Big),$
$\quad\operatorname{\mathbb{E}}\Big[\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big(h(B_{i})-\langle h\rangle_{\mu_{i,m_{i}+c_{i},\theta}}\Big)\langle h\rangle_{\mu_{i,m_{i}+c_{i},\theta}}\Big]$ $=O\Big(\frac{1}{\sqrt{p}}\Big).$
The final technical result provides some basic properties of the design matrix $X$ that will be useful in the sequel. The proof follows from Lemmas 4.1, 4.2 in [34].
Lemma A.11 (Design matrix properties).
Let $X$ be the random design from 3.1 and recall $A_{p}$ from 2.1. Then, for any $r\geq 2$, the following properties hold:
$\|A_{p}\|_{r\to r}=O_{P,X}\Big(\frac{p^{1-1/r}}{\sqrt{n}}\Big),\quad\alpha_{p}=\alpha_{p}(X):=\max_{\mathnormal{i}=1}^{\mathnormal{p}}\sum_{\mathnormal{j}=1}^{\mathnormal{p}}A_{p}(i,j)^{2}=O_{P,X}\Big(\frac{p}{n}\Big).$
Recall from Section 1.3 that $O_{P,X}$ denotes the sole randomness of $X$.
Proof of 4.1.
Recall the notation $u_{i},\hat{u}_{i}$ from 4.2. We also use the notations $s_{i}=\sum_{j\neq i}A_{p}(i,j)u_{j}$ and $\hat{s}_{i}=\sum_{j\neq i}A_{p}(i,j)\hat{u}_{j}$ throughout this proof.
  1. (a)
    We prove a general claim under a general parameter value $\theta$ that may depend on the data $y,X$. Let $h$ be any bounded function with $\|h\|_{\infty}\leq 1$. For $t\in\mathbb{R}$ and fixed indices $i,\theta$, define $H_{i,\theta}(t):=\langle h\rangle_{\mu_{i,t,\theta}}$. Using the properties of the complete conditional distribution of $\beta_{i}$ (see (56)), for $\beta\sim\operatorname{\mathbb{P}}_{\theta}$ and $m_{i}=m_{i}(\beta_{(-i)})=\sum_{j\neq i}A_{p}(i,j)\beta_{j}$, we have
    $\operatorname{\mathbb{E}}[h(\beta_{i})]=\operatorname{\mathbb{E}}[\operatorname{\mathbb{E}}[h(\beta_{i})\mid\beta_{(-i)}]]=\operatorname{\mathbb{E}}\langle h\rangle_{\mu_{i,m_{i}+w_{i},\theta}}=\operatorname{\mathbb{E}}H_{i,\theta}(m_{i}+w_{i}).$ (35)
    Our key argument is noting that
    $|\operatorname{\mathbb{E}}H_{i,\theta}(m_{i}+w_{i})-H_{i,\theta}(s_{i}+w_{i})|=O(\operatorname{\mathbb{E}}|m_{i}-s_{i}|)=O_{P,X}(\sqrt{\operatorname{\alpha_{\mathnormal{p}}}}).$ (36)
    Here, the first equality follows from a Taylor expansion and the fact that $h$ is bounded, and the second follows from writing $m_{i}-s_{i}=\sum_{j\neq i}A_{p}(i,j)(\beta_{j}-u_{j})$ and using a second moment bound for RFIMs (see Lemma A.10(a)). The hidden constants above do not depend on the prior hyperparameter $\theta$, and only on the compact support in 3.5.
    The identities for the mean and variance are immediate by taking $h(b)=b,b^{2}$, respectively in (36). The conclusion for the TV distance follows by spelling out the definition and using (36):
    $\mathrm{D_{TV}}\big(\operatorname{\mathbb{P}}_{\theta}(\beta_{i}),\mathbb{Q}_{\theta}(\beta_{i})\big)=\sup_{h(\cdot)=I(\cdot\in E),~E\subset\mathbb{R}}\Big|\operatorname{\mathbb{E}}\langle h\rangle_{\mu_{i,m_{i}+w_{i},\theta}}-\langle h\rangle_{\mu_{i,s_{i}+w_{i},\theta}}\Big|=O_{P,X}(\sqrt{\operatorname{\alpha_{\mathnormal{p}}}}).$
    Finally we use the bound on $\operatorname{\alpha_{\mathnormal{p}}}$ from Lemma A.11.
  2. (b)
    Following part (a), let $h$ be any bounded function. Use two triangle inequalities (first and third line) to bound
    $|\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}_{\theta_{0}}}h(\beta_{i})$ $-\operatorname{\mathbb{E}}_{\mathbb{Q}_{\hat{\theta}_{\mathnormal{p}}}}h(\beta_{i})|\leq|\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}_{\theta_{0}}}h(\beta_{i})-\operatorname{\mathbb{E}}_{\mathbb{Q}_{\theta_{0}}}h(\beta_{i})|+|\operatorname{\mathbb{E}}_{\mathbb{Q}_{\theta_{0}}}h(\beta_{i})-\operatorname{\mathbb{E}}_{\mathbb{Q}_{\hat{\theta}_{\mathnormal{p}}}}h(\beta_{i})|$
    $=O_{P,X}(\sqrt{\operatorname{\alpha_{\mathnormal{p}}}})+|H_{i,\theta_{0}}(s_{i}+w_{i})-H_{i,\hat{\theta}_{\mathnormal{p}}}(\hat{s}_{i}+w_{i})|$
    $\leq O_{P,X}(\sqrt{\operatorname{\alpha_{\mathnormal{p}}}})+|s_{i}-\hat{s}_{i}|+|H_{i,\theta_{0}}(\hat{s}_{i}+w_{i})-H_{i,\hat{\theta}_{\mathnormal{p}}}(\hat{s}_{i}+w_{i})|$
    $=O_{P,X}(\sqrt{\operatorname{\alpha_{\mathnormal{p}}}})+|s_{i}-\hat{s}_{i}|+O(\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|).$ (37)
    Here, the second line used (36) from part (a), and the fourth line used Lemma A.9.
    To control $|s_{i}-\hat{s}_{i}|$ in (37), let $v_{i}:=\hat{u}_{i}-u_{i}$. Again, by Lemma A.9 (with $h(b)=b$ and $t=\hat{s}_{i}+w_{i}$), we have
    $\hat{u}_{i}=\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime}\big(\hat{s}_{i}+w_{i}\big)=\psi_{i,\theta_{0}}^{\prime}\big(\hat{s}_{i}+w_{i}\big)+E_{i},\quad\text{where}\quad|E_{i}|\lesssim\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|.$
    Hence, using the definition of $\hat{u}_{i},u_{i}$ in 4.2, we can write
    $v_{i}=\psi_{i,\theta_{0}}^{\prime}\big(\hat{s}_{i}+w_{i}\big)-\psi_{i,\theta_{0}}^{\prime}\big(s_{i}+w_{i}\big)$ $=\psi_{i,\theta_{0}}^{\prime\prime}(\xi_{i})(\hat{s}_{i}-s_{i})+E_{i}$
    $=\psi_{i,\theta_{0}}^{\prime\prime}(\xi_{i})\sum_{j\neq i}A_{p}(i,j)v_{j}+E_{i},$ (38)
    for some $\xi_{i}\in(s_{i},\hat{s}_{i})$. By setting $D_{p}$ as a $p\times p$ matrix with $D_{p}(i,j):=\psi_{i,\theta_{0}}^{\prime\prime}(\xi_{i})A_{p}(i,j)$, we get a vector equation $v=D_{p}v+E$ (with $E=(E_{1},\ldots,E_{p}$)). As
    $\|D_{p}\|\leq\|\text{{Diag}}(\psi_{i,\theta_{0}}^{\prime\prime}(\xi_{i}))\|\|A_{p}\|<\frac{1}{2}$
    with high probability (using the operator norm bound for $A_{p}$ in Lemma A.11), we have
    $\|v\|\leq\|(I_{p}-D_{p})^{-1}\|\|E\|\leq 2\|E\|=O_{P,X}(\sqrt{p}\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|).$
    Then, by a Cauchy-Schwartz, we have
    $|s_{i}-\hat{s}_{i}|=\Big|\sum_{j\neq i}A_{p}(i,j)v_{j}\Big|\leq\sqrt{\sum_{\mathnormal{j}=1}^{\mathnormal{p}}A_{p}(i,j)^{2}}\|v\|=O_{P,X}\Big(\frac{p}{\sqrt{n}}\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|\Big).$
    The desired conclusions follow by plugging this into (37). In particular, the TV distance claim follows by taking $h$ as indicator functions, and the claims for the mean and variance follows by taking $h(b)=b$ and $b^{2}$.
The proof of 4.3 builds upon recently developed finite-sample error bounds for the KS distance between the projected posterior and a Gaussian, which we state below.
Lemma A.12 (Theorem 2.4 in [34]).
Suppose $\beta=(\beta_{1},\ldots,\beta_{p})\in[-1,1]^{p}\stackrel{{\scriptstyle d}}{{\equiv}}\tilde{\operatorname{\mathbb{P}}}_{\hat{\theta}_{\mathnormal{p}}}^{A_{p},w}$ (see A.1). Define error terms $R_{1},R_{2}$ as:
$R_{1}:=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big(\sum_{\mathnormal{j}\neq i}^{\mathnormal{p}}A_{p}(i,j)q_{i}(\psi_{j,\hat{\theta}_{\mathnormal{p}}}^{\prime\prime}(w_{j})-\hat{\upsilon}_{p})\Big)^{2},\quad R_{2}:=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big(\sum_{\mathnormal{j}\neq i}^{\mathnormal{p}}A_{p}(i,j)\psi_{j,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{j})\Big)^{2},$
where $\hat{\upsilon}_{p}$ is defined in 4.3. Then, under 3.1, the following bound holds, where the hidden constant is an universal:
$d_{KS}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}(\beta_{i}-\hat{u}_{i}),N(0,\hat{\upsilon}_{p})\mid y,X\right)\lesssim\sqrt{R_{1}}+\sqrt{\alpha_{p}R_{2}}+\frac{\sqrt{p}\alpha_{p}+\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}^{4}}\sqrt{R_{2}}+\|q\|_{\infty}}{\hat{\upsilon}_{p}}+\|A_{p}\|.$
As the original result in [34] allows any choice of prior/base-measure $\mu$, it also holds under the estimated prior $\mu=\mu_{\hat{\theta}_{\mathnormal{p}}}$. Here, we have slightly modified the statement in [34] (i.e. replaced an additional error term $R_{3p}$ there in terms of $R_{2}$) by replacing the bound in the third display in [pg 17, 34] to the following:
$\Big|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}^{2}(\psi_{i}^{\prime\prime}(s_{i}+w_{i})-\psi_{i}^{\prime\prime}(w_{i}))\Big|\leq\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}^{2}|s_{i}|\leq\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}^{4}}\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}s_{i}^{2}}\lesssim\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}^{4}}\sqrt{R_{2}},$
We also need the following technical lemmas.
Lemma A.13.
Under Assumptions 3.1 and 3.4, for any $\hat{\theta}_{\mathnormal{p}}$ with $\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|=O_{P}\Big(\frac{1}{\sqrt{p}}\Big)$, we have
$\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime}(s_{i,\hat{\theta}_{\mathnormal{p}}}+w_{i})=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\psi_{0,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})+o_{P}(1).$
Lemma A.14 (LLN with general weights).
Assume R5. Let $g:\mathbb{R}\to\mathbb{R}$ be a $\mathcal{C}^{1}$ function where $g^{\prime}\in\operatorname{\mathsf{{Poly}}}(r)$, and let $r_{i}$ be constants such that $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}r_{i}\to\gamma\in\mathbb{R}$ and $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}r_{i}^{2}=o(1)$. Then, we have
$\sum_{\mathnormal{i}=1}^{\mathnormal{p}}r_{i}g(w_{i})\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}\gamma\operatorname{\mathbb{E}}g(W_{\theta_{0}}).$
The proof of Lemma A.14 follows from direct modifications from Lemma A.2 (which is when $r_{i}\equiv\frac{1}{p}$). So we omit the proof for brevity.
Proof of 4.3.
  1. (a)
    We show that $\hat{\upsilon}_{p}\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}\upsilon(\theta_{0})$ under the double-dipping of the data $y$ on $w_{i}$ and $\hat{\theta}_{\mathnormal{p}}$. For this goal, we already know from 4.2 that
    $\upsilon_{p}\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}\upsilon(\theta_{0}),$
    so it suffices to show $\hat{\upsilon}_{p}-\upsilon_{p}=o_{P}(1)$. Indeed, this follows from Lemma A.9 as
    $|\hat{\upsilon}_{p}-\upsilon_{p}|=|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}^{2}[\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime\prime}(w_{i})-\psi_{i,\theta_{0}}^{\prime\prime}(w_{i})]|\leq\sup_{w\in\mathbb{R}}|\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime\prime}(w)-\psi_{i,\theta_{0}}^{\prime\prime}(w)|=O(\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|).$
    The RHS is $o_{P}(1)$ as we assume that $\hat{\theta}_{\mathnormal{p}}$ is consistent.
  2. (b)
    It suffices to show that the finite-sample bound in the RHS of Lemma A.12 is small. Here we use the bounds $\operatorname{\alpha_{\mathnormal{p}}}=\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{p}{n}\Big),\|A_{p}\|=\mathnormal{O}_{\mathnormal{P},X}\Big(\sqrt{\frac{p}{n}}\Big)$, which are stated in Lemma A.11. By these bounds, it also follows that
    $R_{1}=\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{p}{n}\Big),\quad R_{2}=\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{p^{2}}{n}\Big)$ (39)
    where the hidden constants do not depend on the estimated prior $\hat{\theta}_{\mathnormal{p}}$. The proof is complete by plugging these into Lemma A.12, alongside the assumption on $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}^{4}$.
  3. (c)
    Recall from 4.2 that
    $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}(\beta_{i}-u_{i})\mid y,X\xrightarrow[\operatorname{\mathbb{P}}_{\theta_{0}}]{d}N(0,\upsilon(\theta_{0})).$
    By the conditional converging together lemma (see Lemma B.5), it suffices to show the following limit:
    $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}(u_{i}-\hat{u}_{i})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(0,\gamma^{2}J(\theta_{0})^{\top}S(\theta_{0})J(\theta_{0})\Big),$ (40)
    as the two displays above imply
    $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}(\beta_{i}-\hat{u}_{i})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(0,\upsilon(\theta_{0})+\gamma^{2}J(\theta_{0})^{\top}S(\theta_{0})J(\theta_{0})\Big).$
    Then, (26) follows by modifying the variance to the sample-based quantities, using the conclusion of part (a) alongside the continuous mapping theorem. The claim under $\gamma=0$ also follows from (40), as (40) implies
    $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}(u_{i}-\hat{u}_{i})\mid y,X\operatorname*{\xrightarrow{\mathnormal{P}}}0.$
    It remains to show (40). By applying Lemma A.13 for $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\hat{u}_{i}$ and $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}u_{i}$, we can write
    $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}(\hat{u}_{i}-u_{i})=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\Big[\psi_{0,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})-\psi_{0,\theta_{0}}^{\prime}(w_{i})\Big]+o_{P}(1).$
    For notational simplicity, set $g(\theta,w):=\psi_{0,\theta}^{\prime}(w)$. Then, a Taylor approximation gives
    $g(\hat{\theta}_{\mathnormal{p}},w_{i})-g(\theta_{0},w_{i})=(\hat{\theta}_{\mathnormal{p}}-\theta_{0})^{\top}\frac{\partial g}{\partial\theta}(\theta_{0},w_{i})+O(\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|^{2}),$
    where we used the fact that $\frac{\partial^{2}g}{\partial\theta^{2}}$ is uniformly bounded under 3.5. By using the LLN in Lemma A.14, we get
    $\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\frac{\partial g}{\partial\theta}(\theta_{0},w_{i})\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}\gamma\operatorname{\mathbb{E}}\frac{\partial g}{\partial\theta}(\theta_{0},W_{\theta_{0}})=\gamma J(\theta_{0}).$
    Combining this alongside the limiting distribution $\sqrt{p}(\hat{\theta}_{\mathnormal{p}}-\theta_{0})\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,S(\theta_{0})),$ we have
    $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\Big[\psi_{0,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})-\psi_{0,\theta_{0}}^{\prime}(w_{i})\Big]=(\hat{\theta}_{\mathnormal{p}}-\theta_{0})^{\top}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\frac{\partial g}{\partial\theta}(\theta_{0},w_{i})+o_{P}(1)\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(0,\gamma^{2}J(\theta_{0})^{\top}S(\theta_{0})J(\theta_{0})\Big),$
    completing the proof of (40).
A.4 Proofs from Section 5
Proof of 5.1.
Throughout the proof, all convergence (in distribution/probability) statements are conditioned on $X$. Consider the asymptotic scaling $p\ll n^{3/4}$ and define a function $G(\theta):=V(\theta)^{-1}\kappa(\theta)$. By writing
$\sqrt{p}(\tilde{\theta}^{\mathsf{d}}-\theta_{0})=\sqrt{p}\big(\hat{\theta}^{\mathsf{vEB}}-\theta_{0}-\frac{p}{n}G(\theta_{0})\big)-\frac{p\sqrt{p}}{n}\big(G(\hat{\theta}^{\mathsf{vEB}})-G(\theta_{0})\big),$
it suffices to show the following limits:
$\sqrt{p}\big(\hat{\theta}^{\mathsf{vEB}}-\theta_{0}-\frac{p}{n}G(\theta_{0})\big)\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,V(\theta_{0})^{-1}),\quad\frac{p\sqrt{p}}{n}\big(G(\hat{\theta}^{\mathsf{vEB}})-G(\theta_{0})\big)\operatorname*{\xrightarrow{\mathnormal{P}}}0.$ (41)
Recall from (32) that
$\sqrt{p}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})=\big(V(\theta_{0})^{-1}+o_{P}(1)\big)\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0}),$
so we can write the left term of (41) as
$\sqrt{p}\big(\hat{\theta}^{\mathsf{vEB}}-\theta_{0}-\frac{p}{n}G(\theta_{0})\big)=V(\theta_{0})^{-1}\Big[\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})-\frac{p\sqrt{p}}{n}\kappa(\theta_{0})\Big]+o_{P}(1).$
The first limit in (41) follows by recalling from part (c) of Lemma A.4 that the limit
$\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})-\frac{p\sqrt{p}}{n}\kappa(\theta_{0})\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,V(\theta_{0})).$
To show the second claim in (41), use a Taylor expansion to write
$G(\hat{\theta}^{\mathsf{vEB}})-G(\theta_{0})=G^{\prime}(\xi)(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})=O_{p}\Big(\frac{1}{\sqrt{p}}+\frac{p}{n}\Big),$
for some $\xi\in(\hat{\theta}^{\mathsf{vEB}},\theta_{0})$. Here we used (a) the fact that $G^{\prime}(\cdot)$ is bounded (since it is a continuous function and $\Theta$ is compact), (b) the bound for $\hat{\theta}^{\mathsf{vEB}}-\theta_{0}$ in $3.$ Now, the proof is complete by multiplying both sides by $\frac{p\sqrt{p}}{n}$. ∎
Appendix B Proof of lemmas
B.1 Proof of Lemmas from the main paper
Proof of Lemma 3.1.
By marginalizing out the joint distribution of $(B_{\theta},W_{\theta})$ in 3.1, we have
$e^{\tilde{\ell}(\theta;t)}=\operatorname{\mathbb{P}}_{W_{\theta}}(t)=\int\frac{e^{-\frac{(t-d_{0}b)^{2}}{2d_{0}}}}{\sqrt{2\pi d_{0}}}d\mu_{\theta}(b)=\frac{e^{-\frac{t^{2}}{2d_{0}}}}{\sqrt{2\pi d_{0}}}\int e^{tb-\frac{d_{0}b^{2}}{2}+\ell(\theta;b)}db.$
Recalling the notation $v^{2}$ for a vector $v\in\mathbb{R}^{k}$ (see Section 1.3), direct computations give
$\nabla^{2}\tilde{\ell}(\theta;t)$ $=\frac{\partial^{2}\big(\log\int e^{tb-\frac{d_{0}b^{2}}{2}+\ell(\theta;b)}db\big)}{\partial\theta^{2}}$
$=\operatorname{\mathbb{E}}\big[\nabla^{2}\ell(\theta;B_{t,d_{0},\theta})\big]+\operatorname{\mathbb{E}}\big[\nabla\ell(\theta;B_{t,d_{0},\theta})^{2}\big]-\big[\operatorname{\mathbb{E}}\nabla\ell(\theta;B_{t,d_{0},\theta})\big]^{2}.$
By taking an outer expectation with respect to $t\stackrel{{\scriptstyle d}}{{\equiv}}W_{\theta}$ and noting that $B_{W_{\theta},d_{0},\theta}\stackrel{{\scriptstyle d}}{{\equiv}}B_{\theta}$, we get
$\operatorname{\mathbb{E}}\big[\nabla^{2}\ell(\theta;B_{W_{\theta},d_{0},\theta})\big]+\operatorname{\mathbb{E}}[\nabla\ell(\theta;B_{W_{\theta},d_{0},\theta})^{2}]-\operatorname{\mathbb{E}}\Big[\operatorname{\mathbb{E}}\big(\nabla\ell(\theta;B_{W_{\theta},d_{0},\theta})\mid W_{\theta}\big)\Big]^{2}$
$=$ $\operatorname{\mathbb{E}}\big[\nabla^{2}\ell(\theta;B_{\theta})\big]+\operatorname{\textnormal{Var}}\nabla\ell(\theta;B_{\theta})-\operatorname{\textnormal{Var}}\Big[\operatorname{\mathbb{E}}\big(\nabla\ell(\theta;B_{\theta})\mid W_{\theta}\big)\Big]=-V(\theta).$ (42)
Here the first equality used the first Bartlett’s identity $\operatorname{\mathbb{E}}\nabla\ell(\theta;B_{\theta})=0$, and the second equality follows from the second Bartlett’s identity $I(\theta)=-\operatorname{\mathbb{E}}\big[\nabla^{2}\ell(\theta;B_{\theta})\big]=\operatorname{\textnormal{Var}}\nabla\ell(\theta;B_{\theta})$. ∎
For Lemma 3.2, we use the following property for sub-Gaussian mixtures, whose proof is postponed to Appendix C.
Lemma B.1.
Let $X$ be a $M$-component mixture of sub-Gaussian random variables, where the $m$-th mixture component has mean $\mu_{m}$ and sub-Gaussian norm $K_{m}$. Then, $X$ is sub-Gaussian with
$\|X\|_{\psi_{2}}^{2}\leq K_{\max}^{2}+\frac{1}{4}(\mu_{\max}-\mu_{\min})^{2},$
where $K_{\max}:=\max_{m=1}^{M}K_{m},~\mu_{\max}:=\max_{m=1}^{K}\mu_{m},~\mu_{\min}:=\min_{m=1}^{K}\mu_{m}$.
Proof of Lemma 3.2.
  1. (a)
    This is immediate as the sub-Gaussian norm of a bounded measure is uniformly bounded.
  2. (b)
    For a $\kappa$-log-concave measure $\nu$, the Bakry-Émery inequality gives the LSI [2], and consequently imply sub-Gaussianity of $\nu$: $\|\nu\|_{\psi_{2}}\leq\frac{1}{\kappa}$. R5 directly follows from this observation, as $B_{\theta_{0}}$ is sub-Gaussian.
    Noting that $B_{t,d,\theta}$ has density proportional to $e^{tb-\frac{db^{2}}{2}+\ell(\theta;b)}$, set $(b):=tb-\frac{db^{2}}{2}+\ell(\theta;b). ff$ is $d$-concave (i.e. strongly concave) as $f(b)+\frac{db^{2}}{2}$ is concave for all $b$, and hence $B_{t,d,\theta}$ is sub-Gaussian with norm bounded by $\frac{1}{d}<\frac{2}{d_{0}}$ as long as $d>\frac{d_{0}}{2}$. Hence, the second claim in R6 holds.
    Proceeding to prove the first claim in R6, define a function $\mathfrak{m}:B_{\epsilon}(d_{0})\times\Theta\to\mathbb{R}$ by setting
    $\mathfrak{m}(d,\theta):=\operatorname{\mathbb{E}}[B_{0,d,\theta}]=\frac{\int be^{-db^{2}/2}d\mu_{\theta}(b)}{\int e^{-db^{2}/2}d\mu_{\theta}(b)},\quad\forall d\in B_{\epsilon}(d_{0}).$
    As this is a continuous function (of $d,\theta$) on a compact domain, it is bounded. For example, regarding the supremum over $d$, direct computations show that
    $|\partial_{d}\mathfrak{m}(d,\theta)|=\frac{1}{2}|\operatorname{\textnormal{Cov}}(B_{0,d,\theta},B_{0,d,\theta}^{2})|\leq\frac{1}{2}\sqrt{\operatorname{\textnormal{Var}}(B_{0,d,\theta})\operatorname{\textnormal{Var}}(B_{0,d,\theta}^{2})}\lesssim_{d_{0}}1,\quad\forall d>\frac{d_{0}}{2},~~\theta,$
    so $\mathfrak{m}(d,\theta)=\mathfrak{m}(d_{0},\theta)+O(1).$ The final inequality uses the sub-Gaussianity $\|B_{0,d,\theta}\|_{\psi_{2}}\leq\frac{2}{d_{0}}$ (note that this implies that $B_{0,d,\theta}^{2}$ is sub-exponential).
  3. (c)
    R5 is immediate by applying Lemma B.1, as each mixture component of $B_{\theta_{0}}$ is sub-Gaussian by parts (a) and (b). The first part of R6 is also bounded by the tower property. To check the second part of R6, let $p_{k}(b)$ denote the conditional density (if we consider a degenerate mixture component, consider the base measure $\nu=\delta_{0}+\lambda$ where $\lambda$ denotes the Lebesgue measure) of the $k$th mixture component, and observe that $B_{t,d,\theta}$ has density:
    $\operatorname{\mathbb{P}}(B_{t,d,\theta}=b)\propto\sum_{k=1}^{K}\pi_{k}p_{k}(b)e^{tb-\frac{db^{2}}{2}}.$
    This is a finite mixture of the component-wise tilts $B^{(k)}_{t,d}$, which have density $p_{k}(b)e^{tb-\frac{db^{2}}{2}}$. By parts (a) and (b), each $B^{(k)}_{t,d}$ are sub-Gaussian with $\|B^{(k)}_{t,d}\|_{\psi_{2}}\leq K$ for some constant $K>0$. Also, using the properties of exponential tilting (in particular, we appropriately modify (54) in Lemma A.5), we can show that $\operatorname{\mathbb{E}}B^{(k)}_{t,d}=O(1+K|t|)$. Now, the conclusion follows by applying Lemma B.1.
B.2 Proof of upper bound lemmas
Proof of Lemma A.3.
Recall the notation $(B_{\theta},W_{\theta})$ from 3.1 and denote the corresponding joint measure as $\operatorname{\mathbb{P}}_{B_{\theta},W_{\theta}}$. Then, we have
$\frac{d\operatorname{\mathbb{P}}_{B_{\theta},W_{\theta}}}{d\mu_{\theta}\otimes d\lambda}(b,t)=\frac{1}{\sqrt{2\pi d_{0}}}e^{-\frac{t^{2}}{2d_{0}}+tb-\frac{d_{0}b^{2}}{2}}.$
Here, $\lambda$ denotes the Lebesgue measure on $\mathbb{R}$. By marginalizing out $B_{\theta}$, we have
$\frac{d\operatorname{\mathbb{P}}_{W_{\theta}}}{d\lambda}(t)=\frac{e^{-\frac{t^{2}}{2d_{0}}}}{\sqrt{2\pi d_{0}}}\int e^{tb-\frac{d_{0}b^{2}}{2}}d\mu_{\theta}(b).$ (43)
Now, by the definition of $\mathrm{D_{KL}}$ and $F_{0}$,
$\mathrm{D_{KL}}(\operatorname{\mathbb{P}}_{W_{\theta_{0}}}\|\operatorname{\mathbb{P}}_{W_{\theta}})$ $=\operatorname{\mathbb{E}}\Big[\log\Big(\int\exp(W_{\theta_{0}}b-\frac{d_{0}b^{2}}{2})d\mu_{\theta_{0}}(b)\Big)-\log\Big(\int\exp(W_{\theta_{0}}b-\frac{d_{0}b^{2}}{2})d\mu_{\theta}(b)\Big)\Big]$
$=F_{0}(\theta_{0})-F_{0}(\theta)\geq 0.$
The conclusion holds if the final inequality is strict for any $\theta\neq\theta_{0}$. This is because the parametric family $\{\operatorname{\mathbb{P}}_{W_{\theta}}\}_{\theta\in\Theta}$ (see (43)) is identifiable (i.e. $\operatorname{\mathbb{P}}_{W_{\theta}}=\operatorname{\mathbb{P}}_{W_{\theta^{\prime}}}$ implies $\theta=\theta^{\prime}$), which follows from the injectivity of the Laplace transform for finite measures, alongside the identifiability of $\mu_{\theta}$ from R0 in 3.2. ∎
Next, we prove Lemma A.2, using some technical lemmas which are proved in Appendix C.
Lemma B.2.
Let $(X_{p},Y_{p})$ be a sequence of random variables. Then, $X_{p}\operatorname*{\xrightarrow{\mathnormal{P}}}0$ if and only if $X_{p}\mid Y_{p}\operatorname*{\xrightarrow{\mathnormal{P}}}0$.
Lemma B.3 (Extension of Lemma B.3 in [34]).
Let $h$ be any $C^{1}$ function where $h^{\prime}\in\operatorname{\mathsf{{Poly}}}_{r}$, (i.e., $h^{\prime}$ is bounded by a degree $r$ polynomial). For random variables $V\sim N(m,d),V_{0}\sim N(m_{0},d_{0})$ with $d\leq 2d_{0}$, we have
$|\operatorname{\mathbb{E}}h(V)-\operatorname{\mathbb{E}}h(V_{0})|\lesssim_{d_{0},r}|m-m_{0}|(1+|m_{0}|^{r}+|m-m_{0}|^{r})+|d-d_{0}|.$
Proof of Lemma A.2.
  1. (a)
    We separate the proof into two steps. Throughout, $K$ denotes absolute constants whose exact value may differ line-by-line.
    Step 1. Recall that $w\mid X,\beta^{\star}\sim N(\Sigma\beta^{\star},\Sigma)$ where $\Sigma=\sigma^{-2}X^{\top}X$. By the Gaussian Poincare inequality [e.g. Theorem 3.25, 57], the growth condition, and Lemma A.7, we have
    $\operatorname{\textnormal{Var}}\Big(\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(w_{i})\mid X,\beta^{\star}\Big)\leq$ $\frac{1}{p^{2}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}[g^{\prime}(w_{i})^{2}\mid X,\beta^{\star}]\leq\frac{1}{p^{2}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[1+\operatorname{\mathbb{E}}\big[|w_{i}|^{2r}\mid X,\beta^{\star}\big]\Big]=O_{P}\Big(\frac{1}{p}\Big).$
    For notational simplicity, let $\phi(m,d):=\operatorname{\mathbb{E}}g(N(m,d))$. This allows us to express the conditional mean of $g(w_{i})$ as:
    $\operatorname{\mathbb{E}}[g(w_{i})\mid X,\beta^{\star}]=\phi(d_{i}\beta_{i}^{\star}-\mathcal{F}_{i},d_{i}),$
    and the variance bound implies
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(w_{i})-\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\phi(d_{i}\beta_{i}^{\star}-\mathcal{F}_{i},d_{i})\mid X,\beta^{\star}\operatorname*{\xrightarrow{\mathnormal{P}}}0.$
    Step 2. Noting that $\operatorname{\mathbb{E}}g(W_{\theta_{0}})=\operatorname{\mathbb{E}}\Big[\operatorname{\mathbb{E}}\big[g(W_{\theta_{0}})\mid B_{\theta_{0}}\big]\Big]=\operatorname{\mathbb{E}}\phi(d_{0}B_{\theta_{0}},d_{0})$, we claim $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\phi(d_{i}\beta_{i}^{\star}-\mathcal{F}_{i},d_{i})\mid X\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}\phi(d_{0}B_{\theta_{0}},d_{0}).$ It suffices to individually show:
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big[\phi(d_{i}\beta_{i}^{\star}-\mathcal{F}_{i},d_{i})-\phi(d_{0}\beta_{i}^{\star},d_{0})\big]\mid X$ $\operatorname*{\xrightarrow{\mathnormal{P}}}0,\quad\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\phi(d_{0}\beta_{i}^{\star},d_{0})\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}\phi(d_{0}B_{\theta_{0}},d_{0}).$ (44)
    The second claim in (44) follows from the LLN for i.i.d. random variables. To show the first claim in (44), we use Lemma B.3 to write
    $|\phi(d_{i}\beta_{i}^{\star}-\mathcal{F}_{i},d_{i})-\phi(d_{0}\beta_{i}^{\star},d_{0})|$ (45)
    $\lesssim_{d_{0},r}$ $|d_{i}-d_{0}|+|(d_{i}-d_{0})\beta_{i}^{\star}-\mathcal{F}_{i}|\Big(1+|d_{0}\beta_{i}^{\star}|^{r}+|(d_{i}-d_{0})\beta_{i}^{\star}-\mathcal{F}_{i}|^{r}\Big)$
    $\lesssim_{d_{0},r}$ $|d_{i}-d_{0}|+\big(|(d_{i}-d_{0})\beta_{i}^{\star}|+|\mathcal{F}_{i}|\big)\Big(1+|\beta_{i}^{\star}|^{r}+|\mathcal{F}_{i}|^{r}\Big).$ (46)
    By Cauchy Schwartz alongside moment bounds,
    $\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big[\phi(d_{i}\beta_{i}^{\star}-\mathcal{F}_{i},d_{i})-\phi(d_{0}\beta_{i}^{\star},d_{0})\big]\Big]^{2}$
    $\lesssim$ $p\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(d_{i}-d_{0})^{2}+\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big(|(d_{i}-d_{0})\beta_{i}^{\star}|^{2}+|\mathcal{F}_{i}|^{2}\Big)\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big(1+|\beta_{i}^{\star}|^{2r}+|\mathcal{F}_{i}|^{2r}\Big)$
    $=$ $O_{P,X}\Big(\frac{p^{2}}{n}\Big)+O_{P}\Big(\frac{p}{n}+\frac{p^{2}}{n}\Big)\Big(p+\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|\beta_{i}^{\star}|^{2r}+\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|\mathcal{F}_{i}|^{2r}\Big)$
    $=$ $O_{P,X}\Big(\frac{p^{2}}{n}\Big)+O_{P}\Big(\frac{p^{3}}{n}\Big)=o_{P}(p^{2}),$
    which implies the first claim. Here, the third line used simple expectation computations: $\operatorname{\mathbb{E}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(d_{i}-d_{0})^{2}=p/n,\operatorname{\mathbb{E}}[\mathcal{F}_{i}^{2}\mid\beta^{\star}]\leq\frac{\|\beta^{\star}\|^{2}}{n}$. The fourth line used Lemma A.7. Finally, the conditional limit in (44) follows from Lemma B.2.
  2. (b)
    Given a function $h\in\operatorname{\mathsf{{Poly}}}(r)$, define a function $H:\mathbb{R}\to\mathbb{R}$ as
    $H(t):=\operatorname{\mathbb{E}}h(B_{t})=\frac{\int h(b)\exp[tb-\frac{d_{0}b^{2}}{2}]d\mu_{\theta_{0}}(b)}{\int\exp\big[tb-\frac{d_{0}b^{2}}{2}\big]d\mu_{\theta_{0}}(b)}.$
    Note that taking the derivative and using Lemma A.5 gives
    $|H^{\prime}(t)|=|\operatorname{\textnormal{Cov}}(B_{t},h(B_{t}))|\lesssim\operatorname{\mathbb{E}}[|B_{t}|+|B_{t}|^{r+1}]\lesssim 1+|t|^{3(r+1)},$
    so we have $H^{\prime}\in\operatorname{\mathsf{{Poly}}}(3(r+1))$.
    Using this notation, we have
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}h(B_{w_{i},d_{i}})$ $=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}h(B_{w_{i},d_{0}})+o_{P,X}(1)$
    $=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}H(w_{i})+o_{P,X}(1)\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}H(W_{\theta_{0}})=\operatorname{\mathbb{E}}h(B_{\theta_{0}}).$
    Here, the first equality used Lemma A.6 (noting that $h$ has polynomial growth), followed by using Lemma A.7 to bound the error term. The limit in the last line follows from part (a) of this lemma (noting that $H^{\prime}$ has polynomial growth). The final equality uses the observation $B_{\theta_{0}}\mid W_{\theta_{0}}\stackrel{{\scriptstyle d}}{{\equiv}}B_{W_{\theta_{0}},d_{0}}$ to write
    $H(W_{\theta_{0}})=\operatorname{\mathbb{E}}\big[h(B_{\theta_{0}})\mid W_{\theta_{0}}\big].$
    The second identity follows by modifying the above argument by setting $H(w):=[\operatorname{\mathbb{E}}h(B_{w})]^{2}.$
Remark B.1.
In part (a) of the above proof, by modifying the computations in (46) and the following line, we can show the following:
$\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\phi(d_{i}\beta_{i}-\mathcal{F}_{i},d_{i})-\phi(d_{0}\beta_{i}^{\star}-\mathcal{F}_{i},d_{0})]\Big]^{2}$
$\lesssim_{d_{0},r}$ $p\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(d_{i}-d_{0})^{2}+\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|(d_{i}-d_{0})\beta_{i}^{\star}|^{2}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(1+|\beta_{i}^{\star}|^{2r}+|\mathcal{F}_{i}|^{2r})=O_{P}\Big(\frac{p^{2}}{n}\Big).$
Note that by replacing the first argument of each $\phi$ in the summand from $d_{i}\beta_{i}^{\star}-\mathcal{F}_{i}$ to $d_{0}\beta_{i}^{\star}-\mathcal{F}_{i}$ (as opposed to replacing to $d_{0}\beta_{i}^{\star}$ as in (46)), we have a better error bound. We will use this fact to prove the CLT Lemma A.4, where a tighter bound is required.
To prove Lemma A.4, we require a crucial technical lemma, which provides an improved moment bound compared to Step 2 in the proof of Lemma A.2.
Lemma B.4.
Suppose R5 holds. Let $\Phi:\mathbb{R}\to\mathbb{R}$ be a $\mathcal{C}^{3}$ function with $\Phi^{\prime},\Phi^{\prime\prime},\Phi^{\prime\prime\prime}\in\operatorname{\mathsf{{Poly}}}(r)$. Then, for $\kappa_{\Phi}(\theta):=\frac{1}{2}\operatorname{\mathbb{E}}[B_{\theta}^{2}]\operatorname{\mathbb{E}}[\Phi^{\prime\prime}(d_{0}B_{\theta})]$, we have
$\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi(d_{0}\beta_{i}^{\star}-\mathcal{F}_{i})=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi(d_{0}\beta_{i}^{\star})+\frac{p^{2}}{n}\kappa_{\Phi}(\theta_{0})+O_{P}\Big(\frac{p}{\sqrt{n}}+\frac{p^{5/2}}{n^{3/2}}\Big).$
Here, the hidden constant only depends on $d_{0},r$, and the moments of $B_{\theta_{0}}$.
Finally, we will need a technical result which is a restatement of Lemma A.13 in [33].
Lemma B.5 (Conditional converging together).
Let $G_{p}=G_{p}(y,X,\beta^{\star}),~\mu_{p}(X,\beta^{\star})$ be $k$-dimensional random vectors, and let $T_{1},T_{2}\succ 0$ be $k\times k$ deterministic matrices. Suppose $G_{p}\mid X,\beta^{\star}\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,T_{1})$ and $\mu_{p}\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,T_{2}).$ Then, we have $G_{p}+\mu_{p}\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,T_{1}+T_{2}).$
Proof of Lemma A.4.
(a) We first prove part (a). Here, all results are stated for general $(p,n)$, and we will specify whenever the assumption $\frac{p^{3/2}}{n}\to\delta$ is used. Throughout this proof, $\theta=\theta_{0}$ is assumed and we abbreviate $B_{t,d}=B_{t,d,\theta_{0}}$.
Since $\nabla F_{i}(\theta_{0})=\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{w_{i},d_{i}})$, using Lemma A.6 followed by Lemma A.7 gives:
$\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{w_{i},d_{i}})$ $=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{w_{i},d_{0}})+O_{P}\Big(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|d_{i}-d_{0}|(1+|w_{i}|^{3(r+2)})\Big)$
$=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{w_{i},d_{0}})+O_{P}\Big(\frac{p}{\sqrt{n}}\Big).$
For any $t\in\mathbb{R}$, define a vector-valued function $h:\mathbb{R}\to\mathbb{R}^{k}$ by
$h(t)=(h_{1}(t),\ldots,h_{k}(t))^{\top}:=\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{t,d_{0}}).$ (47)
Then, it suffices to prove that
$\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}h(w_{i})\operatorname*{\xrightarrow{\mathnormal{d}}}N(\delta\kappa(\theta_{0}),V(\theta_{0}))\quad\text{as}\quad\frac{p^{3/2}}{n}\to\delta.$ (48)
We separate the proof into two main steps. In the remainder of the proof, the quadratic tilt parameter will always be $d_{0}$, so we further abbreviate $B_{t}=B_{t,d_{0},\theta_{0}}$.
Step 1: Use the second-order Poincare inequality to get conditional normality.
Denote the conditional (on $X,\beta^{\star}$) mean and variance of the LHS of (48) as:
$\mu_{p}(X,\beta^{\star}):=\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}\Big[h(w_{i})\mid X,\beta^{\star}\Big],\quad\tau^{2}_{p}(X,\beta^{\star}):=\frac{1}{p}\operatorname{\textnormal{Var}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}h(w_{i})\mid X,\beta^{\star}\Big]$
We claim the conditional limit:
$\frac{1}{\sqrt{p}}\operatorname{\sum_{i=1}^{k}}h(w_{i})-\mu_{p}(X,\beta^{\star})\mid X,\beta^{\star}\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(0,\tau_{p}^{2}(X,\beta^{\star})\Big).$
We use the Cramér–Wold device to show vector-valued convergence. It suffices to show the following for any constant vector $\alpha\in\mathbb{R}^{k}$ with $\|\alpha\|=1$:
$\frac{1}{\sqrt{p}}\alpha^{\top}\Big[\operatorname{\sum_{i=1}^{k}}h(w_{i})-\mu_{p}(X,\beta^{\star})\Big]\mid X,\beta^{\star}\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(0,\alpha^{\top}\tau_{p}^{2}(X,\beta^{\star})\alpha\Big).$
We prove this claim for any fixed vector $\alpha$. By applying the unidimensional Gaussian Poincare inequality (see Theorem 2.2 in [13]), the TV distance between the LHS and the RHS (still conditioned on $X,\beta^{\star}$) is bounded above by $\frac{\kappa_{1}\kappa_{2}}{\tau_{p}^{2}(X,\beta^{\star})},$ where
$\kappa_{1}^{4}:=\frac{1}{p^{2}}\operatorname{\mathbb{E}}\Big[\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big[\sum_{a=1}^{k}\alpha_{a}h_{a}^{\prime}(w_{i})\big]^{2}\Big]^{2}\mid X,\beta^{\star}\Big],\quad\kappa_{2}^{4}:=\frac{1}{p^{2}}\operatorname{\mathbb{E}}\Big[\max_{\mathnormal{i}=1}^{\mathnormal{p}}\big[\sum_{a=1}^{k}\alpha_{a}h_{a}^{\prime\prime}(w_{i})\big]^{4}\mid X,\beta^{\star}\Big].$
$\alpha^{\top}\tau_{p}^{2}(X,\beta^{\star})\alpha$ is bounded away from 0 by computations in the next step (see Step 2-2, which shows that $\tau_{p}^{2}(X,\beta^{\star})$ converges in probability to a positive definite matrix), and it suffices to show $\kappa_{1}=O_{P}(1),\kappa_{2}=o_{P}(1)$.
To bound $\kappa_{1}$, use Cauchy-Schwartz (twice, for each square) alongside $\|\alpha\|=1$ to write
$\kappa_{1}^{4}\leq\frac{1}{p^{2}}\operatorname{\mathbb{E}}\Big[\big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\sum_{a=1}^{k}h_{a}^{\prime}(w_{i})^{2}\big]^{2}\mid X,\beta^{\star}\Big]\leq\frac{k}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\sum_{a=1}^{k}\operatorname{\mathbb{E}}\Big[h_{a}^{\prime}(w_{i})^{4}\mid X,\beta^{\star}\Big]$
Recalling the definition of $h$ from (47) and moving the differential inside the expectation, we get
$h^{\prime}(t)$ $=\operatorname{\textnormal{Cov}}(\nabla\ell(\theta_{0};B_{t}),B_{t}),$
$h^{\prime\prime}(t)$ $=\operatorname{\textnormal{Cov}}(B_{t}\nabla\ell(\theta_{0};B_{t}),B_{t})-\operatorname{\mathbb{E}}[B_{t}]\operatorname{\textnormal{Cov}}(\nabla\ell(\theta_{0};B_{t}),B_{t})-\operatorname{\mathbb{E}}[\nabla\ell(\theta_{0};B_{t})]\operatorname{\textnormal{Var}}(B_{t}).$
Hence, recalling from R4 that $\nabla\ell(\theta_{0};b)\in\operatorname{\mathsf{{Poly}}}(r)$ and using Lemma A.5, we have
$|h_{a}^{\prime}(t)|\leq\operatorname{\mathbb{E}}|B_{t}\partial_{\theta_{a}}\ell(\theta_{0};B_{t})|\lesssim\operatorname{\mathbb{E}}|B_{t}|+\operatorname{\mathbb{E}}|B_{t}|^{r+1}\lesssim_{r}1+|t|^{3(r+1)},\quad\forall a,t.$ (49)
Hence, by taking $t=w_{i}$ and taking an outer expectation over $w_{i}\mid X,\beta^{\star}$, we have
$\operatorname{\mathbb{E}}\big[h_{a}^{\prime}(w_{i})^{4}\mid X,\beta^{\star}\big]\lesssim_{r}1+\operatorname{\mathbb{E}}|w_{i}|^{12(r+1)},\quad\forall a.$ (50)
Plugging in this into the upper bound for $\kappa_{1}^{4}$ and using Lemma A.7 gives
$\kappa_{1}^{4}\lesssim\frac{k^{2}}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(1+\operatorname{\mathbb{E}}|w_{i}|^{12(r+1)})=O_{P}(1).$
Moving on to bounding $\kappa_{2}$, a similar application of Cauchy-Schwartz followed by the naive bound $\max(v_{1},\ldots,v_{p})\leq v_{1}+\ldots+v_{p}$ gives
$\kappa_{2}^{4}$ $\leq\frac{1}{p^{2}}\operatorname{\mathbb{E}}\Big[\max_{\mathnormal{i}=1}^{\mathnormal{p}}\big[\sum_{a=1}^{k}h_{a}^{\prime\prime}(w_{i})^{2}\big]^{2}\mid X,\beta^{\star}\Big]\leq\frac{k}{p^{2}}\operatorname{\mathbb{E}}\Big[\max_{\mathnormal{i}=1}^{\mathnormal{p}}\big[\sum_{a=1}^{k}h_{a}^{\prime\prime}(w_{i})^{4}\big]\mid X,\beta^{\star}\Big]$
$\leq\frac{k}{p^{2}}\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big[\sum_{a=1}^{k}h_{a}^{\prime\prime}(w_{i})^{4}\big]\mid X,\beta^{\star}\Big]=O_{P}\Big(\frac{1}{p}\Big).$
Here, the final bound follows from repeating the computations (49)-(50) for $h_{a}^{\prime\prime}$.
Step 2: Marginalize out $\beta^{\star}$.
Now, to prove (48), it suffices to separately prove the following:
$\mu_{p}(X,\beta^{\star})\operatorname*{\xrightarrow{\mathnormal{d}}}N(\delta\kappa(\theta_{0}),T_{1}),\quad\tau^{2}_{p}(X,\beta^{\star})\operatorname*{\xrightarrow{\mathnormal{P}}}T_{2},\quad T_{1}+T_{2}=V(\theta_{0}),$
where $T_{1},T_{2}\succ 0$ are deterministic $k\times k$ matrices (depending on $\theta_{0}$) defined as follows:
$T_{1}:=\operatorname{\textnormal{Var}}\Big[\operatorname{\mathbb{E}}\big[h(W_{\theta_{0}})\mid B_{\theta_{0}}\big]\Big],\quad T_{2}:=\operatorname{\mathbb{E}}\Big[\operatorname{\textnormal{Var}}\big[h(W_{\theta_{0}})\mid B_{\theta_{0}}\big]\Big].$ (51)
Assuming this, (48) follows from Lemma B.5. We separate out the proof of each claim into three sub-steps.
Step 2-1: $\mu_{p}(X,\beta^{\star})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(\delta\kappa(\theta_{0}),T_{1})$.
For any $m\in\mathbb{R}$, define a random variable $W_{m}:=N(m,d_{0})$ and a vector-valued function $\Phi:\mathbb{R}\to\mathbb{R}^{k}$ as $\Phi(m):=\operatorname{\mathbb{E}}h(W_{m})$. Then, we have
$\Phi^{\prime}(m)=\operatorname{\mathbb{E}}h^{\prime}(N(m,d_{0})),\quad\Phi^{\prime\prime}(m)=\operatorname{\mathbb{E}}h^{\prime\prime}(N(m,d_{0})).$ (52)
Here, note that for each $a\in[k]$, we have $\Phi^{\prime}_{a}\in\operatorname{\mathsf{{Poly}}}(3r+3),\Phi^{\prime\prime}_{a}\in\operatorname{\mathsf{{Poly}}}(3r+6),\Phi^{\prime\prime\prime}_{a}\in\operatorname{\mathsf{{Poly}}}(3r+9)$. To see this, by using (49) followed by Lemma A.5 (conditional on $W_{m}$) and Gaussian moment bounds, we have
$|\Phi_{a}^{\prime}(m)|=|\operatorname{\mathbb{E}}h_{a}^{\prime}(W_{m})|\leq\operatorname{\mathbb{E}}|h_{a}^{\prime}(W_{m})|\lesssim_{r,d_{0}}1+\operatorname{\mathbb{E}}|W_{m}|^{3(r+1)}\lesssim_{r,d_{0}}1+|m|^{3(r+1)}.$
The bound for $\Phi^{\prime\prime}_{a}$ follows from similar computations.
Now, by sequentially applying Remark B.1 and Lemma B.4 for each coordinate of $h$ (with $\Phi$ defined as above), we have
$\mu_{p}(X,\beta^{\star})$ $=\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}\Big[h(w_{i})\mid X,\beta^{\star}\Big]=\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi(d_{0}\beta_{i}^{\star}-\mathcal{F}_{i})+\mathnormal{O}_{\mathnormal{P},X}\Big(\sqrt{\frac{p}{n}}\Big)$ (53)
$=\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi(d_{0}\beta_{i}^{\star})+\frac{p^{3/2}}{n}\kappa(\theta_{0})+O_{P,X}\Big(\sqrt{\frac{p}{n}}+\frac{p^{2}}{n^{3/2}}\Big).$
The limit for $\mu_{p}(\beta^{\star})$ follows by taking the limit $\frac{p^{3/2}}{n}\to\delta$, and using the usual i.i.d. CLT for $\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi(d_{0}\beta_{i}^{\star})$ alongside the moment computations:
$\operatorname{\mathbb{E}}_{\beta_{i}^{\star}\sim B_{\theta_{0}}}\Phi(d_{0}\beta_{i}^{\star})=\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{\theta_{0}})=0,\quad\operatorname{\textnormal{Var}}_{\beta_{i}^{\star}\sim B_{\theta_{0}}}\big[\Phi(d_{0}\beta_{i}^{\star})\big]=\operatorname{\textnormal{Var}}\Big[\operatorname{\mathbb{E}}\big[h(W_{\theta_{0}})\mid B_{\theta_{0}}\big]\Big]=T_{1}.$
Step 2-2: $\tau^{2}_{p}(\beta^{\star})\operatorname*{\xrightarrow{\mathnormal{P}}}T_{2}$.
By expanding the variance as
$\operatorname{\textnormal{Var}}\Big(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}h(w_{i})\mid X,\beta^{\star}\Big)=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\textnormal{Var}}(h(w_{i})\mid X,\beta^{\star})+\sum_{i\neq j}\operatorname{\textnormal{Cov}}(h(w_{i}),h(w_{j})\mid X,\beta^{\star}),$
it suffices to separate show the following:
$\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\textnormal{Var}}(h(w_{i}))=\underbrace{\operatorname{\mathbb{E}}\Big[\operatorname{\textnormal{Var}}\big[h(W_{\theta_{0}})\mid B_{\theta_{0}}\big]\Big]}_{=T_{2}}+o_{P,X}(1),\quad\frac{1}{p}\sum_{1\mathnormal{\leq i\neq j\leq p}}\operatorname{\textnormal{Cov}}(h(w_{i}),h(w_{j}))=o_{P,X}(1).$
In the remainder of this step, we omit the conditioning on $X,\beta^{\star}$ for simplicity.
To show the first claim, define a function $\tilde{\Phi}(m):=\operatorname{\textnormal{Var}}[h(N(m,d_{0}))]$. By similar computations as in (52), the first and second derivatives of $\tilde{\Phi}$ are bounded by polynomials of appropriate degree. Hence, we can apply Remark B.1 and Lemma B.4 to write
$\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\textnormal{Var}}(h(w_{i}))$ $=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\tilde{\Phi}(d_{0}\beta_{i}^{\star}-\mathcal{F}_{i})+\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{1}{\sqrt{n}}\Big)$
$=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\tilde{\Phi}(d_{0}\beta_{i}^{\star})+\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{1}{\sqrt{n}}+\frac{p}{n}\Big)\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}\tilde{\Phi}(d_{0}\beta_{1}^{\star})=T_{2}.$
The final equality follows by plugging-in the definition of $\tilde{\Phi}$.
To show the second claim, define $\tilde{w}_{j}^{i}:=w_{j}+\frac{A_{p}(i,j)}{d_{i}}w_{i}$ so that $\operatorname{\textnormal{Cov}}(\tilde{w}_{j}^{i},w_{i})=0$ and $\tilde{w}_{j}^{i}$ and $w_{i}$ are independent. For any $i\neq j$, a Taylor expansion of $h(w_{j})$ around $w_{j}\approx\tilde{w}_{j}^{i}$ gives
$\operatorname{\textnormal{Cov}}\big(h(w_{i}),h(w_{j})\big)$
$=$ $\operatorname{\textnormal{Cov}}\big(h(w_{i}),h(\tilde{w}_{j}^{i})\big)-\frac{A_{p}(i,j)}{d_{i}}\operatorname{\textnormal{Cov}}\big(h(w_{i}),w_{i}h^{\prime}(\tilde{w}_{j}^{i})\big)+\frac{A_{p}(i,j)^{2}}{2d_{i}^{2}}\operatorname{\textnormal{Cov}}\big(h(w_{i}),w_{i}^{2}h^{\prime\prime}(\xi_{i,j})\big)$
$=$ $-A_{p}(i,j)\operatorname{\mathbb{E}}h^{\prime}(w_{i})\operatorname{\mathbb{E}}h^{\prime}(\tilde{w}_{j}^{i})^{\top}+O\Big(A_{p}(i,j)^{2}\big(\operatorname{\mathbb{E}}\|h(w_{i})\|^{2}+\operatorname{\mathbb{E}}[w_{i}^{2}\|h^{\prime\prime}(\xi_{i,j})\|^{2}]\big)\Big)$
$=$ $-A_{p}(i,j)\operatorname{\mathbb{E}}h^{\prime}(w_{i})\operatorname{\mathbb{E}}h^{\prime}(w_{j})^{\top}+O\Big(A_{p}(i,j)^{2}\big(\operatorname{\mathbb{E}}[\|h(w_{i})\|^{2}+\|h^{\prime}(w_{i})\|^{2}]+\operatorname{\mathbb{E}}[w_{i}^{2}(\|h^{\prime\prime}(\xi_{i,j})\|^{2}$
$\quad+\|h^{\prime\prime}(\eta_{i,j})\|^{2})]\big)\Big)$
$=$ $-A_{p}(i,j)\operatorname{\mathbb{E}}h^{\prime}(w_{i})\operatorname{\mathbb{E}}h^{\prime}(w_{j})^{\top}+O\Big(A_{p}(i,j)^{2}\operatorname{\mathbb{E}}\big[1+\operatorname{\mathbb{E}}|w_{i}|^{r^{\prime}}+\operatorname{\mathbb{E}}|w_{j}|^{r^{\prime}}\big]\Big)$
for $\xi_{i,j},\eta_{i,j}\in(w_{j},\tilde{w}_{j}^{i})$ and some positive integer $r^{\prime}$ that depends on $r$. Here, the second equality follows from expanding the covariance and using Stein’s identity:
$\operatorname{\textnormal{Cov}}\big(h(w_{i}),w_{i}h^{\prime}(\tilde{w}_{j}^{i})\big)$ $=\operatorname{\mathbb{E}}\big[(h(w_{i})-\operatorname{\mathbb{E}}h(w_{i}))w_{i}\big]\operatorname{\mathbb{E}}h^{\prime}(\tilde{w}_{j}^{i})$
$=\operatorname{\mathbb{E}}\big[h(w_{i})(w_{i}-\operatorname{\mathbb{E}}w_{i})\big]\operatorname{\mathbb{E}}h^{\prime}(\tilde{w}_{j}^{i})=d_{i}\operatorname{\mathbb{E}}h^{\prime}(w_{i})\operatorname{\mathbb{E}}h^{\prime}(\tilde{w}_{j}^{i}),$
and using the basic covariance inequality $\operatorname{\textnormal{Cov}}(V,W)\leq\sqrt{\operatorname{\textnormal{Var}}V^{2}\operatorname{\textnormal{Var}}W^{2}}\leq\frac{\operatorname{\mathbb{E}}V^{2}+\operatorname{\mathbb{E}}W^{2}}{2}$ for each entry of the covariance matrix. The fourth line uses another Taylor expansion of $h^{\prime}$: $h^{\prime}(\tilde{w}_{j}^{i})=h^{\prime}(w_{j})+\frac{A_{p}(i,j)w_{i}}{d_{i}}h^{\prime\prime}(\eta_{i,j})$ followed by a Cauchy Schwartz. The final line simplifies the error term $\operatorname{\mathbb{E}}\|h^{\prime}(w_{i})\|^{2}$ using (49)–(50):
$\operatorname{\mathbb{E}}\|h^{\prime}(w_{i})\|^{2}=\sum_{a=1}^{k}\operatorname{\mathbb{E}}[h_{a}^{\prime}(w_{i})^{2}]\lesssim 1+\operatorname{\mathbb{E}}|w_{i}|^{6(r+1)}.$
The remaining error terms can be similarly bounded by $O\big(\operatorname{\mathbb{E}}|{w_{i}}|^{r^{\prime}}+\operatorname{\mathbb{E}}|{w_{j}}|^{r^{\prime}}\big)$ for a large enough $r^{\prime}$, using the naive bound $|A_{p}(i,j)|\leq|d_{i}|$ to write $|\xi_{i,j}|\leq|w_{j}|+\frac{|A_{p}(i,j)|}{d_{i}}|w_{i}|\leq|w_{j}|+|w_{i}|$.
Now, by summing up the long display above for $i\neq j$, we get
$\frac{1}{p}\sum_{1\mathnormal{\leq i\neq j\leq p}}\operatorname{\textnormal{Cov}}(h(w_{i}),h(w_{j}))$
$=$ $-\frac{1}{p}\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)\operatorname{\mathbb{E}}h^{\prime}(w_{i})\operatorname{\mathbb{E}}h^{\prime}(w_{j})^{\top}+\frac{1}{p}O\Big(\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)^{2}\operatorname{\mathbb{E}}|w_{i}|^{r^{\prime}}\Big)$
$=$ $\mathnormal{O}_{\mathnormal{P},X}\Big(\sqrt{\frac{p}{n}}\Big)+O_{P}\Big(\frac{p}{n}\Big)=o_{P,X}(1).$
The last line uses the operator norm bound of $A_{p}$ for the first term (see Lemma A.11), and uses Lemma A.7:
$\frac{1}{p}\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)^{2}\operatorname{\mathbb{E}}|w_{i}|^{r^{\prime}}\leq\frac{\alpha_{p}}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}|w_{i}|^{r^{\prime}}=O_{P}(\alpha_{p})=O_{P}\Big(\frac{p}{n}\Big)$
to control the second term (recall the notation $\alpha_{p}=\max_{\mathnormal{i}=1}^{\mathnormal{p}}\sum_{j\neq i}A_{p}(i,j)^{2}$ from Lemma A.11).
Step 2-3: $T_{1}+T_{2}=V(\theta_{0})$.
Recalling 3.1, define a pair of random variables $(B_{\theta_{0}},W_{\theta_{0}})$. By recalling the definitions of $T_{1},T_{2}$ in (51), and applying the law of total variance, we have
$T_{1}+T_{2}$ $=\operatorname{\textnormal{Var}}\Big[\operatorname{\mathbb{E}}\big[h(W_{\theta_{0}})\mid B_{\theta_{0}}\big]\Big]+\operatorname{\mathbb{E}}\Big[\operatorname{\textnormal{Var}}\big[h(W_{\theta_{0}})\mid B_{\theta_{0}}\big]\Big]$
$=\operatorname{\textnormal{Var}}\big[h(W_{\theta_{0}})\big]$
$=\operatorname{\textnormal{Var}}\Big[\operatorname{\mathbb{E}}\big[\nabla\ell(\theta_{0};B_{W_{\theta_{0}},d_{0}})\mid W_{\theta_{0}}\big]\Big]$
$=\operatorname{\textnormal{Var}}[\nabla\ell(B_{\theta_{0}})]-\operatorname{\mathbb{E}}\Big[\operatorname{\textnormal{Var}}\big[\nabla\ell(\theta_{0};B_{W_{\theta_{0}},d_{0}})\mid W_{\theta_{0}}\big]\Big]=V(\theta_{0}).$
The third line plugs-in the definition of $h$. The last line follows from noting that $B_{\theta_{0}}\equiv B_{W_{\theta_{0}},d_{0}}$, and again applying the law of total variance. This completes the proof of part (a).
(b) Now, we work under the scaling $n^{2/3}\ll p\ll n$. Note that steps 1 and 2-2 above only require $p\ll n$, and still holds. Consequently by Slutsky, it only suffices to modify step 2-1 as follows, under the modified asymptotic scaling:
$\frac{n}{p\sqrt{p}}\mu_{p}(X,\beta^{\star})\operatorname*{\xrightarrow{\mathnormal{P}}}\kappa(\theta_{0}).$
Note that this limit has a slower convergence rate compared to the limit in step 1, and dominates the law of $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})$.
To wit, recall the notation $\Phi(m):=\operatorname{\mathbb{E}}[h(W_{m})]$ from (52) and recall that rescaling (53) by $\frac{n}{p\sqrt{p}}$ gives
$\frac{n}{p\sqrt{p}}\mu_{p}(X,\beta^{\star})$ $=\frac{n}{p^{2}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi(d_{0}\beta_{i}^{\star})+\kappa(\theta_{0})+O_{P}\Big(\frac{\sqrt{n}}{p}+\sqrt{\frac{p}{n}}\Big)=\kappa(\theta_{0})+o_{P}(1).$
Here, the last line used the fact that $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi(d_{0}\beta_{i}^{\star})=O_{P}(\sqrt{p})$ and the asymptotic scaling $p\gg n^{2/3}$. This completes the proof.
(c) Finally, suppose $p\ll n^{3/4}$. Again, by the computations in part (a), it suffices to show:
$\mu_{p}(X,\beta^{\star})-\frac{p\sqrt{p}}{n}\kappa(\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,T_{2}).$
But this is a direct consequence of (53), noting that the error term there is $o_{P,X}(1)$. ∎
Proof of Lemma A.5.
For simplicity, define the c.g.f. $\psi_{d,\theta}(t):=\log\int e^{bt-\frac{db^{2}}{2}}d\mu_{\theta}(b).$ Using the exponential tilt properties and the assumption $\|B_{t,d,\theta}\|_{\psi_{2}}\lesssim 1+|t|$, we have
$\psi_{d,\theta}^{\prime\prime}(t)=\operatorname{\textnormal{Var}}(B_{t,d,\theta})\lesssim 1+t^{2},\quad\forall t\in\mathbb{R}.$
Hence, by integrating over $t$, we get
$\operatorname{\mathbb{E}}B_{t,d,\theta}=\psi_{d,\theta}^{\prime}(t)\leq\operatorname{\mathbb{E}}B_{0,d,\theta}+O(|t|+|t|^{3})=O(1+|t|^{3})$ (54)
and the second claim holds. The simplification in the last equality used the assumption that $\operatorname{\mathbb{E}}B_{0,d,\theta}$ is uniformly bounded. Then, the first claim also follows as
$\psi_{d,\theta}(t)\leq\psi_{d,\theta}(0)+O(|t|^{3}+|t|^{4})=O(|t|^{3}+|t|^{4}).$
The last equality uses the fact $\psi_{d,\theta}(0)=\log\int e^{-\frac{db^{2}}{2}}d\mu_{\theta}(b)\leq 0.$
It remains to show the claims regarding $B_{t,d,\theta}$. Again by the sub-Gaussianity of $B_{t,d,\theta}$ and (54), we have
$\operatorname{\mathbb{E}}|B_{t,d,\theta}|^{r}\lesssim_{r}(1+|t|)^{r}+[\operatorname{\mathbb{E}}B_{t,d,\theta}]^{r}\lesssim_{r}1+|t|^{3r}.$
Now, by taking $t=w_{i}\stackrel{{\scriptstyle d}}{{\equiv}}N(d_{i}\beta_{i}^{\star}-\mathcal{F}_{i},d_{i})$ and using the tower property, we can further write
$\operatorname{\mathbb{E}}\big[|B_{w_{i},d,\theta}|^{r}\mid X,\beta^{\star}\big]\lesssim_{r}1+\operatorname{\mathbb{E}}\big[|w_{i}|^{3r}\mid X,\beta^{\star}\big]\lesssim_{r,d_{0}}1+|\beta_{i}^{\star}|^{3r}+|\mathcal{F}_{i}|^{3r},\quad\forall d,\theta.$
Hence, using Lemma A.7 and the assumption R5 that $\beta_{i}^{\star}$ has finite moments, we have $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}|B_{w_{i},d,\theta}|^{r}=O_{P}(p).$
Proof of Lemma A.6.
Set $G(d):=\log\int e^{tb-\frac{db^{2}}{2}}d\mu_{\theta}(b)$ and recall the random variable $B_{t,d,\theta}$ from 3.1. Then, we have $G^{\prime}(d)=-(1/2)\operatorname{\mathbb{E}}[B^{2}_{t,d,\theta}],$ so Taylor expansion gives
$|G(d)-G(d_{0})|=|d-d_{0}||G^{\prime}(\xi)|\leq|d-d_{0}|(1+|t|^{6}),\quad\exists\xi\in(d,d_{0}).$
The last bound used Lemma A.5 with $r=2$.
To show the second claim, for some $h\in\operatorname{\mathsf{{Poly}}}(r)$, set $H(d):=\operatorname{\mathbb{E}}[h(B_{t,d,\theta})]$. Then, by a similar Taylor expansion noting that $|H^{\prime}(d)|=\Big|\operatorname{\textnormal{Cov}}\Big(h(B_{t,d,\theta}),-\frac{B_{t,d,\theta}^{2}}{2}\Big)\Big|\lesssim\operatorname{\mathbb{E}}|B_{t,d,\theta}|^{r+2}$ and using Lemma A.5, we have
$|H(d)-H(d_{0})|=|d-d_{0}||H^{\prime}(\xi)|\leq|d-d_{0}|\big(1+|t|^{3(r+2)}\big).$
We prove Lemma A.7 using an additional concentration result for $\mathcal{F}_{i}=\sum_{j\neq i}A_{p}(i,j)\beta_{j}^{\star}$, which is proved in Appendix C.
Lemma B.6.
There exists some constant $C>0$ such that $\max_{i=1}^{p}\operatorname{\mathbb{E}}\big[e^{\lambda\mathcal{F}_{i}}\mid\beta^{\star}\big]\leq e^{\frac{Cp\lambda^{2}}{n}}$ for all $|\lambda|\leq\frac{Cn}{\sqrt{p}}$ with high probability (with respect to $\beta^{\star}$).
Proof of Lemma A.7.
We start by showing the first claim. Fix any constant $K>0$. It suffices to show that
$\operatorname{\mathbb{E}}\left(\frac{1}{p}\sum_{i=1}^{p}e^{K|\mathcal{F}_{i}|}\mid\beta^{\star}\right)\operatorname*{\xrightarrow{\mathnormal{P}}}1\qquad\mbox{and}\qquad\operatorname{\mathbb{E}}\left(\frac{1}{p}\sum_{i=1}^{p}e^{K|\mathcal{F}_{i}|}\mid\beta^{\star}\right)^{2}\operatorname*{\xrightarrow{\mathnormal{P}}}1.$ (55)
We will only prove the first display of (55) as the second display follows similarly. Note that $e^{|x|}\leq e^{x}+e^{-x}$. We then have
$1\leq\limsup\frac{1}{p}\operatorname{\mathbb{E}}\left[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}e^{K|\mathcal{F}_{i}|}\mid\beta^{\star}\right]\leq\limsup\max_{i=1}^{p}\operatorname{\mathbb{E}}\exp(K|\mathcal{F}_{i}|)\leq 1,$
with high probability, as a consequence of Lemma B.6. The conclusion follows from a standard sandwiching argument.
We move on to the second claim. Note that Lemma B.6 implies that $\sum_{i=1}^{p}|\mathcal{F}_{i}|^{r}=O_{P}(p)$. Using the Gaussianity of $w_{i}\mid X,\beta^{\star}$ yields
$\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}|w_{i}|^{r}\lesssim_{r}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[|d_{i}\beta_{i}^{\star}-\mathcal{F}_{i}|^{r}+|d_{i}|^{\frac{r}{2}}\Big]\lesssim_{r}p+\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[|d_{i}|^{r}(1+|\beta_{i}^{\star}|^{r})+|\mathcal{F}_{i}|^{r}\Big]=O_{P}(p).$
This completes the proof. ∎
B.3 Proof of lower bound lemmas
This Section is devoted to the proof of Lemma A.8. Throughout the following proof, $B=(B_{1},\ldots,B_{p})$ will denote a sample from the RFIM $\tilde{\operatorname{\mathbb{P}}}=\tilde{\operatorname{\mathbb{P}}}_{\theta}^{A_{p},c}$ (see A.1). The proof proceeds through the use of Stein’s method of exchangeable pairs which has served as a key tool for analyzing Ising models (see e.g. [16, 15, 12] and the references therein).
Proof of Lemma A.8.
We start by deriving general results that will be used to prove both parts of the statement. Let $g:[-1,1]\to[-1,1]$ be a continuous function. Define
$T_{p}=T_{p}(B):=\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big(g(B_{i})-\langle g\rangle_{\mu_{i,s_{i}+c_{i},\theta}}\Big).$
Given an RFIM sample $B\sim\tilde{\operatorname{\mathbb{P}}}$, generate $B^{\prime}$ as follows: let $I$ be a randomly sampled index from $\{1,2,\ldots,p\}$. Given $I=i$, replace $B_{i}$ with an independently generated $B^{\prime}_{i}$ drawn from the conditional distribution of $B_{i}$ given $(B_{j},\ j\neq i)$.
Recalling the notation $m_{i}$ from Lemma A.10, we can write
$B_{i}\mid(B_{j},\ j\neq i)\stackrel{{\scriptstyle d}}{{\equiv}}\mu_{i,m_{i}+c_{i},\theta}.$ (56)
For notational convenience, define
$h_{i}(t):=\langle g\rangle_{\mu_{i,t,\theta}}=\operatorname{\mathbb{E}}_{B_{i}\sim\mu_{i,t,\theta}}g(B_{i}),\quad\forall t\in\mathbb{R},$
and note that (56) implies $\operatorname{\mathbb{E}}[g(B_{i})\mid B_{j}:j\neq i]=h_{i}(m_{i}+c_{i}).$ Also, observe that
$h_{i}^{\prime}(t)=\operatorname{\textnormal{Cov}}_{\mu_{i,t,\theta}}(B_{i},g(B_{i})),\quad h_{i}^{\prime\prime}(t)=\operatorname{\textnormal{Cov}}_{\mu_{i,t,\theta}}\Big((B_{i}-\psi_{i,\theta}^{\prime}(t))^{2},g(B_{i})\Big)$ (57)
are uniformly bounded. By a Taylor’s series expansion of around $m_{i}\approx s_{i}$, we get that for some $\{\xi_{i}\}_{1\leq i\leq p}$:
$\operatorname{\mathbb{E}}(T_{p}-T_{p}^{\prime}$ $\mid B)=\frac{1}{p\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(g(B_{i})-h_{i}(m_{i}+c_{i}))$
$=\frac{1}{p\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\left(g(B_{i})-h_{i}(s_{i}+c_{i})-(m_{i}-s_{i})h_{i}^{\prime}(s_{i}+c_{i})-\frac{1}{2}(m_{i}-s_{i})^{2}h_{i}^{\prime\prime}(\xi_{i}+c_{i})\right)$
$={\frac{T_{p}}{p}}-\underbrace{\frac{1}{p\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(m_{i}-s_{i})h_{i}^{\prime}(s_{i}+c_{i})}_{:=H_{1}}-\underbrace{\frac{1}{2p\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(m_{i}-s_{i})^{2}h_{i}^{\prime\prime}(\xi_{i}+c_{i})}_{:=H_{2}}.$ (58)
We bound the error terms $H_{1},H_{2}$ in the above expression.
  • For $H_{1}$, we expand $m_{i}-s_{i}=\sum_{j\neq i}A_{p}(i,j)(B_{j}-u_{j})$ and write
    $H_{1}=\frac{1}{p\sqrt{p}}\sum_{\mathnormal{j}=1}^{\mathnormal{p}}\gamma_{j}(B_{j}-u_{j}),$
    where $\gamma_{j}:=\sum_{i\neq j}A_{p}(i,j)h_{i}^{\prime}(s_{i}+c_{i})$. Then, by part (a) of Lemma A.10, we have
    $p^{2}\operatorname{\mathbb{E}}H_{1}^{2}\lesssim\frac{1}{p}\big(\sum_{\mathnormal{j}=1}^{\mathnormal{p}}\gamma_{j}^{2}\big)\Big(1+\frac{p^{3}}{n^{2}}\Big)=\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{p}{n}\Big).$
    The final inequality follows from using the operator norm bound for $A_{p}$ in Lemma A.11 to write $\sum_{\mathnormal{j}=1}^{\mathnormal{p}}\gamma_{j}^{2}\lesssim p\|A_{p}\|^{2}=\mathnormal{O}_{\mathnormal{P},X}(p^{2}/n)$.
  • For $H_{2}$, using the fact that $h_{i}^{\prime\prime}$ are uniformly bounded followed by part (b) of Lemma A.10, we have
    $p^{2}\operatorname{\mathbb{E}}H_{2}^{2}\lesssim\frac{1}{p}\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(m_{i}-s_{i})^{2}\Big]^{2}=\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{p^{3}}{n^{2}}\Big).$
Now, we are ready to prove each claims.
(a) We prove the first moment bound. Since $(T_{p},T_{p}^{\prime})$ form an exchangable pair, we have $\operatorname{\mathbb{E}}T_{p}=\operatorname{\mathbb{E}}T_{p}^{\prime}$. Hence, taking an outer expectation on (58) gives the identity
$\operatorname{\mathbb{E}}T_{p}=p\operatorname{\mathbb{E}}H_{1}+p\operatorname{\mathbb{E}}H_{2}.$
The earlier error bounds for $H_{1},H_{2}$ implies that the RHS is $\mathnormal{O}_{\mathnormal{P},X}\left(\sqrt{\frac{p}{n}}+\sqrt{\frac{p^{3}}{n^{2}}}\right)$.
To prove the second conclusion, using a Taylor expansion followed by the operator norm bound of $A_{p}$ (see Lemma A.11) gives
$\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[h_{i}(s_{i}+w_{i})-h_{i}(w_{i})\Big]$ $=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[s_{i}h_{i}^{\prime}(\xi_{i}+w_{i})\Big]\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)s_{i}h_{i}^{\prime}(\xi_{i}+w_{i})u_{j}=\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{p\sqrt{p}}{\sqrt{n}}\Big).$ (59)
Here, we have used the fact that $h_{i}^{\prime}$ is uniformly bounded, which follows from the compact support assumption.
(b) Next, we prove the second moment bound. Let $\Delta_{p}:=T_{p}-T_{p}^{\prime}=\frac{1}{\sqrt{p}}(g(B_{I})-g(B_{I}^{\prime}))$ and note that $|\Delta_{p}|\leq\frac{2}{\sqrt{p}}$. Multiplying (58) by $pT_{p}$ and taking an outer expectation gives
$p\operatorname{\mathbb{E}}[\Delta_{p}T_{p}]=\operatorname{\mathbb{E}}[T_{p}^{2}]-(p\operatorname{\mathbb{E}}[T_{p}H_{1}]+p\operatorname{\mathbb{E}}[T_{p}H_{2}]).$
On the other hand, by exchangeability of $(T_{p},T_{p}^{\prime})$, we can also write
$p\operatorname{\mathbb{E}}[\Delta_{p}T_{p}]=\frac{p\operatorname{\mathbb{E}}[\Delta_{p}^{2}]}{2}\leq 2.$
Hence, by upper bounding
$|p\operatorname{\mathbb{E}}[T_{p}H_{a}]|\leq\sqrt{p^{2}\operatorname{\mathbb{E}}H_{a}^{2}}\sqrt{\operatorname{\mathbb{E}}[T_{p}^{2}]},\quad a=1,2,$ (60)
we get the inequality
$\operatorname{\mathbb{E}}T_{p}^{2}\leq 2+\sum_{a=1}^{2}\sqrt{p^{2}\operatorname{\mathbb{E}}H_{a}^{2}}\sqrt{\operatorname{\mathbb{E}}[T_{p}^{2}]}.$
By plugging-in the earlier bounds for $p^{2}\operatorname{\mathbb{E}}H_{a}^{2}$, we get a preliminary bound $\operatorname{\mathbb{E}}[T_{p}^{2}]=\mathnormal{O}_{\mathnormal{P},X}(1)$.
To compute the precise constant for $\operatorname{\mathbb{E}}[T_{p}^{2}]$, we expand the term $p\operatorname{\mathbb{E}}[\Delta_{p}^{2}]$ as follows:
$p\operatorname{\mathbb{E}}[\Delta_{p}^{2}]$ $=\frac{1}{p}\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}[g(B_{i})-g(B_{i}^{\prime})]^{2}\Big]$
$=\frac{2}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[\operatorname{\mathbb{E}}g(B_{i})^{2}-\operatorname{\mathbb{E}}[g(B_{i})\langle g\rangle_{\mu_{i,m_{i}+c_{i},\theta}}]\Big],$
where the last line follows by the tower property (by conditioning on $(B_{j}:j\neq i)$ and using (56)) and using exchangeability. Now, using the two conclusions of part (c) in Lemma A.10 (with $h\equiv g^{2},g$, respectively) gives the following moment bounds:
$\operatorname{\mathbb{E}}\Big[\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(B_{i})^{2}\Big]$ $=\operatorname{\mathbb{E}}\Big[\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\langle g^{2}\rangle_{\mu_{i,m_{i}+c_{i},\theta}}\Big]+O\Big(\frac{1}{\sqrt{p}}\Big),$
$\operatorname{\mathbb{E}}\Big[\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(B_{i})\langle g\rangle_{\mu_{i,\theta}(m_{i}+c_{i})}\Big]$ $=\operatorname{\mathbb{E}}\Big[\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big(\langle g\rangle_{\mu_{i,m_{i}+c_{i},\theta}}\big)^{2}\Big]+O\Big(\frac{1}{\sqrt{p}}\Big).$
We additionally simplify the RHS expressions using the following fact: for any continuous function $h:[-1,1]\to[-1,1]$, we have
$\operatorname{\mathbb{E}}\Big[\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}h(m_{i}+c_{i})\Big]=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}h(c_{i})+\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|m_{i}|=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}h(c_{i})+\mathnormal{O}_{\mathnormal{P},X}\Big({\sqrt{\frac{p}{n}}}\Big),$
which follows from bounding $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|m_{i}|\leq\sqrt{p}\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}m_{i}^{2}}=\mathnormal{O}_{\mathnormal{P},X}\Big(\sqrt{\frac{p^{3}}{n}}\Big).$ By taking $h$ appropriately, we can further simplify the individual moment bounds as:
$\operatorname{\mathbb{E}}\Big[\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(B_{i})^{2}\Big]$ $=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\langle g^{2}\rangle_{\mu_{i,c_{i},\theta}}+\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{1}{\sqrt{p}}+\sqrt{\frac{p}{n}}\Big),$
$\operatorname{\mathbb{E}}\Big[\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(B_{i})\langle g\rangle_{\mu_{i,\theta}(m_{i}+c_{i})}\Big]$ $=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big(\langle g\rangle_{\mu_{i,c_{i},\theta}}\big)^{2}+\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{1}{\sqrt{p}}+\sqrt{\frac{p}{n}}\Big).$
By combining all computations above, we have the following:
$\operatorname{\mathbb{E}}[T_{p}^{2}]=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[\langle g^{2}\rangle_{\mu_{i,c_{i},\theta}}-\big(\langle g\rangle_{\mu_{i,c_{i},\theta}}\big)^{2}\Big]+(p\operatorname{\mathbb{E}}[T_{p}H_{1}]+p\operatorname{\mathbb{E}}[T_{p}H_{2}])+\mathnormal{O}_{\mathnormal{P},X}\Big({\frac{1}{\sqrt{p}}+\sqrt{\frac{p}{n}}}\Big).$
Using (60) along the preliminary bound $\operatorname{\mathbb{E}}T_{p}^{2}=O(1)$, we have
$|p\operatorname{\mathbb{E}}[T_{p}H_{1}]+p\operatorname{\mathbb{E}}[T_{p}H_{2}]|\leq\sqrt{\operatorname{\mathbb{E}}T_{p}^{2}}\Big(\sqrt{p^{2}\operatorname{\mathbb{E}}H_{1}^{2}}+\sqrt{p^{2}\operatorname{\mathbb{E}}H_{2}^{2}}\Big)=\mathnormal{O}_{\mathnormal{P},X}\Big({\sqrt{\frac{p}{n}}+\sqrt{\frac{p^{3}}{n^{2}}}}\Big).$
This completes the proof. Note that the term $\sqrt{p/n}$ is omitted in the final bound since
$\frac{1}{\sqrt{p}}+\sqrt{\frac{p^{3}}{n^{2}}}\geq 2\sqrt{\frac{p}{n}}.$
B.4 Proof of lemmas for Section 4
Proof of Lemma A.9.
Fix $t$, an index $i\leq p$, a bounded function $h$ with $|h|\leq 1$, and write
$\langle h\rangle_{\mu_{i,t,\theta}}=\frac{\int h(b)e^{tb-\frac{d_{i}b^{2}}{2}+\ell(\theta;b)}db}{\int e^{tb-\frac{d_{i}b^{2}}{2}+\ell(\theta;b)}db}=\frac{A(\theta)}{B(\theta)}.$
Recall that we assume in 3.5 that $\ell(\theta;b)$ is continuous in both $\theta$ and $b$, which implies that it is Lipschitz in both arguments. Hence, for any function $g:[-1,1]\to[-1,1]$, we have
$\Big|\int g(b)e^{tb-\frac{d_{i}b^{2}}{2}+\ell(\theta;b)}(1-e^{\ell(\theta^{\prime};b)-\ell(\theta;b)})db\Big|$
$\leq$ $\sup_{b\in[-1,1]}\Big|1-e^{\ell(\theta^{\prime};b)-\ell(\theta;b)}\Big|\int e^{tb-\frac{d_{i}b^{2}}{2}+\ell(\theta;b)}db\lesssim\|\theta-\theta^{\prime}\|B(\theta).$
Here, the hidden constant only depends on the log-likelihood $\ell$. By taking $g(b)=h(b)$ and $1$ respectively, we have the following bounds:
$|A(\theta)-A(\theta^{\prime})|\lesssim\|\theta-\theta^{\prime}\|B(\theta),\quad|B(\theta)-B(\theta^{\prime})|\lesssim\|\theta-\theta^{\prime}\|B(\theta^{\prime}).$
Now, the claim follows by using the triangle inequality and plugging in the above bounds:
$|\langle h\rangle_{\mu_{i,t,\theta}}-\langle h\rangle_{\mu_{i,t,\theta^{\prime}}}|=\left|\frac{A(\theta)}{B(\theta)}-\frac{A(\theta^{\prime})}{B(\theta^{\prime})}\right|\leq\frac{|B(\theta)-B(\theta^{\prime})|A(\theta)}{B(\theta)B(\theta^{\prime})}+\frac{|A(\theta)-A(\theta^{\prime})|}{B(\theta)}\lesssim\|\theta-\theta^{\prime}\|.$
Note that we have also used the basic identity $|\frac{A(\theta)}{B(\theta)}|\leq 1$ for the final simplification. ∎
To prove Lemma A.13, we use the following fact from the proof of Theorem 3.1, part (a) in [33] (see the last two bullet points there).
Lemma B.7.
Take $\epsilon:=A_{p}q$. For any $\theta$, set
$R_{4}(\theta):=\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)q_{i}\big(\psi_{i,\theta}^{\prime\prime}(w_{i})-\upsilon_{p})\psi_{j,\theta}^{\prime}(w_{j})\big).$
Then, fixing $\theta_{0}$, the following bounds hold:
$\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\epsilon_{i}\psi_{i,\theta_{0}}^{\prime}(w_{i})=o_{P}(1),\quad R_{4}(\theta_{0})=o_{P}(1).$
Proof of Lemma A.13.
By the triangle inequality, it suffices to separately prove the following:
$\Big|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\big(\hat{u}_{i}-\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})\big)\Big|=o_{P}(1),\quad\Big|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\big(\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})-\psi_{0,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})\big)\Big|=o_{P}(1).$ (61)
For simplicity, we use the notations $R_{1},R_{2}$ from Lemma A.12 and also set
$R_{3}:=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[\sum_{\mathnormal{j}=1}^{\mathnormal{p}}A_{p}(i,j)\psi_{j,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{j})\Big]^{4}.$
To show the first bound in (61), apply Theorem 2.5(a) of [34] with $\lambda_{n}\equiv 0$ to write
$\Big|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\hat{u}_{i}-\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\psi^{\prime}_{i,\hat{\theta}_{\mathnormal{p}}}(w_{i})\Big|$ $\lesssim\sqrt{R_{1}R_{2}}+\sqrt{R_{3}}+\sqrt{R_{2}\|A_{p}q\|}+R_{4}(\hat{\theta}_{\mathnormal{p}})+\big|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\epsilon_{i}\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})\big|$
$\lesssim R_{4}(\hat{\theta}_{\mathnormal{p}})+\big|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\epsilon_{i}\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})\big|+o_{P}(1).$ (62)
The simplification in the second line follow from the bounds for $R_{1},R_{2}$ in (39) (and eq. (A.17) in [33] to control $R_{3}$) to write
$R_{1}R_{2}=O_{P,X}\Big(\frac{p^{3}}{n^{2}}\Big),\quad R_{3}=o_{P}(1),\quad R_{2}\|A_{p}q\|=O_{P,X}\Big(\frac{p^{3}}{n^{2}}\Big).$
We show that the first and second term in (B.4) are both $o_{P}(1)$. Our main idea is to write
$\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})=\psi_{i,\theta_{0}}^{\prime}(w_{i})+E_{i},\quad\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime\prime}(w_{i})=\psi_{i,\theta_{0}}^{\prime\prime}(w_{i})+F_{i},$
where the error terms are bounded as $|E_{i}|,|F_{i}|\lesssim\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|$ (by Lemma A.9).
We first claim that the third term in (B.4) is $o_{P}(1)$. By using the triangle inequality, we have
$|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\epsilon_{i}\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})|\leq|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\epsilon_{i}\psi_{i,\theta_{0}}^{\prime}(w_{i})|+\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|\epsilon_{i}||E_{i}|=o_{P}(1).$
In the equality, the first term is controlled by Lemma B.7. The second term is bounded via the assumption $\|\hat{\theta}_{\mathnormal{p}}-\theta_{0}\|=O_{P}(1/\sqrt{p})$ alongside $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|\epsilon_{i}|\leq\sqrt{p}\|A_{p}q\|=o_{P}(\sqrt{p^{2}/n}).$
For the middle term in (B.4), we have
$~~\big|\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)q_{i}(\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime\prime}(w_{i})-\hat{\upsilon}_{p})\psi_{j,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{j})\big|$
$\leq R_{4}(\theta_{0})+\big|\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)q_{i}(\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime\prime}(w_{i})-\hat{\upsilon}_{p})E_{j}\big|+\big|\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)q_{i}(E_{i}+\upsilon_{p}-\hat{\upsilon}_{p})\psi_{j,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{j})\big|$
$=o_{P}(1).$
In the last equality, the second term is bounded using the operator norm bound
$|\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)q_{i}(\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime\prime}(w_{i})-\hat{\upsilon}_{p})E_{j}|\leq\|A_{p}\|\sqrt{\sum_{\mathnormal{j}=1}^{\mathnormal{p}}E_{j}^{2}}=O_{P}\Big(\sqrt{\frac{p}{n}}\Big),$
and the third term also follows similarly by plugging in the size of $\hat{\upsilon}_{p}-\upsilon_{p}$ computed in part (a) of 4.3. The bound $R_{4}(\theta_{0})=o_{P}(1)$ follows from Lemma B.7. Hence, we have shown that the RHS of (B.4) is $o_{P}(1)$.
Finally, the second bound in (61) follows from using Lemma B.3 with $m=m_{0}=w_{i}$ and $d=d_{i}$ to get:
$\Big|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\big(\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})-\psi_{0,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})\big)\Big|\leq\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|q_{i}||d_{i}-d_{0}|\leq\|q\|\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{p}}(d_{i}-d_{0})^{2}}=\mathnormal{O}_{\mathnormal{P},X}\Big(\sqrt{\frac{p}{n}}\Big).$
This completes the proof. ∎
Appendix C Proof of technical concentration lemmas
Proof of Lemma B.1.
Let $\pi_{m}$ denote the probability of the $m$th mixture component. By the tower property and sub-Gaussianity of mixture components, we have
$\operatorname{\mathbb{E}}e^{\lambda X}\leq\sum_{m=1}^{M}\pi_{m}e^{\lambda\mu_{m}+\frac{\lambda^{2}K_{m}^{2}}{2}}\leq e^{\frac{\lambda^{2}K_{\max}^{2}}{2}}\sum_{m=1}^{M}\pi_{m}e^{\lambda\mu_{m}},\quad\forall\lambda\in\mathbb{R}.$
Let $Y$ be a categorical random variable with pmf $\operatorname{\mathbb{P}}(Y=\mu_{m})=\pi_{m}$ for all $m\leq M$. Then, by using Hoeffding’s inequality, we can write
$\sum_{m=1}^{M}\pi_{m}e^{\lambda\mu_{m}}=\operatorname{\mathbb{E}}e^{\lambda Y}\leq e^{\lambda\operatorname{\mathbb{E}}Y+\frac{\lambda^{2}(\mu_{\max}-\mu_{\min})^{2}}{8}}.$
Finally, noting that $\operatorname{\mathbb{E}}X=\operatorname{\mathbb{E}}Y$, we have
$\operatorname{\mathbb{E}}e^{\lambda(X-\operatorname{\mathbb{E}}X)}\leq\exp\Big[\frac{\lambda^{2}}{2}\Big(K_{\max}^{2}+\frac{1}{4}(\mu_{\max}-\mu_{\min})^{2}\Big)\Big]$
and the proof is complete. ∎
Proof of Lemma B.2.
Recall that $X_{p}\mid Y_{p}\operatorname*{\xrightarrow{\mathnormal{P}}}0$ if and only if $\operatorname{\mathbb{P}}(|X_{p}|>\epsilon\mid Y_{p})\operatorname*{\xrightarrow{\mathnormal{P}}}0$ for any $\epsilon>0$. The claim is immediate after using the tower property to write $\operatorname{\mathbb{P}}(|X_{p}|>\epsilon)=\operatorname{\mathbb{E}}\operatorname{\mathbb{P}}(|X_{p}|>\epsilon\mid Y_{p})$, and using the bounded convergence theorem (for the if part) / Markov inequality (only if part). ∎
Proof of Lemma B.3.
It suffices to prove the claim under two cases: (1) $m=m_{0}$, (2) $d=d_{0}$. First, suppose $m=m_{0}=0$. Noting that $\operatorname{\mathbb{E}}h(V)=\operatorname{\mathbb{E}}h(\sqrt{\frac{d}{d_{0}}}V_{0})$, we can write
$|\operatorname{\mathbb{E}}h(V)-\operatorname{\mathbb{E}}h(V_{0})|$ $=|\operatorname{\mathbb{E}}h(\sqrt{\frac{d}{d_{0}}}V_{0})-\operatorname{\mathbb{E}}h(V_{0})|$
$\lesssim\operatorname{\mathbb{E}}\Big|\big(\sqrt{\frac{d}{d_{0}}}-1\big)V_{0}\Big|(1+|V_{0}|^{r})$
$\lesssim_{r}\big|\sqrt{\frac{d}{d_{0}}}-1\big|(d_{0}^{1/2}+d_{0}^{(r+1)/2})\lesssim_{d_{0},r}|d-d_{0}|.$
Here, the second line follows from a Taylor expansion alongside the polynomial growth bound $|h^{\prime}(v)|\lesssim 1+|v|^{r}$, where $v\leq\max(V_{0},\sqrt{d/d_{0}}V_{0})\leq\sqrt{2}V_{0}$. The square root term in the final line is simplified using $\sqrt{\frac{d}{d_{0}}}-1=\frac{d-d_{0}}{\sqrt{d_{0}}(\sqrt{d}+\sqrt{d_{0}})}.$
Next, suppose $d=d_{0}=1$. Then, by similar computations using the bound
$\sup_{v\in[v_{0}+(m-m_{0}),v_{0}]}|h^{\prime}(v)|\lesssim\sup_{v\in[v_{0}+(m-m_{0}),v_{0}]}\big[1+|v|^{r}\big]\lesssim_{r}1+|v_{0}|^{r}+|m-m_{0}|^{r},$
we have
$|\operatorname{\mathbb{E}}h(V)-\operatorname{\mathbb{E}}h(V_{0})|$ $=|\operatorname{\mathbb{E}}h(V_{0}+(m-m_{0}))-\operatorname{\mathbb{E}}h(V_{0})|$
$\leq|m-m_{0}|(1+\operatorname{\mathbb{E}}|V_{0}|^{r}+|m-m_{0}|^{r})$
$\lesssim_{r}|m-m_{0}|(1+m_{0}^{r}+|m-m_{0}|^{r}).$
Finally, we prove Lemma B.4 using another technical moment bound as a separate lemma, which will be proved at the end of this appendix.
Lemma C.1.
Assuming $\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}(\beta_{1}^{\star})^{4}<\infty$ and $\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}g(\beta_{1}^{\star})^{2}<\infty$, it holds that
$\operatorname{\mathbb{E}}\Big[\frac{n}{p^{2}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})-\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}g(\beta_{1}^{\star})\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}(\beta_{1}^{\star})^{2}\Big]^{2}=O\Big(\frac{1}{p}+\frac{p}{n}\Big).$
Proof of Lemma B.4.
(a) Let $\Phi$ be any $\mathcal{C}^{3}$ function such that $|\Phi^{\prime}(m)|,|\Phi^{\prime\prime}(m)|,|\Phi^{\prime\prime\prime}(m)|=\operatorname{\mathsf{{Poly}}}(r)$ for some constant $r>0$. By a Taylor expansion, we can write
$\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi(d_{0}\beta_{i}^{\star}-\mathcal{F}_{i})-\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi(d_{0}\beta_{i}^{\star})$
$=$ $-\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}\Phi^{\prime}(d_{0}\beta_{i}^{\star})+\frac{1}{2}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}\Phi^{\prime\prime}(d_{0}\beta_{i}^{\star})-\frac{1}{6}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{3}\Phi^{\prime\prime\prime}(d_{0}\beta_{i}^{\star}+\xi_{i}),$ (63)
where $\xi_{i}\in(-\mathcal{F}_{i},0)$. We separately control the second moment of each term in the RHS.
We start by control the first term in (63). By computing the moment (c.f. Lemma A.9(a) in [33]), we have
$\operatorname{\mathbb{E}}_{X}\Big[\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}\Phi^{\prime}(d_{0}\beta_{i}^{\star})\Big]^{2}\mid\beta^{\star}\Big]\lesssim\frac{\|\beta^{\star}\|^{2}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi^{\prime}(d_{0}\beta_{i}^{\star})^{2}}{n}\lesssim\frac{\|\beta^{\star}\|^{2}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|\beta_{i}^{\star}|^{2r}}{n}=O_{P,\beta^{\star}}\Big(\frac{p^{2}}{n}\Big).$
The last bound used assumption R5 that $\beta_{i}^{\star}\stackrel{{\scriptstyle d}}{{\equiv}}B_{\theta_{0}}$ has finite moments. Hence, we have
$\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}\Phi^{\prime}(d_{0}\beta_{i}^{\star})=O_{P}\Big(\frac{p}{\sqrt{n}}\Big).$
Next, we move on to the second term in (63). By the moment bound in Lemma C.1, we can write
$\frac{1}{2}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}\Phi^{\prime\prime}(d_{0}\beta_{i}^{\star})=\frac{p^{2}}{n}\kappa_{\Phi}(\theta_{0})+O_{P}\Big(\frac{p^{3/2}}{n}+\frac{p^{5/2}}{n^{3/2}}\Big).$
Finally, we control the third term in (63). By Holder’s inequality (with exponents $4/3,4$), using the growth condition on $\Phi^{\prime\prime\prime}$, and recalling $|\xi_{i}|\leq|\mathcal{F}_{i}|$, we have
$\operatorname{\mathbb{E}}_{X}\Big|\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{3}\Phi^{\prime\prime\prime}(d_{0}\beta_{i}^{\star}+\xi_{i})\Big|\leq\Big[\operatorname{\mathbb{E}}_{X}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{4}\Big]^{3/4}\Big[\operatorname{\mathbb{E}}_{X}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big[|d_{0}\beta_{i}^{\star}|^{4r}+|\mathcal{F}_{i}|^{4r}\big]\Big]^{1/4}=O_{P,\beta^{\star}}\Big(\frac{p^{5/2}}{n^{3/2}}\Big).$
Here, the last bound follows as (i) $\operatorname{\mathbb{E}}_{X}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{4}\Big]=O_{P,\beta^{\star}}(p^{3}/n^{2})$ by Lemma A.9(b) in [33], and (ii) $\operatorname{\mathbb{E}}_{X}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}|d_{0}\beta_{i}^{\star}|^{4r}+|\mathcal{F}_{i}|^{4r}\Big]=O_{P,\beta^{\star}}(p)$ by the moment assumption for $\beta_{i}^{\star}$ and the moment bounds for $\mathcal{F}_{i}$ in Lemma B.6.
The final bound follows by combining the above bounds and noting that
$\frac{p}{\sqrt{n}}+\frac{p^{5/2}}{p^{3/2}}\gg\frac{p^{3/2}}{n}.$
Note that the final form in Lemma B.4, which is conditional on $X$, follows from Lemma B.2. ∎
Proof of Lemma B.6.
Throughout the proof, all expectations will be over $X$, and always be conditional on $\beta^{\star}$. Fix any $i\leq p$. For $k\leq n$, let $Z_{k,i}:=\sqrt{n}X_{k,i}$ and recall from 3.1 that $Z_{k,i}$ are mean zero sub-Gaussian random variables with sub-Gaussian norm bounded by $K>0$. Let us assume By definition, we can write $\mathcal{F}_{i}=\frac{1}{n}\sum_{k=1}^{n}\sum_{1\leq j\neq i\leq p}Z_{k,i}Z_{k,j}\beta_{j}^{\star}$. Setting $Y_{k,i}:=\sum_{1\leq j\neq i\leq p}Z_{k,j}\beta_{j}^{\star}$, $Y_{k,i}$ is a sub-Gaussian random variable whose norm is bounded by $K\|\beta^{\star}\|$ by Hoeffding’s inequality.
Now, the product property for sub-Gaussians (e.g. Lemma 2.8.6 in [58]) implies that $Z_{k,i}Y_{k}$ is sub-exponential with the corresponding sub-exponential norm bounded by $K^{2}\|\beta^{\star}\|$. This can be equivalently written out as
$\operatorname{\mathbb{E}}e^{\lambda^{\prime}Z_{k,i}Y_{k}}\leq e^{K^{4}\|\beta^{\star}\|^{2}\lambda^{\prime 2}},\quad\forall\,\,|\lambda^{\prime}|\leq\frac{1}{K^{2}\|\beta^{\star}\|}.$
Now, by spelling out $\mathcal{F}_{i}=\frac{1}{n}\sum_{k=1}^{n}Z_{k,i}Y_{k,i}$ and setting $\lambda=\lambda^{\prime}n$, we have
$\operatorname{\mathbb{E}}e^{\lambda\mathcal{F}_{i}}=\prod_{i=1}^{n}\operatorname{\mathbb{E}}e^{\frac{\lambda}{n}Z_{k,i}Y_{k}}\leq e^{K^{4}\|\beta^{\star}\|^{2}\lambda^{2}/n}$
as long as $|\lambda|\leq\frac{n}{K^{2}\|\beta^{\star}\|}$. The final conclusion follows by noting that $\|\beta^{\star}\|=\mathnormal{O}_{\mathnormal{P},\beta^{\star}}(\sqrt{p}).$
Proof of Lemma C.1.
For notational convenience, define a $n\times p$ matrix $Z$ with entries $Z_{k,i}:=\sqrt{n}X_{k,i}$, so that each $Z_{k,i}$ are independent sub-Gaussians with mean 0 and variance 1. We will write $\operatorname{\mathbb{E}}_{\operatorname{Z}}$ and $\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}$ for expectations with respect to the matrix $Z$ and $\operatorname{\boldsymbol{\beta}^{\star}}$. Observe that
$\operatorname{\mathbb{E}}_{\operatorname{Z}}\mathcal{F}_{i}^{2}=\operatorname{\mathbb{E}}_{\operatorname{Z}}\left[\frac{1}{n^{2}}\sum_{k_{1},k_{2}}\sum_{j_{1},j_{2}\neq i}Z_{k_{1},i}Z_{k_{1},j_{1}}Z_{k_{2},i}Z_{k_{2},j_{2}}\beta_{j_{1}}^{\star}\beta_{j_{2}}^{\star}\right]=\frac{1}{n}\sum_{j\neq i}(\beta_{j}^{\star})^{2}.$ (64)
By using (64), we get:
$\frac{n}{p^{2}}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}\operatorname{\mathbb{E}}_{\operatorname{Z}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})$ $=\frac{1}{p^{2}}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}\sum_{i\neq j}(\beta_{j}^{\star})^{2}g(\beta_{i}^{\star})$
$=\frac{p-1}{p}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}(\beta_{1}^{\star})^{2}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}g(\beta_{1}^{\star})\to\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}(\beta_{1}^{\star})^{2}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}g(\beta_{1}^{\star}).$
Having established the limit for the mean, we proceed to show the variance converges to 0. Using the law of total variance, we will separately show that
$\frac{n^{2}}{p^{4}}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}\operatorname{\textnormal{Var}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\boldsymbol{\beta}^{\star}}\right)\to 0\quad\mbox{and}\quad\frac{n^{2}}{p^{4}}\operatorname{\textnormal{Var}}\operatorname{\mathbb{E}}_{\operatorname{Z}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\boldsymbol{\beta}^{\star}}\right)\to 0.$ (65)
By (64), we have:
$\;\;\;\;\frac{n^{2}}{p^{4}}\operatorname{\textnormal{Var}}\operatorname{\mathbb{E}}_{\operatorname{Z}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\boldsymbol{\beta}^{\star}}\right)$
$=\frac{1}{p^{4}}\operatorname{\textnormal{Var}}\left(\sum_{i\neq j}(\beta_{j}^{\star})^{2}g(\beta_{i}^{\star})\right)$
$=\frac{1}{p^{4}}\sum_{i\neq j}\operatorname{\textnormal{Var}}(g(\beta_{i}^{\star})(\beta_{j}^{\star})^{2})+\frac{1}{p^{4}}\sum_{i_{!}\neq j_{1}}\sum_{i_{2}\neq j_{2}}\operatorname{\textnormal{Cov}}(g(\beta_{i_{1}}^{\star})(\beta_{j_{1}}^{\star})^{2},g(\beta_{i_{2}}^{\star})(\beta_{j_{2}}^{\star})^{2})$
$\lesssim p^{-2}+p^{-1}\to 0,$
where the last inequality follows from the fact that the covariance term is 0 if all the indices $i_{1},i_{2},j_{1},j_{2}$ are distinct. For the other term in (65), let us define
$A(i,j_{1},j_{2}):=\beta_{j_{1}}^{\star}\beta_{j_{2}}^{\star}g(\beta_{i}^{\star}),\,\,\,\,1\leq i,j_{1},j_{2}\leq p.$
We then note that
$\;\;\;\;\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}\operatorname{\textnormal{Var}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\boldsymbol{\beta}^{\star}}\right)$
$=\frac{1}{n^{4}}\sum_{k_{1},k_{2}=1}^{n}\sum_{i=1}^{p}\sum_{j_{1},j_{2}\neq i}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}[A^{2}(i,j_{1},j_{2})]\operatorname{\textnormal{Var}}(Z_{k_{1},i}Z_{k_{1},j_{1}}Z_{k_{2},i}Z_{k_{2},j_{2}})$
$\qquad+\frac{1}{n^{4}}\sum_{k_{1},k_{2},k_{3},k_{4}=1}^{n}\sum_{i_{1},i_{2}=1}^{p}\sum_{\begin{subarray}{c}j_{1},j_{2}\neq i_{1},\\ j_{3},j_{4}\neq i_{2}\end{subarray}}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}[A(i_{1},j_{1},j_{2})A(i_{2},j_{3},j_{4})]\operatorname{\textnormal{Cov}}(Z_{k_{1},i_{1}}Z_{k_{1},j_{1}}Z_{k_{2},i_{1}}Z_{k_{2},j_{2}},$
$\qquad\qquad Z_{k_{3},i_{2}}Z_{k_{3},j_{3}}Z_{k_{4},i_{2}}Z_{k_{4},j_{4}}).$
As $\operatorname{\mathbb{E}}A^{2}(i,j_{1},j_{2})<\infty$ uniformly, for $j_{1},j_{2}\neq i$, we have:
$\frac{1}{n^{4}}\sum_{k_{1},k_{2}=1}^{n}\sum_{i=1}^{p}\sum_{j_{1},j_{2}\neq i}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}[A^{2}(i,j_{1},j_{2})]\operatorname{\textnormal{Var}}(Z_{k_{1},i}Z_{k_{1},j_{1}}Z_{k_{2},i}Z_{k_{2},j_{2}})=\frac{p^{3}}{n^{2}}=o(p^{4}/n^{2}).$
To bound the covariance part, we will first isolate terms for which
$\operatorname{\textnormal{Cov}}(Z_{k_{1},i_{1}}Z_{k_{1},j_{1}}Z_{k_{2},i_{1}}Z_{k_{2},j_{2}},Z_{k_{3},i_{2}}Z_{k_{3},j_{3}}Z_{k_{4},i_{2}}Z_{k_{4},j_{4}})$
is not equal to 0. If the set $\{k_{1},k_{2},k_{3},k_{4}\}$ consists of $3$ or more distinct elements, then clearly the above covariance term equals 0. If $|\{k_{1},k_{2},k_{3},k_{4}\}|=2$, once again if $k_{1}=k_{2}$, $k_{3}=k_{4}$, but $k_{1}\neq k_{3}$, then the covariance term will be 0. So let us consider the case $k_{1}=k_{3}$, $k_{2}=k_{4}$. In that case, the covariance term reduces to
$\operatorname{\textnormal{Cov}}(Z_{k_{1},i_{1}}Z_{k_{1},j_{1}}Z_{k_{2},i_{1}}Z_{k_{2},j_{2}},Z_{k_{1},i_{2}}Z_{k_{1},j_{3}}Z_{k_{2},i_{2}}Z_{k_{2},j_{4}}).$
Now if $i_{1}\neq i_{2}$, then simple calculation shows that the above covariance term can be non-zero only when $j_{3}=j_{4}=i_{1}$ and $j_{1}=j_{2}=i_{2}$. The contribution of such terms is given by
$\frac{1}{n^{4}}\sum_{k_{1}\neq k_{2}}\sum_{i_{1}\neq i_{2}}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}[A(i_{1},i_{2},i_{2})A(i_{2},i_{1},i_{1})]\operatorname{\textnormal{Var}}(Z_{k_{1},i_{1}}Z_{k_{1},i_{2}}Z_{k_{2},i_{1}}Z_{k_{2},i_{2}})\lesssim\frac{p^{2}}{n^{2}}=o(p^{4}/n^{2}).$
Next suppose $i_{1}=i_{2}$. Then we can check that the covariance term will be 0 if either $j_{1}\neq j_{3}$ or $j_{2}\neq j_{4}$. Therefore, to get non-zero covariance we must have $j_{1}=j_{3}$ and $j_{2}=j_{4}$. The contribution of such terms is given by
$\frac{1}{n^{4}}\sum_{k_{1}\neq k_{2}}\sum_{i=1}^{p}\sum_{j_{1},j_{2}\neq i}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}[A^{2}(i,j_{1},j_{2})]\operatorname{\textnormal{Var}}(Z_{k_{1},i}Z_{k_{1},j_{1}}Z_{k_{2},i}Z_{k_{2},j_{2}})=\frac{p^{3}}{n^{2}}=o(p^{4}/n^{2}).$
The final remaining case is when $|\{k_{1},k_{2},k_{3},k_{4}\}|=1$. In this case, it is not hard to check that unless $|\{i_{1},i_{2},j_{1},j_{2},j_{3},j_{4}\}|\leq 4$, the covariance term will be 0. The contribution of such terms is given by
$\;\;\;\;\frac{1}{n^{4}}\sum_{k=1}^{n}\sum_{\begin{subarray}{c}i_{1},i_{2},(j_{1},j_{2})\neq i_{1},(j_{3},j_{4})\neq i_{2},\\ |\{i_{1},i_{2},j_{1},j_{2},j_{3},j_{4}\}|\leq 4\end{subarray}}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}[A(i_{1},j_{1},j_{2})A(i_{2},j_{3},j_{4})]\operatorname{\textnormal{Cov}}(Z_{k,i_{1}}^{2}Z_{k,j_{1}}Z_{k,j_{2}},Z_{k,i_{2}}^{2}Z_{k,j_{3}}Z_{k,j_{4}})$
$\lesssim\frac{p^{4}}{n^{3}}=o(p^{4}/n^{2}).$
This establishes (65) and completes the proof. ∎
Appendix D Simulation details
Parameter space
For each parametric family defined in Section 5 from the main text, we consider the following compact parameter space $\Theta$:
  1. (1)
    for $\textsf{Ber}(\pi)$, consider $\pi\in[0,1]$,
  2. (2)
    for $\textsf{Spike-Slab}(\pi,\tau^{2})$ consider $\pi\in[0,1],\tau\in[0.2,5]$,
  3. (3)
    for $\textsf{location-GMM}(\theta_{1},\theta_{2})$, consider $\theta_{1},\theta_{2}\in[-2,2]$.
Implementation details
The three methods are implemented as follows.
  • Variational EB: We computed the vEB estimator by using the optimize (one-dimensional) and optim (higher-dimensional) functions in R to solve the minimization (11). For the latter, we used the “L-BFGS-B” method with the above parameter space constraints.
  • Method of Moments: For each parametric family, we solve the following moment equations using the LSE $\hat{\beta}=(X^{\top}X)^{-1}X^{\top}y$:
    $\operatorname{\mathbb{E}}_{\textsf{Ber}}\beta=\pi,\qquad\operatorname{\mathbb{E}}_{\textsf{Spike-Slab}}\big[\beta^{2},~\beta^{4}\big]=(1-\pi)\big[\tau^{2},~3\tau^{4}\big],$
    $\operatorname{\mathbb{E}}_{\textsf{location-GMM}}\big[\beta,~\beta^{2}\big]=\frac{1}{2}\big[\theta_{1}+\theta_{2},~\theta_{1}^{2}+\theta_{2}^{2}+1.25\big].$
    The final estimates are thresholded to take values in the compact parameter space.
  • Langevin diffusion: We adaptively sample the prior parameters $\theta=(\theta_{1},\theta_{2})$ and regression coefficients $\beta$ $4000$ times via the discretized Langevin algorithm in [21]. The algorithm is initialized under the constraints $\theta_{1}<\theta_{2}$, with the default learning rate $10^{-2}$ (for both $\theta,\beta$). The final estimates are thresholded to take values in the compact parameter space.
The Mean-Field approximation in (7) is computed by iterating the fixed point equation for $u$ in 2.2. To evaluate coverage guarantees under the Oracle posterior $\operatorname{\mathbb{P}}_{\theta_{0}}$, we draw samples from $\operatorname{\mathbb{P}}_{\theta_{0}}$ as follows.
  • Bernoulli: We use the IsingSampler package in R [19], using the Metropolis-Hastings algorithm. We collect 500 samples.
  • Spike-Slab, GMM: For both models, we implemented a parallel Gibbs sampler (initialized at the true coefficients $\beta^{\star}$) and collected 5000 samples. We used the last half of the samples to evaluate the coverage probabilities.
Computation time
We report the average computation time per scenario in Table D.1 (which corresponds to Table 1). Across all settings, the vEB estimator is the fastest. In particular, the MoM estimator suffers in high-dimensions, which is a consequence of having to evaluate an matrix inversion to compute the least squares estimator. In contrast, the vEB estimator just has to solve a fixed-dimensional optimization.
Prior Ber Spike-Slab location-GMM
$p$ \ Algo vEB MoM vEB MoM vEB MoM Langevin
25 0.001 0.001 0.001 0.001 0.001 0.001 0.21
50 0.001 0.009 0.002 0.005 0.001 0.006 0.20
100 0.005 0.069 0.009 0.103 0.007 0.093 0.34
200 0.059 1.103 0.053 1.278 0.051 1.580 0.58
Table D.1: Elapsed time (in seconds) for varying $p$ and $n=p^{2}$ (averaged over 400 replications). The Langevin estimator is implemented in Python while all other estimators (vEB and MoM) are implemented in R.
Additional results
In Table D.2, we additionally report the accuracy of the MoM and Langevin estimators under an uninformative random initialization. As mentioned as a footnote in the main paper, these estimators suffer from fundamental non-identifiability and multiple local maximizers, respectively. We see that the MoM estimator results in an inconsistent result whereas the Langevin estimator only mildly suffers compared to the analogous results in Table 1.
Table D.2 also reports simulations results under the heavy-tailed Cauchy prior with location parameter $\theta$. While our theory does not allow this due to a violation of the moment condition R5, the results indicate consistency but with an empirical convergence rate slower than $\sqrt{p}$, and it would be interesting to investigate whether the convergence rate fundamentally suffers under heavy-tailed priors.
[…]
$p$ \ Algo MoM Langevin vEB
25 0.587 0.730 0.832
50 0.623 0.422 0.584
100 0.596 0.192 0.366
200 0.589 0.093 0.342
Table D.2: MSE $\|\hat{\theta}^{\mathsf{vEB}}-\theta_{0}\|^{2}$ for varying $p$ and $n=p^{2}$ (averaged over 400 replications). The location-GMM results are computed completely uninformatively.